### v1

# On the Optimality of Gaussian Code-books for Signaling over a Two-Users Weak Gaussian Interference Channel

Abstract: This article shows that the capacity region of a 2-users weak Gaussian interference channel is achieved using Gaussian code-books. The approach relies on traversing the boundary in incremental steps. Starting from a corner point with Gaussian code-books, and relying on calculus of variation, it is shown that the end point in each step is achieved using Gaussian code-books.
1 Introduction
Consider a two-users weak Gaussian interference channel with parameters shown in Fig. 1.
Figure 1: Two-users Gaussian Interference Channel (GIC) with $a<1$ and $b<1$.
Boundary is traversed by changing the power allocation between public and private message(s), refereed to as “power reallocation” hereafter. Each step starts from a point on the boundary, and then optimum code-books are found such that the corresponding step ends in another point on the boundary. Power reallocation values corresponding to such a step satisfy
[…]
With some misuse of notations, hereafter power reallocation vectors are denoted as
$(\delta P_{1},\delta P_{2})\leavevmode\nobreak\ \mbox{where}\leavevmode\nobreak\ \leavevmode\nobreak\ \delta P_{1}=|\kappa_{1}|=|\eta_{1}|,\leavevmode\nobreak\ \leavevmode\nobreak\ \delta P_{2}=|\kappa_{2}|=|\eta_{2}|.$ (3)
In other words, $\delta P_{1}$ denotes the increase in the power of $U_{1}$ or $V_{1}$, depending on which of the two has a higher power at the end point vs. the starting point, and likewise for $\delta P_{2}$ in relation to $U_{2}$ and $V_{2}$. Figure 2 depicts an example where notations $\pm$ vs. $\mp$ are used to emphasize that the signs of $\delta P_{1}$ and $\delta P_{2}$ depend on the step and power reallocation is zero-sum. Power reallocation vector is selected to: (i) support a counter-clockwise move along the boundary, and (ii) guarantee the solution achieving each end point is unique. To achieve the latter criterion, while moving continuously along the boundary, power reallocation vector is selected relying on a notation of admissibility called Pareto minimal (see Theorem 10), or relying on a milder condition in which the power reallocation vector is linearly increased (see Theorem 11). Referring to Fig. 2, for a given a power reallocation vector $(\delta P_{1},\delta P_{2})$, the following measure of optimality is used in selecting code-books’ density functions: Given $(\delta P_{1},\delta P_{2})$, maximize the length of the step, i.e., $\mathbf{\Gamma}$, over all possible values of the slope $\mathbf{\Upsilon}$.
Figure 2: An example for power reallocation and its corresponding step along the boundary.
Remark 1: It is known that capacity region of two-users Gaussian Interference Channel (GIC) may include segments achieved by power allocation among different two-users GICs, called component GICs hereafter. The overall capacity region is obtained by computing the convex hull of regions corresponding to all possible dividing of power among component GICs. The optimum allocation of power among component GICs, to enlarge the convex hull, is not discussed here. In other words, this article restricts the power constraint for each user to be satisfied with equality, resulting in a single component GIC. Forcing power constraints to be satisfied with equality may result in code-books with a non-zero mean (to limit the impact of the interference). Results are established which guarantee optimum code-books’ density functions are zero mean. Under these conditions, it is shown that boundary points for a single component GIC are achieved using unique zero mean Gaussian code-books. $\blacksquare$
2 Literature Survey
The problem of Gaussian interference channel has been the subject of numerous outstanding prior works, paving the way to the current point and moving beyond. A subset of these works, reported in [2] to [40], are briefly discussed in this section. A more complete and detailed literature survey will be provided in subsequent revisions of this article
Reference [2] discusses degraded Gaussian interference channel (degraded means one of the two receivers is a degraded version of the other one) and presents multiple bounds and achievable rate regions. Reference [3] studies the capacity of 2-users GIC for the class of strong interference and shows the capacity region is at the intersection of two MAC regions, consistent with the current article. Reference [4] establishes optimality for two extreme points in the achievable region of the general 2-users GIC. [4] also proves that the class of degraded Gaussian interference channels is equivalent to the class of Z (one-sided) interference channels.
References [5] to [7] present achievable rate regions for interference channel. In particular, [5] presents the well-known Han-Kobayashi (HK) achievable rate region. HK rate region coincides with all results derived previously (for Gaussian 2-users GIC), and is shown to be optimum for the class of weak 2-users GIC in the current article. References [8][10] have further studied the HK rate region. [10] shows that HK achievable rate region is strictly sub-optimum for a class of discrete interference channels.
References [11] to [17] have studied the problem of outer bounds for the interference channel. Among these, [13][14][15] have also provided optimality results in some special cases of weak 2-users GIC.
References [18][19] have studied the problem of interference channel with common information. References [20] to [22] have studied the problem of interference channel with cooperation between transmitters and/or between receivers. References [23][24] have studied the problem of interference channel with side information. Reference [25] has studied the problem of interference channel assuming cognition, and reference [26] has studied the problem assuming cognition, with or without secret messages.
Reference [27] has found the capacity regions of vector Gaussian interference channels for classes of very strong and aligned strong interference. [27] has also generalized some known results for sum-rate of scalar Z interference, noisy interference, and mixed interference to the case of vector channels. Reference [28] has addressed the sum-rate of the parallel Gaussian interference channel. Sufficient conditions are derived in terms of problem parameters (power budgets and channel coefficients) such that the sum-rate can be realized by independent transmission across sub-channels while treating interference as noise, and corresponding optimum power allocations are computed. Reference [29] studies a Gaussian interference network where each message is encoded by a single transmitter and is aimed at a single receiver. Subject to feeding back the output from receivers to their corresponding transmitter, efficient strategies are developed based on the discrete Fourier transform signaling.
Reference [30] computes the capacity of interference channel within one bit. References [31][32] study the impact of interference in GIC. [32] shows that treating interference as noise in 2-users GIC achieves the closure of the capacity region to within a constant gap, or within a gap that scales as O(log(log(.)) with signal to noise ratio. Reference [33] relies on game theory to define the notion of a Nash equilibrium region of the interference channel, and characterizes the Nash equilibrium region for: (i) 2-users linear deterministic interference channel in exact form, and (ii) 2-users GIC within 1 bit/s/Hz in an approximate form.
Reference [34] studies the problem of 2-users GIC based on a sliding window superposition coding scheme.
References [35] and [36], independently, introduce the new concept of non-unique decoding as an intermediate alternative to “treating interference as noise”, or “canceling interference”. Reference [37] further studies the concept on non-unique decoding and proves that (in all reported cases) it can be replaced by a special joint unique decoding without penalty.
Reference [38] studies the degrees of freedom of the K-user Gaussian interference channel, and, subject to a mild sufficient condition on the channel gains, presents an expression for the degrees of freedom of the scalar interference channel as a function of the channel matrix.
Reference [39] studies the problem of state-dependent Gaussian interference channel, where two receivers are affected by scaled versions of the same state. The state sequence is (non-causally) known at both transmitters, but not at receivers. Capacity results are established (under certain conditions on channel parameters) in the very strong, strong, and weak interference regimes. For the weak regime, the sum-rate is computed. Reference [40] studies the problem of state-dependent Gaussian interference channel under the assumption of correlated states, and characterizes (either fully or partially) the capacity region or the sum-rate under various channel parameters.
3 Problem Formulation
Consider a two-users weak Gaussian interference channel with inputs $X_{1}$, $X_{2}$ and outputs $Y_{1}$, $Y_{2}$, defined as
$Y_{1}$ $=$ $X_{1}+\sqrt{b}X_{2}+Z_{1}$
$Y_{2}$ $=$ $\sqrt{a}X_{1}+X_{2}+Z_{2}$ (4)
where $a,b<1$ and $Z_{1}$, $Z_{2}$ are additive white Gaussian noise terms of zero mean and unit variance. Finding the corresponding capacity region narrows down to:
[…]
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ P_{U_{2}}+P_{V_{2}}$ $=$ $P_{2}.$ (5)
Solving optimization problem in 5 entails: (i) For each user, allocating the power to public and private messages, called power allocation. (ii) Finding the optimum density functions for each message code-book. (iii) Finding encoding/decoding procures for each user. The term coding strategy is used to specify encoding/decoding procedures for each user at a respective point on the boundary.
Capacity region is traversed by starting from the point with maximum $R_{1}$ and moving counterclockwise along the lower part of the boundary, i.e., for $\mu<1$. It is known that the point maximizing $R_{1}$ is achieved using Gaussian code-books, where message $X_{1}$ is entirely private, message $X_{2}$ is entirely public, $Y_{1}$ uses successive decoding and $Y_{2}$ treats the interference as noise. Starting from the point with maximum $R_{1}$, in a sequence of infinitesimal steps, $R_{2}$ is gradually increased at the expense of reducing $R_{1}$. Each step involves changing the power allocation values by infinitesimal amounts. Amounts of reallocated power, $\delta P_{1}$ and $\delta P_{2}$, are small enough such that the coding strategy does not change within the step (can potentially change at the start of the next step).
Let us consider an infinitesimal step from a starting point, specified by superscript $s$, to an end point specified by superscript $e$. The slope $\mathbf{\Upsilon}$ of such a step defined as
$\mathbf{\Upsilon}=\frac{\Delta\!R_{2}}{\Delta\!R_{1}}=\frac{R^{e}_{V_{2}}+R^{e}_{U_{2}}-R^{s}_{V_{2}}-R^{s}_{U_{2}}}{R^{s}_{V_{1}}+R^{s}_{U_{1}}-R^{e}_{V_{1}}-R^{e}_{U_{1}}}\triangleq\frac{\mathbf{N}}{\mathbf{D}}$ (6)
where $(R^{s}_{U_{1}},R^{s}_{V_{1}})$, $(R^{s}_{U_{2}},R^{s}_{V_{2}})$ are public and private rates of user 1 and user 2, respectively, at the starting point, likewise, $(R^{e}_{U_{1}},R^{e}_{V_{1}})$, $(R^{e}_{U_{2}},R^{e}_{V_{2}})$ are public and private rates at the end point. Note that $\Delta\!R_{1}$ and $\Delta\!R_{2}$ are defined to be positive, in particular $\Delta\!R_{1}$ is defined as the rate $R_{1}$ at the starting point, minus the rate $R_{1}$ at the end point. Optimality of boundary points is captured in $\mathbf{\Gamma}$ defined as
$\mathbf{\Gamma}=\sqrt{(\Delta\!R_{1})^{2}+(\Delta\!R_{2})^{2}}.$ (7)
This work focuses on $\mu<1$ by starting from a point with maximum $R_{1}$ and moving counterclockwise along the boundary. The case of $\mu>1$ follows similarly by starting from a point with maximum $R_{2}$ and moving clockwise along the boundary. The case of $\mu=1$ is obtained by time sharing between the end points for segments corresponding to $\mu<1$ and $\mu>1$. Hereafter, $U_{1}$, $U_{2}$, $V_{1}$, $V_{2}$ are called core random variables. Linear combinations of core random variables appearing in mutual information terms forming 6 and 7 are called composite random variables.
Remark 2: The problem of finding the capacity region is complex, since: (i) Power reallocation affects the selection of code-books’ densities. (ii) The value of weight $\mu$ changes as one moves along the boundary. (iii) One needs to define the infinitesimal steps such that the boundary is covered continuously, and there are unique optimizing code-books for each boundary point. This article does not claim that the coding strategy and its associated code-books’ densities (including power allocation) for realizing an achievable rate pair $(R_{1},R_{2})$ are unique, nor that corresponding density functions are limited to be zero-mean Gaussian. The main result to be established is as follows: For power reallocation vectors which satisfy condition of Theorem 10, or a milder condition of Theorem 11, zero-mean Gaussian code-books for public and private messages provide a unique solution maximizing $\gamma$ for $\mathbf{\Upsilon}=\upsilon$ in $(\mathbf{\Upsilon},\mathbf{\Gamma})=(\upsilon,\gamma)$. This results in a unique point on the boundary. $\blacksquare$
4 Main Results
Theorems 1 shows that, starting from any point on the boundary and moving counterclockwise for $\mu<1$, the value of $\mathbf{\Upsilon}$ in 6 is non-increasing, and the value of $\mathbf{\Gamma}$ in 7 is monotonically increasing. Theorems 2 and 3 show that, in 6 and 7, due to successive decoding in at least one of the receivers, each composite random variable contributes to an entropy term of the form appearing in successive decoding over an additive noise channel. Theorem 4 shows that there is a system of invertible linear equations relating composite random variables to core random variables. This means each core random variable can be expressed as a (unique) linear combination of composite random variables. Theorem 5 establishes that a stationary solution in optimizing a weighted sum of entropy terms, obtained using calculus of variation, is either global maximum or global minimum. Theorem 6 shows that, given $(P_{1},P_{2})$, the dividing of power between public and private messages of each user is such that the mean of each code-book will be zero. As a result, the application of calculus of variation is formulated in terms of zero-mean destinies. Theorem 7 shows that there is a single power allocation achieving a point on the boundary for a given $\mu$. Theorem 8 shows that, to achieve a stationary solution, each composite random variable should have a zero mean Gaussian density. Since, from Theorem 4, there is a one-to-one linear mapping between core and composite random variables, it follows that core random variables will be zero mean Gaussian as well. It remains to impose a condition on power reallocation vector such that each end point is achieved in a unique manner, and the boundary can be traversed in a continuous manner starting from any end point. Such power allocation is called boundary achieving hereafter. Theorems 9, 10 and 11 address this issue for all boundary points with $\mu<1$. Note that $\mu<1$ entails $\mathbf{\Upsilon}<1$.
Theorem 1 establishes how $\mathbf{\Gamma}$ and $\mathbf{\Upsilon}<1$ change as one moves counterclockwise along the boundary.
Theorem 1.
For $\mu<1$, consider a set of consecutive steps, in counterclockwise direction, with end points that fall on the boudnary. Corresponding values for $\mathbf{\Upsilon}$ in 6 will be monotonically decreasing, while $\mathbf{\Gamma}$ in 7 will be monotonically increasing.
Proof.
Proof follows noting that: (1) the capacity region is convex, and (2) lower part starts from a point with maximum $R_{1}$. Let us consider two consecutive infinitesimal steps from point $\mathbb{U}$ to point $\mathbb{V}$ and from point $\mathbb{V}$ to point $\mathbb{W}$. Let us assume $\Delta R_{1}$ for the first and second steps are equal to $\delta$, and corresponding $\Delta R_{2}$ values are equal to $\hat{\delta}$ and $\check{\delta}$, respectively. Since the boundary is continuous, it is possible to form such two consecutive steps. Noting that the lower part of the boudnary starts from a point with maximum $R_{1}$, and then moves counter-clock wise, we can conclude
$\delta>0,\leavevmode\nobreak\ \leavevmode\nobreak\ \hat{\delta}>0,\leavevmode\nobreak\ \leavevmode\nobreak\ \check{\delta}>0.$ (8)
Noting boundary is convex, for $\mu<1$, we have
$\hat{\delta}>\check{\delta}\Longrightarrow\frac{\hat{\delta}}{\delta}>\frac{\check{\delta}}{\delta}$ (9)
otherwise, $\mathbb{V}$ would fall strictly inside the capacity region. From 6, 7, 8 and 9, it follows that
$\hat{\mathbf{\Upsilon}}$ $>$ $\check{\mathbf{\Upsilon}}$ (10)
$\hat{\mathbf{\Gamma}}$ $>$ $\check{\mathbf{\Gamma}}$ (11)
where $(\hat{\mathbf{\Upsilon}},\hat{\mathbf{\Gamma}})$ and $(\check{\mathbf{\Upsilon}},\check{\mathbf{\Gamma}})$ correspond to the first step and the second step, respectively. ∎
Theorem 2 presents results that will be used in Theorem 3 to estabslih some inequalities on mutual information terms. The aim is to determine conditions where $U_{1}$ and $U_{2}$ should be jointly decoded.
Theorem 2.
In successive decoding, message $U_{2}$ at $Y_{1}$ is a degraded version of message $U_{2}$ at $Y_{2}$, message $U_{1}$ at $Y_{2}$ is a degraded version of message $U_{1}$ at $Y_{1}$, message $U_{1}$ at $Y_{2}$ after decoding of $U_{2}$ is a degraded version of message $U_{1}$ at $Y_{1}$ after decoding of $U_{2}$, and message $U_{2}$ at $Y_{1}$ after decoding of $U_{1}$ is a degraded version of message $U_{2}$ at $Y_{2}$ after decoding of $U_{1}$.
Proof.
[…]
Figure 3: Configurations used in the proof of Theorems 2 and 3. Notation $\eth\trianglerighteq\wp$ means capacity of a channel with input $\mathfrak{I}$ and additive noise term $\eth$ (independent of $\mathfrak{I}$) is smaller as compared to that of a channel with input $\mathfrak{I}$ and additive noise term Weierstrass-$\wp$ (independent of $\mathfrak{I}$).
Theorem 3.
In at least one of the receivers, $Y_{1}$ and/or $Y_{2}$, public messages $U_{1}$ and $U_{2}$ are recovered using successive decoding.
Proof.
First, note that regardless of code-books’ densities and the method used in recovering $U_{1}$ and $U_{2}$, i.e., joint or successive decoding, we have
$R_{U_{1}}$$\leq$ $I(U_{1};Y_{1}|U_{2})$ (12)
$R_{U_{1}}$$\leq$ $I(U_{1};Y_{2}|U_{2})$ (13)
$R_{U_{2}}$$\leq$ $I(U_{2};Y_{1}|U_{1})$ (14)
$R_{U_{2}}$$\leq$ $I(U_{2};Y_{2}|U_{1})$ (15)
$R_{U_{1}}+R_{U_{2}}$$\leq$ $I(U_{1},U_{2};Y_{1})$ (16)
$R_{U_{1}}+R_{U_{2}}$$\leq$ $I(U_{1},U_{2};Y_{2}).$ (17)
Conditions of this theorem will be violated if $U_{1}$ and $U_{2}$ should be jointly decoded at both $Y_{1}$ and $Y_{2}$. This corresponds to the necessary condition
$R_{U_{1}}+R_{U_{2}}=I(U_{1},U_{2};Y_{1})=I(U_{1},U_{2};Y_{2}).$ (18)
Applying chain rule for mutual information to 18, we have
$R_{U_{1}}+R_{U_{2}}$ $=$ $I(U_{2};Y_{1})+I(U_{1};Y_{1}|U_{2})$ (19)
$=$ $I(U_{1};Y_{1})+I(U_{2};Y_{1}|U_{1})$ (20)
$=$ $I(U_{1};Y_{2})+I(U_{2};Y_{2}|U_{1})$ (21)
$=$ $I(U_{2};Y_{2})+I(U_{1};Y_{2}|U_{2}).$ (22)
Noting Fig. 3, from Theorem 2, we have
[…]
Note that by swapping
$U_{1}\longleftrightarrow U_{2}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ Y_{1}\longleftrightarrow Y_{2}$ (25)
one can reach from 23 to 24, and vice versa. As a result, without loss of generality, let us focus on 23. From 19, 22 and 23, we conclude
$I(U_{1};Y_{1}|U_{2})\geq I(U_{1};Y_{2}|U_{2}).$ (26)
From 13 and 26, and noting that public messages should be decoded at both $Y_{1}$ and $Y_{2}$ prior to decoding of private messages, we have
$R_{U_{1}}\leq I(U_{1};Y_{2}|U_{2})\leq I(U_{1};Y_{1}|U_{2}).$ (27)
From 22 and 27, we obtain
$R_{U_{1}}$ $=$ $I(U_{1};Y_{2}|U_{2})-\varrho$ (28)
$R_{U_{2}}$ $=$ $I(U_{2};Y_{2})+\varrho\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{where}\leavevmode\nobreak\ \leavevmode\nobreak\ \varrho\geq 0.$ (29)
The method of decoding for recovering $U_{1}$ and $U_{2}$ at $Y_{1}$ does not affect $R_{V_{1}}$, likewise, method of decoding for recovering $U_{1}$ and $U_{2}$ at $Y_{2}$ does not affect $R_{V_{2}}$. This means
$R_{V_{1}}$ $=$ $I(V_{1};Y_{1}|U_{1},U_{2})$ (30)
$R_{V_{2}}$ $=$ $I(V_{2};Y_{2}|U_{1},U_{2}).$ (31)
From 28, 29, 30 and 31, we obtain
$R_{1}+\mu R_{2}=I(V_{1};Y_{1}|U_{1},U_{2})+\mu I(V_{2};Y_{2}|U_{1},U_{2})+I(U_{1};Y_{2}|U_{2})+\mu I(U_{2};Y_{2})-(1-\mu)\varrho.$ (32)
For $\mu<1$, we have $(1-\mu)>0$ and 32 is maximized by: (1) selecting $\varrho=0$, and (2) maximizing
$I(V_{1};Y_{1}|U_{1},U_{2})+\mu I(V_{2};Y_{2}|U_{1},U_{2})+I(U_{1};Y_{2}|U_{2})+\mu I(U_{2};Y_{2}).$ (33)
For $\varrho=0$, we have
$R_{U_{2}}$ $=$ $I(U_{2};Y_{2})$ (34)
$R_{U_{1}}$ $=$ $I(U_{1};Y_{2}|U_{2})$ (35)
$R_{U_{1}}+R_{U_{2}}$ $=$ $I(U_{1},U_{2};Y_{1}).$ (36)
Expressions 34, 35, 36 indicate that: (i) $U_{1}$ and $U_{2}$ are successively decoded at $Y_{2}$ where $U_{2}$ is decoded prior to $U_{1}$, and (ii) $U_{1}$ and $U_{2}$ are jointly decoded at $Y_{1}$. Conditions of Theorem 3 are depicted in Fig. 3(b). In summary, for $\mu<1$, the corner point $\mathbb{Q}$ in Fig. 3(b) is superior to the middle point $\mathbb{M}$. Relying on 24 instead of 23, point $\mathbb{R}$ in Fig. 3(b) establishes a similar result for $\mu>1$. ∎
Remark 3: Since conclusions in Theorem 2 are valid for all power allocations, it follows that the structure shown in Fig. 3(b) is valid for all points on the boundary where both users have public messages. This means, for $(U_{1}\neq 0,U_{2}\neq 0)$, point $\mathbb{Q}$, using joint decoding at $Y_{1}$ and successive decoding at $Y_{2}$, optimizes $R_{1}+\mu R_{2}$ for $\mu<1$, likewise, point $\mathbb{R}$, using joint decoding at $Y_{2}$ and successive decoding at $Y_{1}$, optimizes $R_{1}+\mu R_{2}$ for $\mu>1$. On the other hand, for initial parts of the lower boundary, we have $U_{1}=0$. In such a case, additive noise term (g) in Fig. 3(a) governs the rate of $U_{2}$, which is determined by additive noise term at $Y_{1}$. Likewise, for initial parts on the upper boundary, i.e., starting from the point with maximum $R_{2}$ and moving clockwise for $\mu>1$, we have $U_{2}=0$ and the additive noise term (e) in Fig. 3(a) governs the rate of $U_{1}$. $\blacksquare$
Theorem 4 establishes that core random variables $U_{1}$,$V_{1}$,$U_{2}$,$V_{2}$ are a unique linear combination of composite random variables occurring in successive decoding at $Y_{1}$ or at $Y_{2}$. This property will be used to show that if such composite random variables are jointly Gaussian, then $U_{1}$, $V_{1}$, $U_{2}$, $V_{2}$ will be Gaussian as well.
Theorem 4.
There exits at least one invertible $4\times 4$ matrix allowing to express core random variables in terms of composite random variables.
Proof.
Let us focus on $\mu<1$, i.e., successive decoding of public messages is performed at $Y_{2}$. Consider composite random variables $C_{1}$ to $C_{4}$ involved in successive decoding at $Y_{2}$. We have
$C_{1}$ $=$ $\sqrt{a}U_{1}+\sqrt{a}V_{1}+U_{2}+V_{2}$ (37)
$C_{2}$ $=$ $\sqrt{a}U_{1}+\sqrt{a}V_{1}+V_{2}$ (38)
$C_{3}$ $=$ $\sqrt{a}V_{1}+V_{2}$ (39)
$C_{4}$ $=$ $\sqrt{a}V_{1}.$ (40)
Matrix of linear coefficients forming 37, 38, 39, 40 is equal to
$\begin{bmatrix}\sqrt{a}&\sqrt{a}&1&1\\ \sqrt{a}&\sqrt{a}&0&1\\ 0&\sqrt{a}&0&1\\ 0&\sqrt{a}&0&0\end{bmatrix}.$ (41)
It easily follows that the matrix in 41 is invertible $\forall a\neq 0$. Note that 41 also means power of composite and core random variables are related by an invertible matrix, obtained by changing $\sqrt{b}$ to $b$ in 41. Since the corresponding matrix is not block diagonal, it follows that a change in power allocation for user 1 (and/or for user 2) results in changing all the rate values.
For $a=0$, $b\neq 0$ (similarly for $a\neq 0$, $b=0$), both $Y_{1}$ and $Y_{2}$ rely on successive decoding. It follows that core random variables can be expressed as a unique linear combination of composite random variables. Again, since matrices defining linear combinations are not block diagonal, it follows that a change in the power allocation for user 1 (and/or for user 2) results in changing all rate values.
Finally, $a=b=0$ corresponds to the trivial case of parallel channels. In this case, the optimum power allocation is not unique since $(U_{1},V_{1})$, and $(U_{2},V_{2})$, each form a two-level Gaussian code-book. As a result, $(P_{U_{1}},P_{V_{1}})$ can take all values satisfying $P_{U_{1}}\geq 0,P_{V_{1}}\geq 0\!:\,P_{U_{1}}+P_{V_{1}}=P_{1}$, and likewise, $(P_{U_{2}},P_{V_{2}})$ can take all values satisfying $P_{U_{2}}\geq 0,P_{V_{2}}\geq 0\!:\,P_{U_{2}}+P_{V_{2}}=P_{2}$. ∎
Without loss of generality, let us assume that the decoding strategy in Theorem 4 for $a\neq 0$ and $b\neq 0$ applies throughout this article. This means $R_{U_{1}}$, $R_{V_{1}}$, $R_{U_{2}}$ are governed by a cascade of additive noise channels due to successive decoding at $Y_{2}$ and $R_{V_{1}}$ is governed by an additive noise channel at $Y_{1}$. As a result, rate values contributing to $\mathbf{\Upsilon}$, $\mathbf{\Gamma}$ in 6, 7, respectively, correspond to independent additive noise channels depicted in Fig. 4. Note that Theorem 4 includes all core random variables $U_{1},V_{1},U_{2},V_{2}$. A similar result concerning Gaussianity of core random variables follows if $U_{1}$ or $U_{2}$ is zero.
Figure 4: Channel models depicting decoding methods discussed in Theorem 4 (assuming $a\neq 0$ and $b\neq 0$) where 4(a) corresponds to successive decoding of $(U_{1},U_{2})$ followed by decoding of $V_{2}$ at $Y_{2}$, and 4(b) corresponds to joint decoding of $(U_{1},U_{2})$ followed by decoding of $V_{1}$ at $Y_{1}$.
Theorem 5 concerns applying calculus of variation to entropy terms appearing in such a weighted sum-rate. To motivate derivations that follow (using calculus of variation), let us consider independent probability density functions $f_{1}$, $f_{2}$ and $f_{3}$ which are zero-mean with variances $\vartheta_{1}$, $\vartheta_{2}$, $\vartheta_{3}$, respectively, and let us define functionals ϝ$\digamma_{\!\!1}$ and ϝ$\digamma_{\!\!2}$ as
ϝ$\digamma_{\!\!1}$ $=$ $f_{1}\ast f_{2}\ast f_{3}$ (42)
ϝ$\digamma_{\!\!2}$ $=$ $f_{2}\ast f_{3}$ (43)
where $\ast$ denotes convolution. Entropy terms for ϝ$\digamma_{\!\!1}$, ϝ$\digamma_{\!\!2}$, denoted as $\mathsf{H}^{\digamma_{\!\!1}}$, $\mathsf{H}^{\digamma_{\!\!2}}$, respectively, are
$\mathsf{H}^{\digamma_{\!\!1}}$ $=$ ϝ$-\int\digamma_{\!\!1}\log\digamma_{\!\!1}$ (44)
$\mathsf{H}^{\digamma_{\!\!2}}$ $=$ ϝ$-\int\digamma_{\!\!2}\log\digamma_{\!\!2}.$ (45)
Consider applying calculus of variation to
Maximize $\mathsf{H}^{\digamma_{\!\!1}}+\xi\mathsf{H}^{\digamma_{\!\!2}}$ (46)
Subject to : $\displaystyle\int x^{2}f_{1}=\vartheta_{1}$ (47)
$\displaystyle\int x^{2}f_{2}=\vartheta_{2}$ (48)
$\displaystyle\int x^{3}f_{3}=\vartheta_{3}$ (49)
$\displaystyle\int f_{1}=1$ (50)
$\displaystyle\int f_{2}=1$ (51)
$\displaystyle\int f_{3}=1$ (52)
where $\xi$ is a weight factor. As will become clear in later parts of this article, relying on calculus of variation, it is concluded that Gaussian densities result in a stationary solution for the optimization problem in 46 to 52. Theorem 5 establishes a key property of such a stationary solution.
Theorem 5.
In applying calculus of variation to find a stationary solution for the optimization problem in 46 to 52, the second order variation will be non-zero.
Proof.
Let us apply perturbations $\varpi_{1}h_{1}$, $\varpi_{2}h_{2}$, $\varpi_{3}h_{3}$ to $f_{1}$, $f_{2}$, $f_{3}$, respectively. Applying derivations similar to Appendix A.3.2, it is follows that the second order variation of the problem defined in 46 to 52 is equal to:
$\frac{h_{1}^{2}}{f_{1}\ast f_{2}\ast f_{3}}+\frac{h_{2}^{2}}{f_{1}\ast f_{2}\ast f_{3}}+\frac{h_{3}^{2}}{f_{1}\ast f_{2}\ast f_{3}}+\frac{h_{2}^{2}}{f_{2}\ast f_{3}}+\frac{h_{3}^{2}}{f_{2}\ast f_{3}}\neq 0,\leavevmode\nobreak\ \forall h_{1},h_{2},h_{3}.$ (53)
The summation of the first three terms in 53 are second order variation of $\mathsf{H}^{\digamma_{\!\!1}}$ and the summation of the last two terms are second order variation of $\mathsf{H}^{\digamma_{\!\!2}}$. Since perturbations $\varpi_{1}h_{1}$, $\varpi_{2}h_{2}$, $\varpi_{3}h_{3}$ are arbitrary functions, it follows that the second order variation in 53 is non-zero. ∎
Theorem 5 indicates that Gaussian densities either maximize or minimize functional $\mathsf{H}^{\digamma_{\!\!1}}+\xi\mathsf{H}^{\digamma_{\!\!2}}$ defined in 46. Since power constraints in 47, 48, 49 are forced to be satisfied with equality, such a solution may include cases that the Gaussian densities have a non-zero statistical mean. Following example aims to clarify this point.
Example:
Figure 5: Example of a channel where the stationary solution for mutual information may result in a maximum or a minimum, according to the statistical mean of $\tilde{Z}$.
[…]
$\hat{f}$ $=$ $f_{\tilde{X}}\ast f_{\tilde{Z}}\ast{\cal N}(0,1)$ (54)
$=$ $f_{\tilde{Z}}+{\cal N}(0,2)$ (55)
$\check{f}$ $=$ $f_{\tilde{Z}}\ast{\cal N}(0,1)$ (56)
where ${\cal N}(\mathfrak{u},\mathfrak{s})$ is a Gaussian density with statistical average $\mathfrak{u}$ and variance $\mathfrak{s}$. We have
$I(\tilde{X};\tilde{Y})=\mathsf{H}^{\hat{f}}-\mathsf{H}^{\check{f}}.$ (57)
It follows that
$\min_{\hat{f},\check{f}}I(\tilde{X};\tilde{Y})\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{is achieved for}$ $\!\!\!\!\!\!\!\!\!\!f_{\tilde{Z}}={\cal N}(0,P_{\!\tilde{Z}})$ (58)
$\max_{\hat{f},\check{f}}I(\tilde{X};\tilde{Y})\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{is achieved for}$ $\!\!f_{\tilde{Z}}={\cal N}(\sqrt{P_{\!\tilde{Z}}},0).$ (59)
A non-zero statistical mean entails the power $P_{\!\tilde{Z}}$ is intentionally wasted to avoid interference. $\blacksquare$
Note that the reason for having two stationary solutions for 57, one being the minimum and the other one being the maximum, is the possibility of having code-books’ densities with non-zero means. This observation does not contradict the statements of theorems 10 and 11 concerning the uniqueness of the stationary solution achieved using Gaussian code-books. Indeed, Theorem 6 establishes that code-books densities for $U_{1}$, $V_{1}$, $U_{2}$ and $V_{2}$ are zero-mean. This entails a case similar to the above example will not be encountered in code-books’ densities forming the capacity region in this work.
Theorem 6.
Code-books’ densities for $U_{1}$, $V_{1}$, $U_{2}$, $V_{2}$ are zero mean.
Proof.
First, note that code-books’ densities for public messages $U_{1}$ and $U_{2}$ are zero mean. The reason is that, instead of wasting the allocated power values $P_{U_{1}}$ and/or $P_{U_{2}}$ relying on a non-zero mean value, the variance of corresponding code-book(s) can be increased which in turn increases $R_{U_{1}}$ and/or $R_{U_{2}}$ while satisfying the condition that public messages should be recoverable at both receivers. On the other hand, if the code-books’ densities for private messages $V_{1}$ and/or $V_{2}$ have a non-zero mean, the wasted power can be allocated to the corresponding public message, increasing $R_{U_{1}}$ and/or $R_{U_{2}}$, while guaranteeing public and private messages can be decoded. ∎
Relying on Theorem 6, all optimization problems involving application of calculus of variation are formulated in terms of zero-mean destinies.
Theorem 7.
There is a single power allocation achieving a point on the boundary for any given $\mu$.
Proof.
Given $\mu<1$, let us use $\mathsf{W}(\mu)$ to refer to the corresponding optimum weighted sum-rate, i.e.,
$\mathsf{W}(\mu)\equiv\max(R_{1}+\mu R_{2}).$ (60)
Consider the following two power allocations for users $X_{1}$, $X_{2}$, refereed to as $1^{st}$ and $2^{nd}$, and distinguished by superscripts 1,2:
$1^{st}\mbox{\leavevmode\nobreak\ power allocation for user\leavevmode\nobreak\ }X_{1}:(P^{1}_{U_{1}},P^{1}_{V_{1}})=(\mathsf{t}^{1}_{1},1-\mathsf{t}^{1}_{1})P_{1}$$\equiv$ $\mathbf{p}_{1}^{1}$ (61)
$1^{st}\mbox{\leavevmode\nobreak\ power allocation for user\leavevmode\nobreak\ }X_{2}:(P^{1}_{U_{2}},P^{1}_{V_{2}})=(\mathsf{t}^{1}_{2},1-\mathsf{t}^{1}_{2})P_{2}$$\equiv$ $\mathbf{p}_{2}^{1}$ (62)
$2^{nd}\mbox{\leavevmode\nobreak\ power allocation for user\leavevmode\nobreak\ }X_{1}:(P^{2}_{U_{1}},P^{2}_{V_{1}})=(\mathsf{t}^{2}_{1},1-\mathsf{t}^{2}_{1})P_{1}$$\equiv$ $\mathbf{p}_{1}^{2}$ (63)
$2^{nd}\mbox{\leavevmode\nobreak\ power allocation for user\leavevmode\nobreak\ }X_{2}:(P^{2}_{U_{2}},P^{2}_{V_{2}})=(\mathsf{t}^{2}_{2},1-\mathsf{t}^{2}_{2})P_{2}$$\equiv$ $\mathbf{p}_{2}^{2}$ (64)
where $\mathsf{t}^{1}_{1},\mathsf{t}^{1}_{2},\mathsf{t}^{2}_{1},\mathsf{t}^{2}_{2}\in[0,1]$. Consider applying calculus of variation in conjunction with power allocation 4-tuples $(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})$, as well as in conjunction with power allocation 4-tuples $(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})$. According to Theorem 6, the corresponding stationary solutions rely on zero-mean Gaussian code-books for core random variables $U_{1}$, $V_{1}$, $U_{2}$, $V_{2}$. Let us assume the two solutions result in the same point on the boundary, i.e.,
$\mathsf{W^{1}}(\mu)=\mathsf{W^{2}}(\mu)=\max(R_{1}+\mu R_{2})$ (65)
where superscripts 1,2 correspond to power allocations $(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})$ and $(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})$, respectively. Consider power allocation 4-tuples obtained by time sharing between $(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})$ and $(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})$, i.e.,
$\mathsf{T}(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})+(1-\mathsf{T})(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2}),\leavevmode\nobreak\ \mathsf{T}\in[0,1].$ (66)
Time-sharing between $1^{st}$ and $2^{nd}$ points result in the same value of $\mathsf{W^{1}}(\mu)=\mathsf{W^{2}}(\mu)$ for the weighted sum-rate. On the other hand, if
$(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})\neq(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})$ (67)
it follows that
$(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})\neq(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})\neq\mathsf{T}(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})+(1-\mathsf{T})(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})\leavevmode\nobreak\ \mbox{for}\leavevmode\nobreak\ \mathsf{T}\neq 0,1.$ (68)
For given $\mu$, let us apply calculus of variation in conjunction with power allocation $\mathsf{T}(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})+(1-\mathsf{T})(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})\leavevmode\nobreak\ \mbox{for}\leavevmode\nobreak\ \mathsf{T}\neq 0,1$. This results in a solution, using zero-mean Gaussian densities for core random variables, with a weighted sum-rate $\mathsf{\breve{W}}(\mu)$ larger than $\mathsf{W^{1}}(\mu)=\mathsf{W^{2}}(\mu)$. This contradicts the initial assumption, entailing $(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})$ and $(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})$, where $(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})\neq(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})$, cannot result in the same point on the boundary. ∎
Theorem 8 shows that Gaussian code-books result in a stationary solution for $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$.
Theorem 8.
Gaussian densities for $U_{1}$, $V_{1}$, $U_{2}$, $V_{2}$ result in a stationary solution for $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$.
Proof.
Appendix B establishes that Gaussian densities for composite random variables result in a stationary solution for $\mathbf{\Upsilon}$, as well as for $\mathbf{\Gamma}$. In the following, it is established that densities for core random variables will be Gaussian as well. Let us focus on $\mathbf{N}$, i.e.,
$\mathbf{N}\equiv R^{e}_{V_{2}}+R^{e}_{U_{2}}-R^{s}_{V_{2}}-R^{s}_{U_{2}}$ (69)
where $R^{s}_{V_{2}},R^{s}_{U_{2}}$ are fixed and $R^{e}_{V_{2}},R^{e}_{U_{2}}$ should be optimized. For $\mu<1$, one relies on successive decoding at $Y_{2}$ (see channel models in Fig. 4). This means $R_{U_{2}}^{e}$ and $R_{V_{2}}^{e}$, forming $\mathbf{N}$ in 69, are mutual information terms across two channels formed at $Y_{2}$, each with an additive noise independent of its input. Mutual information terms forming $R^{e}_{V_{2}}$ and $R^{e}_{U_{2}}$ are each composed of two entropy terms (likewise for $R^{e}_{V_{1}}$ and $R^{e}_{U_{1}}$ appearing in $\mathbf{D}$). For simplicity, formulations do not explicitly include the role of Gaussian noise terms added at $Y_{1}$ and $Y_{2}$. Let us use notations $\mathbf{p}_{i}$, $i=1,2,3,4$ to refer to densities of composite random variables appearing in entropy terms in $R^{e}_{V_{2}}$ and $R^{e}_{U_{2}}$. From Fig. 4, we have
$\mathbf{p}_{1}:$ density function of composite random variable $\!\!\sqrt{a}U_{1}+\sqrt{a}V_{1}+U_{2}+V_{2}$ (70)
$\mathbf{p}_{2}:$ density function of composite random variable $\!\!\sqrt{a}U_{1}+\sqrt{a}V_{1}+V_{2}$ (71)
$\mathbf{p}_{3}:$ density function of composite random variable $\!\!\sqrt{a}V_{1}+V_{2}$ (72)
$\mathbf{p}_{4}:$ density function of composite random variable $\!\!\sqrt{a}V_{1}.$ (73)
It is observed that each of the corresponding composite random variables is a linear combination of $U_{1}$, $U_{2}$, $V_{1}$, $V_{2}$. Likewise, in $\mathbf{D}$, term $R_{U_{1}}^{e}$ is governed by an additive noise channel formed at $Y_{2}$, and $R_{V_{1}}^{e}$ is governed by an additive noise channel formed at $Y_{1}$ (after $U_{1},U_{2}$ are jointly decoded). Relevant entropy terms include two additional composite random variables with densities $\mathbf{p}_{5}$, $\mathbf{p}_{6}$ where
$\mathbf{p}_{5}:$ density function of composite random variable $\!\!V_{1}+\sqrt{b}V_{2}$ (74)
$\mathbf{p}_{6}:$ density function of composite random variable $\!\!\sqrt{b}V_{2}.$ (75)
Since $U_{1}$, $U_{2}$, $V_{1}$, $V_{2}$ are independent of each other, each $\mathbf{p}_{i}$, $i=1,2,3,4,5,6$ can be expressed in terms of a convolution. In applying calculus of variation, densities are assumed to be zero mean, and constraints on “power" and “area under each density function" are added to the objective function using Lagrange multipliers. Then, the density functions of core random variables $U_{1}$, $U_{2}$, $V_{1}$, $V_{2}$ are perturbed using $\epsilon_{1}h_{1}$, $\epsilon_{2}h_{2}$, $\epsilon_{3}h_{3}$ and $\epsilon_{4}h_{4}$. Setting the derivatives of 6 with respect to $\epsilon_{i}$, $i=1,2,3,4$ equal to zero results in
$\frac{\partial\mathbf{\Upsilon}}{\partial\epsilon_{i}}\Big{|}_{\epsilon_{i}=0}=0\leavevmode\nobreak\ \leavevmode\nobreak\ \Longrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \left(\frac{\partial\mathbf{N}}{\partial\epsilon_{i}}\mathbf{D}-\frac{\partial\mathbf{D}}{\partial\epsilon_{i}}\mathbf{N}\right)\Big{|}_{\epsilon_{i}=0}=0.$ (76)
Constraints on powers of core random variables are
$P_{U_{1}}+P_{V_{1}}$ $=$ $P_{1}$ (77)
$P_{U_{2}}+P_{V_{2}}$ $=$ $P_{2}.$ (78)
Power constraints in 77, 78 are expressed in terms a larger set, with each constraint limiting the power of a composite random variable. Power constraints in this larger set are linearly dependent, causing redundancy. However, since constraints in the enlarged set are consistent, imposing redundancy does not affect the validity of the final solution. A similar set of redundant constraints are used in imposing the restriction that the area under each density function should be equal to one. Under these conditions, relying on a formulation similar to [41] (see page 335), it follows that
$\frac{\partial\mathbf{N}}{\partial\epsilon_{i}}\Big{|}_{\epsilon_{i}=0}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \frac{\partial\mathbf{D}}{\partial\epsilon_{i}}\Big{|}_{\epsilon_{i}=0}$ (79)
in 76 will be zero if densities of $U_{1}$, $U_{2}$, $V_{1}$, $V_{2}$ are zero-mean Gaussian. Appendix B includes some details in applying calculus of variation to $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$. It follows that the same Gaussian densities for core random variables which result in a stationary solution for $\mathbf{\Upsilon}$, also result in a stationary solution for $\mathbf{\Gamma}$. Relying on Theorem 5, such stationary solutions should be either the maximum or the minimum. ∎
Next, the condition for a power reallocation vector to be boundary achieving is discussed. Let us consider a step along the boundary which is small enough such that the coding strategy remains the same within the step. Let us assume $(\hat{\Delta}\!P_{1},\hat{\Delta}\!P_{2})$ is the power reallocation vector corresponding to an end point beyond which a change in strategy is needed, and consider
$(\Delta\!P_{1},\Delta\!P_{2}):\Delta\!P_{1}\leq\hat{\Delta}\!P_{1}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \Delta\!P_{2}\leq\hat{\Delta}\!P_{2}.$ (80)
Let us define $\upsilon\leq\mu^{s}$, where $\mu^{s}$ is the value of $\mu$ at the starting point, and the set $\bar{\mathsf{S}}_{\upsilon}$ as
$\bar{\mathsf{S}}_{\upsilon}=\left\{f_{U_{1}},f_{V_{1}},f_{U_{2}},f_{V_{2}}\!:\!\leavevmode\nobreak\ \mbox{outgoing slope at the starting point is}\leavevmode\nobreak\ \upsilon\triangleq\min_{(\delta P_{1},\delta P_{2})\in[0,\Delta\!P_{1}]\times[0,\Delta\!P_{2}]}\mathbf{\Upsilon}\right\}.$ (81)
Set $\bar{\mathsf{S}}_{\upsilon}$ is defined over all possible code-books’ densities, including Gaussian. Each member of 81 corresponds to a power reallocation vector $(\delta P_{1},\delta P_{2})\in[0,\Delta\!P_{1}]\times[0,\Delta\!P_{2}]$. This correspondence is potentially many-to-one since multiple choices for densities $(f_{U_{1}},f_{V_{1}},f_{U_{2}},f_{V_{2}})$, with the same $(\delta P_{1},\delta P_{2})$, may achieve the same $\mathbf{\Upsilon}=\upsilon$. Given $\mathbf{\Upsilon}=\upsilon$, the size of the set $\bar{\mathsf{S}}_{\upsilon}$ is reduced by limiting it to choice(s) which maximize $\mathbf{\Gamma}$. Maximum value of $\mathbf{\Gamma}$ over the set $\bar{\mathsf{S}}_{\upsilon}$ is denoted as $\varkappa_{\upsilon}$. Let us consider a second set $\bar{\bar{\mathsf{S}}}_{\upsilon}$ where
$\bar{\bar{\mathsf{S}}}_{\upsilon}\subseteq\bar{\mathsf{S}}_{\upsilon}:\leavevmode\nobreak\ \mathbf{\Gamma}=\varkappa_{\upsilon}.$ (82)
The set $\bar{\bar{\mathsf{S}}}_{\upsilon}$ includes a point on the boundary with
$\mathbf{\Upsilon}=\upsilon\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\Gamma}=\varkappa_{\upsilon}\triangleq\max_{\mathbf{\Upsilon}=\upsilon}\mathbf{\Gamma}.$ (83)
We are interested in establishing that the size of $\bar{\bar{\mathsf{S}}}_{\upsilon}$ can be reduced, by increasing $\upsilon$, such that the shrunken set includes a single element, say $\zeta$. Since $\bar{\bar{\mathsf{S}}}_{\upsilon}$ always includes a point on the boundary, it follows that $\zeta$ falls on the boundary. In addition, we need to show that: (i) $\zeta$ is realized using Gaussian code-books, and (ii) the rest of the boundary can be covered starting from $\zeta$. Theorem 9 addresses these requirements.
Theorem 9.
Cardinality of the set $\bar{\bar{\mathsf{S}}}_{\upsilon}$ can be reduced, by increasing $\upsilon<\mu^{s}$, in a recursive manner, such that the final set is associated with a single $(\delta P_{1},\delta P_{2})$.
Proof.
Let us assume the original set $\bar{\bar{\mathsf{S}}}_{\upsilon}$ is associated with $M$ distinct vectors $(\delta^{m}P_{1},\delta^{m}P_{2})$, $m=1,\ldots,M$. Each of these $M$ vectors is associated with a respective set of code-books’ densities. Consider
$(\breve{\delta}P_{1},\breve{\delta}P_{2})=(\min_{m}\delta^{m}P_{1},\min_{m}\delta^{m}P_{2}).$ (84)
The pair $(\breve{\delta}P_{1},\breve{\delta}P_{2})$ is called the Pareto minimal point corresponding to the set $(\delta^{m}P_{1},\delta^{m}P_{2})$, $m=1,\ldots,M$. Let us use $(\breve{\delta}P_{1},\breve{\delta}P_{2})$ to compute new values for $(\mathbf{\Upsilon},\mathbf{\Gamma})$ and select the subset with smallest value of $\upsilon$ denoted as $\breve{\upsilon}$. Accordingly, let us form the sets $\bar{\mathsf{S}}_{\breve{\upsilon}}$ and $\bar{\bar{\mathsf{S}}}_{\breve{\upsilon}}$. Starting from the power reallocation vector $(\breve{\delta}P_{1},\breve{\delta}P_{2})$, each of the pairs $(\delta^{m}P_{1},\delta^{m}P_{2})$, $m=1,\ldots,M$, can be reached relying on a step with power reallocation $(\delta^{m}P_{1}-\breve{\delta}P_{1},\delta^{m}P_{2}-\breve{\delta}P_{2})$. This is possible since $\delta^{m}P_{1}-\breve{\delta}P_{1}\geq 0$ and $\delta^{m}P_{2}-\breve{\delta}P_{2}\geq 0$. This means relying on $\bar{\bar{\mathsf{S}}}_{\breve{\upsilon}}$ to achieve the next point on the boundary does not contradict the possibility of further moving counterclockwise to achieve the boundary point corresponding to $\bar{\bar{\mathsf{S}}}_{\upsilon},\leavevmode\nobreak\ \upsilon<\breve{\upsilon}$. Now let us shrink the range for power reallocation vector by setting
${\Delta}P_{1}=\breve{\delta}P_{1}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ {\Delta}P_{2}=\breve{\delta}P_{2}.$ (85)
Accordingly, let us construct new sets following 81 and 82. Having multiple elements in $\bar{\bar{\mathsf{S}}}_{\breve{\upsilon}}$ allows recursively moving in clockwise direction, where $\mathbf{\Upsilon}$ increases and $\mathbf{\Gamma}$ decreases in each step. This procure can continue until one of the following cases occurs. Case (i): The value of $\mathbf{\Gamma}$ at the final point is zero. Case (ii): The final set includes a single Pareto minimal power reallocation vector achieving a single point on the boundary. Case (i) entails no further counterclockwise step along the boundary is feasible, requiring a change in the strategy. In Case (ii), from theorems 1, 8 and 9, it follows that there is a Pareto minimal power reallocation which, in conjunction with zero-mean Gaussian code-books for composite random variables, results in a unique point on the boundary. ∎
In summary, referring to Theorem 9, using $(\breve{\delta}P_{1},\breve{\delta}P_{2})$ instead of $(\delta^{m}P_{1},\delta^{m}P_{2})$, $m=1,\ldots,M$ is accompanied by a movement in clockwise direction, i.e., reaching from $(\upsilon,\gamma)$ to $(\breve{\upsilon},\breve{\gamma})$, where
$(\upsilon,\varkappa_{\upsilon})\leadsto(\breve{\upsilon},\varkappa_{\breve{\upsilon}}):\leavevmode\nobreak\ \leavevmode\nobreak\ \breve{\upsilon}>\upsilon\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \varkappa_{\breve{\upsilon}}<\varkappa_{\upsilon}.$ (86)
Such a movement can continue in a recursive manner until the step size is small enough to include a single power reallocation vector, i.e.,
$\exists i\in[1,\ldots,M]:\leavevmode\nobreak\ (\breve{\delta}P_{1},\breve{\delta}P_{2})=(\delta^{i}P_{1},\delta^{i}P_{2})$ (87)
with the resulting $(\breve{\delta}P_{1},\breve{\delta}P_{2})$ achieving to a unique point on the boundary. Theorem 9 entails, relying on Pareto minimal power reallocation, the past history in moving counterclockwise along the boundary is captured solely by the starting point in each step. This means, considering two nested Pareto minimal power reallocation vectors $(\dot{\delta}P_{1},\dot{\delta}P_{2})$ and $(\ddot{\delta}P_{1},\ddot{\delta}P_{2})$, where
$\dot{\delta}P_{1}\leq\ddot{\delta}P_{1}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \dot{\delta}P_{2}\leq\ddot{\delta}P_{2}.$ (88)
These power reallocation vectors, in conjunction with Gaussian code-books, achieve two successive points on the boundary
$(\mathbf{\Upsilon}_{\!1},\mathbf{\Gamma}_{\!1})=(\dot{\upsilon},\varkappa_{\dot{\upsilon}})\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ (\mathbf{\Upsilon}_{\!2},\mathbf{\Gamma}_{\!2})=(\ddot{\upsilon},\varkappa_{\ddot{\upsilon}})$ (89)
satisfying
$\ddot{\upsilon}\leq\dot{\upsilon}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \varkappa_{\ddot{\upsilon}}\geq\varkappa_{\dot{\upsilon}}.$ (90)
It remains to show that Gaussian densities for composite random variables entail that core random variables will be Gaussian as well. This is established in Theorem 10.
Theorem 10.
Assume power reallocation vector is Pareto minimal. Then, the stationary solution obtained using Gaussian densities for core random variables results in an end-point which falls on the boundary.
Proof.
Consider a power reallocation vector achieving a unique end point on the boundary. For such a power reallocation vector, consider applying calculus of variation to $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$ by perturbing densities of core random variables. Setting the derivatives of underlying functionals to zero results in a system of equality constraints, which are satisfied if composite random variables are jointly Gaussian. Each composite random variable is a linear combination of core random variables, and linear expressions obtained using different sets of composite random variables are consistent with each other. Theorem 4 shows that the matrix of corresponding linear coefficients is invertible. This in turn means core random variables can be expressed as a unique linear combination of composite random variables. This means core random variables should be Gaussian, and the correspondence is unique. From theorems 5, 8, the stationary solution based on Gaussian densities for composite random variables either maximizes or minimizes $\mathbf{\Upsilon}$. A similar conclusion applies to $\mathbf{\Gamma}$. Combining these arguments with the result of Theorem 1, it is concluded that for the Gaussian code-books in conjunction with a Pareto minimal power reallocation vector, $\mathbf{\Upsilon}$ is minimized while $\mathbf{\Gamma}$ is maximized. ∎
Figure 6: $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$ as a function of time sharing factor $\omega$ (related to Theorem 11).
Remark 4: Note that optimum Pareto minimal power reallocation vector is not unique. However, the corresponding set has a nested structure, and relying on any element of the set will be associated with a unique set of Gaussian code-books (see Theorem 10), achieving a point on the boundary. Different elements in the set of Pareto minimal power reallocation pairs correspond to different step sizes. This property allows covering the boundary in a continuous manner. Theorem 11 shows that conclusions relying on the concept of Pareto minimal power reallocation can be also reached by linearly changing the power reallocation vector to cover a segment on the boundary. $\blacksquare$
Next, Theorem 11, in conjunction with Fig. 6, establishes that, given power reallocation vector $\omega({\delta}P^{\mathbb{E}}_{U_{1}},{\delta}P^{\mathbb{E}}_{V_{1}},{\delta}P^{\mathbb{E}}_{U_{2}},{\delta}P^{\mathbb{E}}_{V_{2}})$, Gaussian code-books minimize $\mathbf{\Upsilon}^{\mathbb{B}}(\omega)$ and maximize $\mathbf{\Gamma}^{\mathbb{B}}(\omega)$. This results in a unique point on the boundary. With some misuse of notation, superscripts are used to refer to points inside or on the capacity region. Consider a segment on the boundary from a starting point $\mathbb{S}$ to an end point $\mathbb{E}$ as depicted in Fig. 6. Assume the power reallocation vector for point $\mathbb{E}$ is equal to $({\delta}P^{\mathbb{E}}_{U_{1}},{\delta}P^{\mathbb{E}}_{V_{1}},{\delta}P^{\mathbb{E}}_{U_{2}},{\delta}P^{\mathbb{E}}_{V_{2}})$. Consider time sharing between points $\mathbb{S}$ and $\mathbb{E}$ with a time sharing factor $\omega\in[0,1]$ where $\omega=0$ and $\omega=1$ correspond to points $\mathbb{S}$ and $\mathbb{E}$, respectively. Time sharing achieves point $\mathbb{T}$ inside the capacity region corresponding to a power reallocation vector $\omega({\delta}P^{\mathbb{E}}_{U_{1}},{\delta}P^{\mathbb{E}}_{V_{1}},{\delta}P^{\mathbb{E}}_{U_{2}},{\delta}P^{\mathbb{E}}_{V_{2}})$. Let us assume the power reallocation vector $\omega({\delta}P^{\mathbb{E}}_{U_{1}},{\delta}^{\mathbb{E}}P_{V_{1}},{\delta}P^{\mathbb{E}}_{U_{2}},{\delta}P^{\mathbb{E}}_{V_{2}})$, with optimum codebooks’ densities, results in the point $\mathbb{B}$ on the boudnary corresponding to $(\mathbf{\Upsilon}^{\mathbb{B}},\mathbf{\Gamma}^{\mathbb{B}})$. This means $\mathbf{\Upsilon}^{\mathbb{B}}$ and $\mathbf{\Gamma}^{\mathbb{B}}$ are both unique functions of $\omega$, denoted as $\mathbf{\Upsilon}^{\mathbb{B}}(\omega)$ and $\mathbf{\Gamma}^{\mathbb{B}}(\omega)$, respectively. Relying on codebooks’ densities obtained through time sharing for point $\mathbb{T}$ and optimum codebooks’ densities for points $\mathbb{E}$ and $\mathbb{B}$, we have
$\mathbf{\Upsilon}^{\mathbb{B}}$ $>$ $\mathbf{\Upsilon}^{\mathbb{T}}\leavevmode\nobreak\ =\leavevmode\nobreak\ \mathbf{\Upsilon}^{\mathbb{E}}$ (91)
$\mathbf{\Gamma}^{\mathbb{B}}$ $>$ $\mathbf{\Gamma}^{\mathbb{T}}\leavevmode\nobreak\ =\leavevmode\nobreak\ \omega\mathbf{\Gamma}^{\mathbb{E}}.$ (92)
Theorem 11.
As functions of $\omega$, $\mathbf{\Upsilon}^{\mathbb{B}}(\omega),\,\omega\in[0,1]$ is monotonically decreasing and $\mathbf{\Gamma}^{\mathbb{B}}(\omega),\,\omega\in[0,1]$ is monotonically increasing.
Proof.
If $\mathbf{\Upsilon}^{\mathbb{B}}(\hat{\omega})$ increases for $\hat{\omega}>\omega$, time sharing coefficient $\hat{\omega}$ would result in a new point on the boundary prior to point $\mathbb{B}$, and a point on the time sharing line prior to point $\mathbb{T}$. This procure can be repeated until one of the following two cases occur: Case (i) the new points move counterclockwise, i.e., direction of movement is reversed. Case (ii) new points fall on $\mathbb{S}$. Case (i) cannot occur since it entails there are two overlapping points on the time sharing line which correspond to two different values of time sharing coefficient. Case (ii) contradicts the basic assumption that, starting from point $\mathbb{S}$, counterclockwise movement along the boundary is feasible. Case (ii) occurs if the starting point $\mathbb{S}$ overlaps with the end point $\mathbb{E}$, requiring a change in the strategy. ∎
Appendix
In the following, to simplify expressions, entropy values are computed in base “e".
Appendix A Constrained Maximization of Entropy Functions
A.1 Entropy Term Involving a Single Density Function
Consider the following constrained optimization problem:
$\mbox{Find function}\leavevmode\nobreak\ \leavevmode\nobreak\ f>0\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{to maximize}$ $\displaystyle-\int f\log(f)$ (93)
subject to: $\displaystyle\int x^{2}f=P$ (94)
and $\displaystyle\int f=1.$ (95)
Using Lagrange multipliers to add 94 and 95 to 93, we obtain
$-\int f\log(f)+\lambda\int x^{2}f+\gamma\int f.$ (96)
Using a perturbation term $\epsilon h$ in 96 results in
$-\int(f+\epsilon h)\log(f+\epsilon h)+\lambda\int x^{2}(f+\epsilon h)+\gamma\int f+\epsilon h.$ (97)
Derivative of 97 with respect to $\epsilon$ is equal to
$-\int h\left[\log(f+\epsilon h)+1-\lambda x^{2}-\gamma\right].$ (98)
Setting 98 to zero for $\epsilon=0$, it follows that a Gaussian density for $f$ results in a stationary solution for constrained optimization problem in 93, 94 and 95. Next, it is shown that such a stationary solution is the maximum by using second order perturbation. Derivative of 98 with respect to $\epsilon$ at $\epsilon=0$ is equal to
$-\int\frac{h^{2}}{f+\epsilon h}\Big{|}_{\epsilon=0}=-\int\frac{h^{2}}{f}\,<\,0\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{for}\leavevmode\nobreak\ \leavevmode\nobreak\ h\neq 0\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{since}\leavevmode\nobreak\ \leavevmode\nobreak\ h^{2}>0\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ f\geq 0.$ (99)
Referring to reference [1], the condition in 99 implies that Gaussian density for $f$, computed relying on calculus of variation, is the global maximum solution for optimization problem in 93, 94 and 95.
A.2 Entropy Term Involving a Convolution of Density Functions
Let us consider functional $\digamma$ defined as
ϝ$\digamma=f_{1}\ast f_{2}.$ (100)
Entropy of $\digamma$ is
$\mathsf{H}^{\digamma}=-\int{\digamma}\ln({\digamma}).$ (101)
Perpetuation of $\digamma$, denoted a $p\digamma$, is equal to
$p\digamma=(f_{1}+\epsilon_{1}h_{1})\ast(f_{2}+\epsilon_{2}h_{2})$ (102)
with an entropy of
$\mathsf{H}^{p\digamma}=-\int(f_{1}+\epsilon_{1}h_{1})\ast(f_{2}+\epsilon_{2}h_{2})\ln[(f_{1}+\epsilon_{1}h_{1})\ast(f_{2}+\epsilon_{2}h_{2})].$ (103)
To have a stationary solution for $\digamma$, density functions $f_{1}$ and $f_{2}$ should satisfy
$\frac{\partial\mathsf{H}^{p\digamma}}{\partial\epsilon_{1}}\Big{|}_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ 00 0 (104)
$\frac{\partial\mathsf{H}^{p\digamma}}{\partial\epsilon_{2}}\Big{|}_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ 00 0 (105)
$\frac{\partial\mathsf{H}^{p\digamma}}{\partial\epsilon_{1}}\Big{|}_{\epsilon_{1}=0,\epsilon_{2}=0}=-\int(h_{1}\ast f_{2})\ln(f_{1}\ast f_{2})-\int(h_{1}\ast f_{2})$ $=$ (106)
$-\int(h_{1}\ast f_{2})[\ln(f_{1}\ast f_{2})+1].$ (107)
Constraints on power and probability density function are expressed as:
[…]
We have
$\frac{\partial\mathsf{E}^{f_{1}\ast f_{2}}}{\partial\epsilon_{1}}\Big{|}_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $\int x^{2}(h_{1}\ast f_{2})$ (110)
$\frac{\partial\mathsf{A}^{f_{1}\ast f_{2}}}{\partial\epsilon_{1}}\Big{|}_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $\int(h_{1}\ast f_{2}).$ (111)
Adding 110 and 111 with Lagrange multipliers $\lambda_{1}$ and $\lambda_{2}$ to 107, we obtain
$-\int(h_{1}\ast f_{2})[\ln(f_{1}\ast f_{2})+1-\lambda_{1}x^{2}-\lambda_{2}].$ (112)
Similarly, for derivative with respect to $\epsilon_{2}$, we obtain
$-\int(f_{1}\ast h_{2})[\ln(f_{1}\ast f_{2})+1-\lambda_{3}x^{2}-\lambda_{4}].$ (113)
Setting 112 and 113 to zero, it follows that a Gaussian density for $f_{1}(x_{1})\ast f_{2}(x_{2})$ is a stationary point for the entropy of ϝ$\digamma=f_{1}(x_{1})\ast f_{2}(x_{2})$. Derivation is very similar to [41] (see page 335). The final conclusion is that $x_{1}+x_{2}$ is Gaussian. However, having a Gaussian density for $x_{1}+x_{2}$ does not mean $x_{1}$ and $x_{2}$ should be Gaussian as well. This problem does not occur in the case of interest here, since, having a Gaussian density for composite random variables can occur only if core random variables are Gaussian. This point is established in Theorem 4.
A.3 Effect of Scaling of Random Variables
Let us consider
$\mathsf{H}^{\digamma_{\!\!1}}+\mathsf{H}^{\digamma_{\!\!2}}$ (114)
with
ϝ$\digamma_{\!\!1}$ $=$ $f_{1}(x)\ast\frac{1}{\gamma}f_{2}\left(\frac{x}{\gamma}\right)\ast n$ (115)
ϝ$\digamma_{\!\!2}$ $=$ $f_{2}(x)\ast n$ (116)
where $f_{1}$ and $f_{2}$ are densities of $x_{1}$ and $x_{2}$, respectively, and $n$ is Gaussian. Let us consider perturbing $f_{2}$ with $\epsilon_{2}h_{2}(x)$. We have
$f_{2}(x)\ast n$$\Longrightarrow$ $\left[f_{2}(x)+\epsilon_{2}h_{2}(x)\right]\ast n$ (117)
$f_{1}(x)\ast\frac{1}{\gamma}f_{2}\left(\frac{x}{\gamma}\right)\ast n$$\Longrightarrow$ $f_{1}(x)\ast\frac{1}{\gamma}\left[f_{2}\left(\frac{x}{\gamma}\right)+\epsilon_{2}h_{2}\left(\frac{x}{\gamma}\right)\right]\ast n.$ (118)
It turns out that the effect of $n$ does not impact conclusions (see Appendix A.4). For simplicity of notation, $n$ is ignored in the following derivations. As a result, 117 and 118 are simplified to
$f_{2}(x)$$\Longrightarrow$ $f_{2}(x)+\epsilon_{2}h_{2}(x)$ (119)
$\frac{1}{\gamma}f_{1}(x)\ast f_{2}\left(\frac{x}{\gamma}\right)$$\Longrightarrow$ $\frac{1}{\gamma}f_{1}(x)\ast f_{2}\left(\frac{x}{\gamma}\right)+\frac{\epsilon_{2}}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right).$ (120)
Corresponding entropy terms are:
$-\int[f_{2}(x)+\epsilon_{2}h_{2}(x)]\ln[f_{2}(x)+\epsilon_{2}h_{2}(x)]\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}$ (121)
$-\int\left[\frac{1}{\gamma}f_{1}(x)\ast f_{2}\left(\frac{x}{\gamma}\right)+\frac{\epsilon_{2}}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right]\ln\left[\frac{1}{\gamma}f_{1}(x)\ast f_{2}\left(\frac{x}{\gamma}\right)+\frac{\epsilon_{2}}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right].$ (122)
A.3.1 First Order Variations
Derivatives of 121, 122 with respect to $\epsilon_{2}$ are, respectively, equal to
$-\int h_{2}(x)\ln\left[f_{2}(x)+\epsilon_{2}h_{2}(x)\right]+h_{2}(x)\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}$ (123)
$-\int\left[\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right]\ln\left(\frac{1}{\gamma}f_{1}(x)\ast\left[f_{2}\left(\frac{x}{\gamma}\right)+\epsilon_{2}h_{2}\left(\frac{x}{\gamma}\right)\right]\right)+\left[\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right].$ (124)
Setting $\epsilon_{2}=0$ in 123, 124, we obtain
$-\int h_{2}(x)\ln f_{2}(x)+h_{2}(x)\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}$ (125)
$-\int\left(\left[\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right]\ln\left[\frac{1}{\gamma}f_{1}(x)\ast f_{2}\left(\frac{x}{\gamma}\right)\right]+\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right).$ (126)
Corresponding constraints on power are expressed as
$\int x^{2}f_{2}(x)$$\Longrightarrow$ $\int x^{2}[f_{2}(x)+\epsilon_{2}h_{2}(x)]\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}$ (127)
$\int x^{2}\left[f_{1}(x)\ast\frac{1}{\gamma}f_{2}\left(\frac{x}{\gamma}\right)\right]$$\Longrightarrow$ $\int x^{2}\left[f_{1}(x)\ast\frac{1}{\gamma}\left(f_{2}\left(\frac{x}{\gamma}\right)+\epsilon_{2}h_{2}\left(\frac{x}{\gamma}\right)\right)\right].$ (128)
Likewise, constraints on areas under density functions are expressed as
$\int f_{2}(x)$$\Longrightarrow$ $\int f_{2}(x)+\epsilon_{2}h_{2}(x)\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}$ (129)
$\int\left[f_{1}(x)\ast\frac{1}{\gamma}f_{2}\left(\frac{x}{\gamma}\right)\right]$$\Longrightarrow$ $\int\left[f_{1}(x)\ast\frac{1}{\gamma}\left(f_{2}\left(\frac{x}{\gamma}\right)+\epsilon_{2}h_{2}\left(\frac{x}{\gamma}\right)\right)\right].$ (130)
Computing derivatives of 127 and 128 with respect to $\epsilon_{2}$ and setting $\epsilon_{2}=0$ in the results, we obtain
$\int x^{2}h_{2}(x)\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}$ (131)
$\int x^{2}\left[\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right].$ (132)
Similar to 131 and 132, constraints on areas under density functions result in
$\int h_{2}(x)\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}$ (133)
$\int\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right).$ (134)
Then, using Lagrange multipliers, 131, 133 are added to 125 and 132, 134 to 126. Note that the term $h_{2}(x)$ is common in 125, 131 and 133 and can be factored out. Likewise, the term $\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)$ is common in 126, 132 and 134 and can be factored out. It follows that relying on Gaussian densities with proper variances for $f_{1}$ and $f_{2}$ results in a stationary point for the entropy terms in 121 and 122.
A.3.2 Second Order Variations
Noting 123 and 124, it follows that the second order derivative of 114 with respect to $\epsilon_{2}$, at $\epsilon_{2}=0$, is equal to
$-\,\,\frac{\left[f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right]^{2}}{\gamma f_{1}(x)\ast f_{2}\left(\frac{x}{\gamma}\right)}-\frac{\left[h_{2}(x)\right]^{2}}{f_{2}(x)}<0\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{since}\leavevmode\nobreak\ \leavevmode\nobreak\ h_{2}\neq 0.$ (135)
As a result, Gaussian density function (computed relying on calculus of variation) maximizes 114.
As will be discussed in Appendix B, for the objective function $\mathbf{\Upsilon}$ defined in 6, perturbations are formed using functions $\epsilon_{i}h_{i}$. Each second order derivative of the form
$\frac{\partial^{2}\mathbf{\Upsilon}}{\partial\epsilon_{i}^{2}}$ (136)
is composed of multiple terms, each of the form given in 135. The term corresponding to perturbation $\epsilon_{i}h_{i}$ will be zero only if $h_{i}=0$. This means collection of Gaussian density functions for composite random variables, each obtained from
$\frac{\partial\mathbf{\Upsilon}}{\partial\epsilon_{i}}=0\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{at}\leavevmode\nobreak\ \leavevmode\nobreak\ \epsilon_{i}=0$ (137)
result in a non-zero value for 136. This means the corresponding stationary solution is either a minimum or a maximum.
A.4 Functional of Composite Random Variables
Let us assume $f_{1}(x_{1})$ and $f_{2}(x_{2})$ are density functions for two core random variables, forming composite random variables $x_{1}+x_{2}$ and $x_{2}$. Let us define
ϝ$\digamma_{\!\!1}$ $=$ $f_{1}\ast f_{2}\ast n$ (138)
ϝ$\digamma_{\!\!2}$ $=$ $f_{2}\ast n$ (139)
where $n$ is the probability density function of the additive Gaussian noise. Then,
$\mathsf{H}^{\digamma_{\!\!1}}-\mathsf{H}^{\digamma_{\!\!2}}$ (140)
is the mutual information over an additive noise channel where $f_{1}\ast f_{2}\ast n$ is the channel output, and $f_{2}\ast n$ is the additive noise. We are interested to find a stationary solution for 140. In the following, variations of ϝ$\digamma_{\!\!1}$, ϝ$\digamma_{\!\!2}$ are denoted as $p\digamma_{\!\!1}$, $p\digamma_{\!\!2}$, respectively. Constrains on power are expressed as:
$\mathsf{E}^{f_{1}\ast f_{2}\ast n}$ $=$ $\int x^{2}(f_{1}\ast f_{2}\ast n)\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{is a constant}$ (141)
$\mathsf{E}^{f_{2}\ast n}$ $=$ $\int x^{2}(f_{2}\ast n)\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{is a constant}$ (142)
$\mathsf{A}^{f_{1}\ast f_{2}\ast n}$ $=$ $\int(f_{1}\ast f_{2}\ast n)=1.$ (143)
$\mathsf{A}^{f_{2}\ast n}$ $=$ $\int(f_{2}\ast n)=1.$ (144)
We have
$\frac{\partial\mathsf{H}^{p\digamma_{\!\!1}}}{\partial\epsilon_{1}}\Big{|}_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $-\int(h_{1}\ast f_{2}\ast n)[\ln(f_{1}\ast f_{2}\ast n)+1]$ (145)
$\frac{\partial\mathsf{E}^{p\digamma_{\!\!1}}}{\partial\epsilon_{1}}\Big{|}_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $\int x^{2}(h_{1}\ast f_{2}\ast n)$ (146)
$\frac{\partial\mathsf{A}^{p\digamma_{\!\!1}}}{\partial\epsilon_{1}}\Big{|}_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $\int h_{1}\ast f_{2}\ast n$ (147)
$\frac{\partial\mathsf{H}^{p\digamma_{\!\!1}}}{\partial\epsilon_{2}}\Big{|}_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $-\int(f_{1}\ast h_{2}\ast n)[\ln(f_{1}\ast f_{2}\ast n)+1]$ (148)
$\frac{\partial\mathsf{E}^{p\digamma_{\!\!1}}}{\partial\epsilon_{2}}\Big{|}_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $\int x^{2}(f_{1}\ast h_{2}\ast n)$ (149)
$\frac{\partial\mathsf{A}^{p\digamma_{\!\!1}}}{\partial\epsilon_{2}}\Big{|}_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $\int f_{1}\ast h_{2}\ast n$ (150)
and
$\frac{\partial\mathsf{H}^{p\digamma_{\!\!2}}}{\partial\epsilon_{2}}\Big{|}_{\epsilon_{2}=0}$ $=$ $-\int(h_{2}\ast n)\ln(f_{2}\ast n)$ (151)
$\frac{\partial\mathsf{E}^{p\digamma_{\!\!2}}}{\partial\epsilon_{2}}\Big{|}_{\epsilon_{2}=0}$ $=$ $\int x^{2}(h_{2}\ast n)$ (152)
$\frac{\partial\mathsf{A}^{p\digamma_{\!\!2}}}{\partial\epsilon_{2}}\Big{|}_{\epsilon_{2}=0}$ $=$ $\int h_{2}\ast n.$ (153)
Adding 146, 147 with Lagrange multipliers to 145; 149, 150 with Lagrange multipliers to 148; and 152, 153 with Lagrange multipliers to 151, then setting the results to zero, it follows that Gaussian distributions for $f_{1}\ast f_{2}$ and $f_{2}$ result in a stationary solution for 140. Denoting arguments of $f_{1}$, $f_{2}$, $n$ as $x_{1}$, $x_{2}$, $z$, respectively, this entails random variables $y_{1}=x_{1}+x_{2}+z$ and $y_{2}=x_{2}+z$ are jointly Gaussian, and consequently, $w_{1}=x_{1}+x_{2}$ and $w_{2}=x_{2}$ are jointly Gaussian as well. In general, if a linear combination of some random variables is Gaussian, it does not necessarily mean each random variable should be Gaussian as well. However, in this example, we can uniquely express $x_{1}$, $x_{2}$ in terms of $w_{1}$, $w_{2}$, i.e., $x_{1}=w_{1}-w_{2}$ and $x_{2}=w_{2}$. This entails $x_{1}$ and $x_{2}$ should be Gaussian as well.
Remark 5: Let use rely on indices $\{1,\ldots,\mathbf{c}_{1}\}$, $\{1,\ldots,\mathbf{c}_{2}\}$ to specify elements of core and composite random variables, respectively. In this work, $\mathbf{c}_{2}\geq\mathbf{c}_{1}$. To obtain a stationary solution, it is shown that composite random variables should be jointly Gaussian. In addition, there are subset(s) of $\{1,\ldots,\mathbf{c}_{2}\}$ of size $\mathbf{c}_{1}$ such that corresponding matrix of linear coefficients is full rank. This property allows expressing core random variables as a linear combination of a subset of composite random variables. Also, expressions corresponding to different subsets of size $\mathbf{c}_{1}$ from $\{1,\ldots,\mathbf{c}_{2}\}$ are consistent. The conclusion is that each core random variable should be Gaussian. $\blacksquare$
Appendix B Stationary Solutions for $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$
This appendix shows that density functions resulting in stationary solutions for $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$ are Gaussian. Let us focus on $\mathbf{\Upsilon}$, since the derivation for $\mathbf{\Gamma}$ is very similar. Terms forming the numerator and the denominator of $\mathbf{\Upsilon}$ are rate across channels with additive noise (see Fig. 4). Each entropy term, corresponding to a composite random variable, is based on the convolution of densities of the underlying core random variables. As will be shown in Appendix A.3, scale factors for core random variables do not affect the derivations to follow. For this reason, such scale factors are not included in this Appendix. Notation $\circledast_{i=1}^{\mathfrak{q}}\mathfrak{p}_{i}$ denotes the convolution $\mathfrak{p}_{1}\ast\mathfrak{p}_{2}\ast...\ast\mathfrak{p}_{q}$, called multi-convolution hereafter. Recall that calculus of variation is applied by perturbing density function of each core random variable. For this reason, multi-convolution terms which involve same core random variables appear in the derivations. Since derivatives related to perturbation of different core random variables are handled separately, one can limit derivations to multi-convolution terms which have (at least) one common core random variable, denoted by the generic notation $\mathfrak{g}$ hereafter. If multi-convolution terms include two or more common core random variables, say $\mathfrak{g}_{1}$ and $\mathfrak{g}_{2}$, since such common terms are perturbed separately, derivations for each term will be similar to what is presented here. It is also enough to consider only four entropy terms (to define a reduced/generic expression for $\mathbf{\Upsilon}$) as given in 154. Derivations for more general cases will be similar (due to linearity of multi-convolution with respect to its terms).
$\frac{\mathfrak{N}}{\mathfrak{D}}=\frac{\displaystyle-\int\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}\right)+\int\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}\right)}{\displaystyle-\int\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}\right)+\int\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}\right)}.$ (154)
Let us perturb $\mathfrak{g}\Rightarrow\mathfrak{g}+\ell\mathfrak{h}$ resulting in
[…]
Derivatives of right hand terms in 155, with respect to $\ell$, are equal to
 $\textbf{\ T}_{1}(\ell)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\partial}{\partial\ell}\int\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{h}\right)\log\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{h}\right)$ $=$
$\int\circledast_{i=1}^{\mathfrak{m}}(\mathfrak{a}_{i}\ast\mathfrak{h})\log\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{h}\right)+\circledast_{i=1}^{\mathfrak{m}}(\mathfrak{a}_{i}\ast\mathfrak{h})$ (156)
 $\textbf{\ T}_{2}(\ell)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\partial}{\partial\ell}\int\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{h}\right)\log\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{h}\right)$ $=$
$\int\circledast_{i=1}^{\mathfrak{n}}(\mathfrak{b}_{i}\ast\mathfrak{h})\log\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{h}\right)+\circledast_{i=1}^{\mathfrak{n}}(\mathfrak{b}_{i}\ast\mathfrak{h})$ (157)
 $\textbf{\ T}_{3}(\ell)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\partial}{\partial\ell}\int\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{h}\right)\log\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{h}\right)$ $=$
$\int\circledast_{i=1}^{\mathfrak{p}}(\mathfrak{e}_{i}\ast\mathfrak{h})\log\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{h}\right)+\circledast_{i=1}^{\mathfrak{p}}(\mathfrak{e}_{i}\ast\mathfrak{h})$ (158)
 $\textbf{\ T}_{4}(\ell)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\partial}{\partial\ell}\int\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{h}\right)\log\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{h}\right)$ $=$
$\int\circledast_{i=1}^{\mathfrak{q}}(\mathfrak{f}_{i}\ast\mathfrak{h})\log\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{h}\right)+\circledast_{i=1}^{\mathfrak{q}}(\mathfrak{f}_{i}\ast\mathfrak{h}).$ (159)
It follows that
 T1(0)subscript $\textbf{\ T}_{1}(0)$ $=$ $\int(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{h})[1+\log(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g})]$ (160)
 T2(0)subscript $\textbf{\ T}_{2}(0)$ $=$ $\int(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{h})[1+\log(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g})]$ (161)
 T3(0)subscript $\textbf{\ T}_{3}(0)$ $=$ $\int(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{h})[1+\log(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g})]$ (162)
 T4(0)subscript $\textbf{\ T}_{4}(0)$ $=$ $\int(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{h})[1+\log(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g})].$ (163)
Noting the expression for $\mathfrak{D}$ in 154, let us define
 D1subscript $\textbf{\ D}_{1}$ $=$ $-\int\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}\right)$ (164)
 D2subscript $\textbf{\ D}_{2}$ $=$ $\int\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}\right)$ (165)
$\mathfrak{D}$ $=$  $\textbf{\ D}_{1}+\textbf{\ D}_{2}.$ (166)
It follows that
$\frac{\partial}{\partial\ell}\frac{\partial\mathfrak{N}}{\partial\mathfrak{D}}\mid_{\ell=0}=\frac{\displaystyle\frac{\partial\mathfrak{N}}{\partial\ell}\mathfrak{D}-\frac{\partial\mathfrak{D}}{\partial\ell}\mathfrak{N}}{\mathfrak{D}^{2}}\mid_{\ell=0}=\frac{-\!\!\textbf{\ T}_{1}(0)+\!\!\textbf{\ T}_{2}(0)}{\!\!\textbf{\ D}_{1}+\!\!\textbf{\ D}_{2}}-\frac{-\!\!\textbf{\ T}_{3}(0)+\!\!\textbf{\ T}_{4}(0)}{(\!\!\textbf{\ D}_{1}+\!\!\textbf{\ D}_{2})^{2}}=$ (167)
[…]
where
[…]
Constraints on power corresponding to terms  T1()subscript $\!\!\textbf{\ T}_{1}(\ell)$,  T2()subscript $\!\!\textbf{\ T}_{2}(\ell)$,  T3()subscript $\!\!\textbf{\ T}_{3}(\ell)$ and  T4()subscript $\!\!\textbf{\ T}_{4}(\ell)$ are
$\int x^{2}(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g})$$\Rightarrow$ $\int x^{2}(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{h})$ (172)
$\int x^{2}(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g})$$\Rightarrow$ $\int x^{2}(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{h})$ (173)
$\int x^{2}(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g})$$\Rightarrow$ $\int x^{2}(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{h})$ (174)
$\int x^{2}(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g})$$\Rightarrow$ $\int x^{2}(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{h}).$ (175)
Computing the derivatives of above terms with respect to $\ell$ for $\ell=0$, and including Lagrange multipliers $\varsigma_{1}$, $\varsigma_{2}$, $\varsigma_{3}$ and $\varsigma_{4}$, we obtain
$\varsigma_{1}\int x^{2}(\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i})$ (176)
$\varsigma_{2}\int x^{2}(\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i})$ (177)
$\varsigma_{3}\int x^{2}(\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i})$ (178)
$\varsigma_{4}\int x^{2}(\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}).$ (179)
Likewise, constraints on areas under density functions can be expressed as
$\iota_{1}\int\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}$ (180)
$\iota_{2}\int\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}$ (181)
$\iota_{3}\int\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}$ (182)
$\iota_{4}\int\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{q}}\mathfrak{h}_{i}$ (183)
[…]
Appendix C Detailed Derivations - First Step
The point on the capacity region with maximum $R_{1}$ is achieved at a corner point using Gaussian densities where user 1 allocates its power $P_{1}$ to a private message and user 2 allocates its power $P_{2}$ to a public message. Stating from this corner point, density functions at the end point of the first incremental step are studied. With some misuse of notations, in specifying the entropy of a composite random variable, the subscript in $H$ shows the corresponding linear combination, the superscripts $s$ and $e$ show if it is a starting point or an end point, and the argument shows the total power, e.g., $H^{s}_{V_{1}+\sqrt{b}U_{2}+Z}(P_{1}+bP_{2}+1)$ denotes the entropy of $V_{1}+\sqrt{b}U_{2}+Z$ at the starting point, where the power values of $V_{1}$, $U_{2}$ are equal to $P_{1}$, $P_{2}$, respectively. Notations $R^{s}_{U_{1}}(.)$, $R^{s}_{U_{2}}(.)$, $R^{s}_{V_{1}}(.)$, $R^{s}_{V_{2}}(.)$ and $R^{e}_{U_{1}}(.)$, $R^{e}_{U_{2}}(.)$, $R^{e}_{V_{1}}(.)$, $R^{e}_{V_{2}}(.)$ refer to the rate associated with ${U_{1}}$, ${U_{2}}$, ${V_{1}}$, ${V_{2}}$ at the starting point and at the end point on a step, respectively (function of relevant power values). Movement is achieved by reallocating a small power value of $\delta P_{2}$ form $U_{2}$ to $V_{2}$. Figure 7 shows such a power reallocation. For the first step, we have:
$R^{s}_{U_{1}}$ $=$ 00 0 (184)
$R^{s}_{V_{1}}$ $=$ ${\complement}(P_{1},1)$ (185)
$R^{s}_{U_{2}}$ $=$ $\complement(bP_{2},P_{1}+1)$ (186)
$R^{s}_{V_{2}}$ $=$ 00 0 (187)
where
$\complement(\alpha,\beta)=0.5\log_{2}\left(1+\frac{\alpha}{\beta}\right).$ (188)
At the end point, $U_{2}$ at $Y_{1}$ is subject to the noise
$\frac{1}{\sqrt{b}}V_{1}+V_{2}+\frac{1}{\sqrt{b}}Z$ (189)
while $U_{2}$ at $Y_{2}$ is subject to the noise
$\sqrt{a}V_{1}+V_{2}+Z.$ (190)
Comparing 189 with 190, since $a<1$ and $b<1$, it is concluded that the rate of $U_{2}$ is governed by the mutual information between $U_{2}$ and $Y_{1}$. As a result
$R^{e}_{U_{1}}$ $=$ 00 0 (191)
$R^{e}_{V_{1}}$ $=$ $H^{e}_{V_{1}+\sqrt{b}V_{2}+Z}(P_{1}+b\delta P_{2}+1)-H^{e}_{\sqrt{b}V_{2}+Z}(b\delta P_{2}+1)$ (192)
$R^{e}_{U_{2}}$ $=$ $H^{e}_{V_{1}+\sqrt{b}V_{2}+\sqrt{b}U_{2}+Z}(P_{1}+bP_{2}+1)-H^{e}_{V_{1}+\sqrt{b}V_{2}+Z}(P_{1}+b\delta P_{2}+1)$ (193)
$R^{e}_{V_{2}}$ $=$ $H^{e}_{\sqrt{a}V_{1}+V_{2}+Z}(aP_{1}+\delta P_{2}+1)-H^{e}_{\sqrt{a}V_{1}+Z}(aP_{1}+1)$ (194)
and
$\mathbf{\Upsilon}=\frac{R^{e}_{U_{2}}+R^{e}_{V_{2}}-R^{s}_{U_{2}}-R^{s}_{V_{2}}}{R^{s}_{U_{1}}+R^{s}_{V_{1}}-R^{e}_{U_{1}}-R^{e}_{V_{1}}}.$ (195)
We have
$R^{e}_{U_{2}}+R^{e}_{V_{2}}-R^{s}_{U_{2}}-R^{s}_{V_{2}}$ $=$ (196)
$H^{e}_{V_{1}+\sqrt{b}U_{2}+\sqrt{b}V_{2}+Z}(P_{1}+bP_{2}+1)-H^{e}_{V_{1}+\sqrt{b}V_{2}+Z}(P_{1}+b\delta P_{2}+1)$ $+$ (197)
$H^{e}_{\sqrt{a}V_{1}+V_{2}+Z}(aP_{1}+\delta P_{2}+1)-H^{e}_{\sqrt{a}V_{1}+Z}(aP_{1}+1)$displaystyle-- (198)
$\complement(bP_{2},P_{1}+1)$ (199)
and
$R^{s}_{U_{1}}+R^{s}_{V_{1}}-R^{e}_{U_{1}}-R^{e}_{V_{1}}=\complement(P_{1},1)-H^{e}_{V_{1}+\sqrt{b}V_{2}+Z}(P_{1}+b\delta P_{2}+1)+H^{e}_{\sqrt{b}V_{2}+Z}(b\delta P_{2}+1).$ (200)
Composite random variables appearing in 191200 are
$\hat{C}_{1}$ $=$ $\sqrt{a}V_{1}+V_{2}$ (201)
$\hat{C}_{2}$ $=$ $\sqrt{a}V_{1}$ (202)
$\hat{C}_{3}$ $=$ $V_{1}+\sqrt{b}U_{2}+\sqrt{b}V_{2}$ (203)
$\hat{C}_{4}$ $=$ $V_{1}+\sqrt{b}V_{2}$ (204)
$\hat{C}_{5}$ $=$ $\sqrt{b}V_{2}.$ (205)
To find density functions for the end point, we rely on calculus of variations. Each composite random variable is accompanied by a constraint on its second moment, and a constraint on the area under its density. Relying on calculus of variation, it is concluded that densities of composite random variables are Gaussian. Using equations 201, 202, 203; or 202, 203, 204; or 202, 203, 205; or 203, 204, 205 from 201, 202, 203, 204, 205, one can express $V_{1},U_{2},V_{2}$ in terms of three of the composite random variables. The existence of such an invertible mapping means $V_{1},U_{2},V_{2}$ should be Gaussian as well. Figure 7 depicts the structure of Gaussian code-books for the start and end points. Note that, since $\delta P_{1}=0$, the condition of Pareto minimality is satisfied.
Figure 7: First step moving counterclockwise from the corner point with maximum $R_{1}$. (a),(b) correspond to the starting point, and (c) corresponds to the end point on the first step.