### v4

# On the Optimality of Gaussian Code-books for Signaling over a Two-Users Weak Gaussian Interference Channel

Abstract: This article shows that the capacity region of a two users weak Gaussian interference channel can be achieved using single letter Gaussian code-books. The approach relies on traversing the boundary in incremental steps. Starting from a corner point with Gaussian code-books, and relying on calculus of variation, it is shown that the end point in each step is achieved using Gaussian code-books. Optimality of Gaussian code-books is first established by limiting the random coding to independent and identically distributed scalar (single-letter) samples. Then, it is shown that the value of any optimum solution for vector inputs does not exceed that of the single-letter case. It is also shown that the maximum number of phases needed to realize the optimum time-sharing is two. It is established that the solution to the Han-Kobayashi achievable rate region, with single letter Gaussian code-books, achieves the optimum boundary. Even though the article focuses on weak interference, the results are applicable to the general case.
1 Introduction
Consider a two users weak Gaussian interference channel with parameters shown in Fig. 1, where random coding is limited to independent and identically distributed (i.i.d.) scalars [^1] Theorems 1 to 10 focus on the single-letter code-books. Arguments in Theorems 1 to 10 are valid regardless of the users’ power values. Relying on this property, Theorems 11, 12, 13 establish that optimum time-sharing can be realized by dividing the time axis into (at most) two phases; one phase occupied by both users, and the other phase, if it exists, occupied by one of the two users. It is shown that the capacity region for vector inputs coincides with the single-letter case. This result is established without the need to compute the optimum time-sharing (upper concave envelope). This means: (a) “finding which user should be active over the (potential) single user phase”, (b) “how the total power of the user active over both phases should be divided between the two”, and (c) “how the total time should be divided between the two phases” are not discussed in this work. Arguments in Theorems 1 to 10 are based on single letter code-books, limited to the phase where both users are active. $U_{1},V_{1},U_{2},V_{2}$.
Figure 1: Two users Gaussian Interference Channel (GIC) with $a<1$ and $b<1$.
Focusing on the single-letter case, boundary is traversed by changing the power allocation between public and private message(s), refereed to as “power reallocation” hereafter. Each step starts from a point on the boundary, and then optimum code-books are found such that the corresponding step ends in another point on the boundary. Power reallocation values corresponding to such a step satisfy
$(P_{U_{1}},P_{V_{1}})$ $\xRightarrow[\text{}]{\text{Power reallocation:\,}(\kappa_{1},\eta_{1})}$ $(P_{U_{1}}+\kappa_{1},P_{V_{1}}+\eta_{1}):\penalty 10000\ \kappa_{1}+\eta_{1}=0$ (1)
$(P_{U_{2}},P_{V_{2}})$ $\xRightarrow[\text{}]{\text{Power reallocation:\,}(\kappa_{2},\eta_{2})}$ $(P_{U_{2}}+\kappa_{2},P_{V_{2}}+\eta_{2}):\penalty 10000\ \kappa_{2}+\eta_{2}=0.$ (2)
With some misuse of notations, hereafter power reallocation vectors are denoted as
$(\delta P_{1},\delta P_{2})\penalty 10000\ \mbox{where}\penalty 10000\ \penalty 10000\ \delta P_{1}=|\kappa_{1}|=|\eta_{1}|,\penalty 10000\ \penalty 10000\ \delta P_{2}=|\kappa_{2}|=|\eta_{2}|.$ (3)
In other words, $\delta P_{1}$ denotes the increase in the power of $U_{1}$ or $V_{1}$, depending on which of the two has a higher power at the end point vs. the starting point, and likewise for $\delta P_{2}$ in relation to $U_{2}$ and $V_{2}$. Figure 2 depicts an example where notations $\pm$ vs. $\mp$ are used to emphasize that the signs of $\delta P_{1}$ and $\delta P_{2}$ depend on the step and power reallocation is zero-sum. Power reallocation vector is selected to: (i) support a counter-clockwise move along the boundary, and (ii) guarantee the solution achieving each end point is unique. To achieve the latter criterion, while moving continuously along the boundary, power reallocation vector is selected relying on a notation of admissibility called Pareto minimal [^2] See Theorem 9. Referring to Fig. 2, for a given a power reallocation vector $(\delta P_{1},\delta P_{2})$, the following measure of optimality is used in selecting code-books’ density functions: Given $(\delta P_{1},\delta P_{2})$, maximize [^3] Theorem 10 presents a number of equivalent formulations for the optimization problem achieving boundary points. the length of the step, i.e., $\mathbf{\Gamma}$, over all possible values of the slope $\mathbf{\Upsilon}$ depicted in Fig. 2. Hereafter, $U_{1}$, $U_{2}$, $V_{1}$, $V_{2}$ are called core random variables. Linear combinations of core random variables appearing in mutual information terms forming $\mathbf{\Gamma}$ and $\mathbf{\Upsilon}$ are called compound random variables.
Figure 2: An example for power reallocation and its corresponding step along the boundary.
2 Literature Survey
The problem of Gaussian interference channel has been the subject of numerous outstanding prior works, paving the way to the current point and moving beyond. A subset of these works, reported in [2] to [40], are briefly discussed in this section. A more complete and detailed literature survey will be provided in subsequent revisions of this article
Reference [2] discusses degraded Gaussian interference channel (degraded means one of the two receivers is a degraded version of the other one) and presents multiple bounds and achievable rate regions. Reference [3] studies the capacity of two users GIC for the class of strong interference and shows the capacity region is at the intersection of two MAC regions, consistent with the current article. Reference [4] establishes optimality for two extreme points in the achievable region of the general two users GIC. [4] also proves that the class of degraded Gaussian interference channels is equivalent to the class of Z (one-sided) interference channels.
References [5] to [7] present achievable rate regions for interference channel. In particular, [5] presents the well-known Han-Kobayashi (HK) achievable rate region. HK rate region coincides with all results derived previously (for Gaussian two users GIC), and is shown to be optimum for the class of weak two users GIC in the current article. References [8][10] have further studied the HK rate region. Reference [10] shows that HK achievable rate region is strictly sub-optimum for a class of discrete interference channels.
References [11] to [17] have studied the problem of outer bounds for the interference channel. Among these, [13][14][15] have also provided optimality results in some special cases of weak two users GIC.
References [18][19] have studied the problem of interference channel with common information. References [20] to [22] have studied the problem of interference channel with cooperation between transmitters and/or between receivers. References [23][24] have studied the problem of interference channel with side information. Reference [25] has studied the problem of interference channel assuming cognition, and reference [26] has studied the problem assuming cognition, with or without secret messages.
Reference [27] has found the capacity regions of vector Gaussian interference channels for classes of very strong and aligned strong interference. [27] has also generalized some known results for sum-rate of scalar Z interference, noisy interference, and mixed interference to the case of vector channels. Reference [28] has addressed the sum-rate of the parallel Gaussian interference channel. Sufficient conditions are derived in terms of problem parameters (power budgets and channel coefficients) such that the sum-rate can be realized by independent transmission across sub-channels while treating interference as noise, and corresponding optimum power allocations are computed. Reference [29] studies a Gaussian interference network where each message is encoded by a single transmitter and is aimed at a single receiver. Subject to feeding back the output from receivers to their corresponding transmitter, efficient strategies are developed based on the discrete Fourier transform signaling.
Reference [30] computes the capacity of interference channel within one bit. References [31][32] study the impact of interference in GIC. [32] shows that treating interference as noise in two users GIC achieves the closure of the capacity region to within a constant gap, or within a gap that scales as O(log(log(.)) with signal to noise ratio. Reference [33] relies on game theory to define the notion of a Nash equilibrium region of the interference channel, and characterizes the Nash equilibrium region for: (i) two users linear deterministic interference channel in exact form, and (ii) two users GIC within 1 bit/s/Hz in an approximate form.
Reference [34] studies the problem of two users GIC based on a sliding window superposition coding scheme.
References [35] and [36], independently, introduce the new concept of non-unique decoding as an intermediate alternative to “treating interference as noise”, or “canceling interference”. Reference [37] further studies the concept on non-unique decoding and proves that (in all reported cases) it can be replaced by a special joint unique decoding without penalty.
Reference [38] studies the degrees of freedom of the K-user Gaussian interference channel, and, subject to a mild sufficient condition on the channel gains, presents an expression for the degrees of freedom of the scalar interference channel as a function of the channel matrix.
Reference [39] studies the problem of state-dependent Gaussian interference channel, where two receivers are affected by scaled versions of the same state. The state sequence is (non-causally) known at both transmitters, but not at receivers. Capacity results are established (under certain conditions on channel parameters) in the very strong, strong, and weak interference regimes. For the weak regime, the sum-rate is computed. Reference [40] studies the problem of state-dependent Gaussian interference channel under the assumption of correlated states, and characterizes (either fully or partially) the capacity region or the sum-rate under various channel parameters.
Reference [41] settles the noiseberg conjecture[42] regarding the Han-Kobayashi region of the Gaussian Z-Interference channel with Gaussian signaling.
A summary of main results are provided in Section 3.
3 Summary of Main Results
Section 4 is limited to single letter code-books with power constraints forced to be satisfied with equality. It is shown that independent and identically distributed Gaussian code-books maximize the corresponding weighted sum-rate. Then, in Section 5, it is shown that such single letter code-books, with power constraints forced to be satisfied with equality, are adequate for forming the optimum time sharing (determining the upper concave envelope), thereby for achieving the boundary points.
Section 4.1 presents the main formulation for single letter inputs, followed by some results in Section 4.2. Theorem 1 in Section 4.2 shows that, starting from any point on the boundary and moving counterclockwise for $\mu<1$, the value of $\mathbf{\Upsilon}$ is monotonically decreasing, and the value of $\mathbf{\Gamma}$ is monotonically increasing (see Fig. 2 for definition of $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$). Theorem 2 in Section 4.2 shows that, due to successive decoding in at least one of the receivers, each compound random variable contributes to an entropy term of the form appearing in successive decoding over an additive noise channel. Theorem 3 in Section 4.2 shows that there is a system of invertible linear equations relating compound random variables to core random variables. This means each core random variable can be expressed as a (unique) linear combination of compound random variables. Theorem 4 in Section 4.2 shows that, given $(P_{1},P_{2})$, the dividing of power between public and private messages of each user is such that the mean of each code-book will be zero. As a result, the application of calculus of variation is formulated in terms of zero-mean densities. Theorem 5 in Section 4.2 shows that to achieve a stationary solution for single letter code-books, each compound random variable can be selected based on a zero mean Gaussian density. Since, from Theorem 3, there is a one-to-one linear mapping between core and compound random variables, it follows that core random variables will be zero mean Gaussian as well. Theorem 6 in Section 4.2 shows that the second order variation for stationary solution achieved using Gaussian densities are non-zero. This means the stationary solution is either local maximum or local minimum. Theorem 7 in Section 4.2 shows that, for any given $\mu$, there is a single solution for power allocation over messages (dividing the power of each user between its public and private messages) corresponding to each boundary point. From Theorem 1 (also see Theorem 10), it is concluded that the stationary solution based on Theorems 5, 6, 7 is a (potentially local) maximizer. It remains to impose a condition on power reallocation vector such that each end point is achieved in a unique manner, and the boundary can be traversed continuously. Such a power reallocation is called boundary achieving hereafter. Theorems 8, 9 in Section 4.2 address this issue for all boundary points with $\mu<1$. Theorems 10 in Section 4.2 provides a number of equivalent formulations, as constrained optimization problems, for achieving boundary points of the single-letter rate region. It is concluded that independent and identical Gaussian code-books provide a (potentially local) optimum solution for the corresponding constrained optimization problem.
In Section 5, Theorem 11 establishes that boundary points can be achieved without using vector code-books. Theorems 12, 13 in Section 5 establish that at most two phases are needed (in time-sharing) to realize the upper concave envelope. In one phase both users are active, and in the other phase, if it exists, a single user is active. This is consistent with the result of [41] in optimizing Han-Kobayashi region [5], with Gaussian inputs, for the Z-channel.
Converse results are established in Section 6.
Finally, in Section 7, it is shown that the solution to Han-Kobayashi achievable rate region, with Gaussian random code-books, achieves the optimum boundary.
Remark 1: It is known that capacity region of two users Gaussian Interference Channel (GIC) may include segments achieved by dividing power (of each user) among different two users GICs, called component GICs hereafter. Then, the overall capacity region is obtained by computing the convex hull of regions corresponding to all possible dividing of time and power among component GICs. This results in realizing the upper concave envelope of weighted sum-rate curves corresponding to all possible formations of component GICs. The need for such a formulation is avoided in this work as explained next: The problem formulation presented in Section 4.1, followed by the solution method in Section 4.2, are limited to independent and identically distributed (i.i.d.) scalar (single-letter) samples for $U_{1},V_{1},U_{2},V_{2}$, Then, in Section 5, it will be shown that optimum time-sharing requires at most two phases. Excluding the trivial case of $a=b=0$, in one phase both users are active, and in another phase, only one of the users is active. If power is optimally allocated over time, it would allow forcing the power constraints over each phase to be satisfied with equality. However, in Sections 4.1 and 4.2, the issue of optimum power allocation over time is not explicitly addressed/solved. As a result, forcing power constraints to be satisfied with equality may result in code-books with a non-zero mean (to limit the impact of the interference). Results are established which guarantee code-books’ density functions are zero mean. In summary, since Theorems 1 to 10 in Section 4 are applicable for all values of $P_{1}$ and $P_{2}$, it will be concluded (in Section 5) that the details of the optimum time-sharing and upper concave envelope do not affect the conclusions. Optimum dividing of time/power to realize the upper concave envelope should be applied over (at most) two phases with: (a) independent and identical Gaussian code-books’ densities over time samples in each phase and (b) subject to optimally dividing each user’s power between its public and private messages. $\blacksquare$
Section 4 focuses on single letter code-books, limited to the phase where both users are active.
4 Analysis Limited to Single Letter Code-books
4.1 Formulation for Single Letter Inputs
Consider a two users Gaussian interference channel with inputs $X_{1}$, $X_{2}$ and outputs $Y_{1}$, $Y_{2}$, defined as
$Y_{1}$ $=$ $X_{1}+\sqrt{b}X_{2}+Z_{1}$ (4)
$Y_{2}$ $=$ $\sqrt{a}X_{1}+X_{2}+Z_{2}$ (5)
where $a,b<1$, $Z_{1}$, $Z_{2}$ are additive white Gaussian noise of zero mean and unit variance, and
$X_{1}$ $=$ $U_{1}+V_{1}$ (6)
$X_{2}$ $=$ $U_{2}+V_{2}.$ (7)
Random code-books are formed relying on i.i.d. samples for $U_{1},V_{1},U_{2},V_{2}$. Maximization of weighted sum-rate is expressed as:
$\mbox{Maximize:\penalty 10000\ \,}\penalty 10000\ \penalty 10000\ \penalty 10000\ R_{1}+\mu R_{2}$ $=$ $R_{U_{1}}+R_{V_{1}}+\mu(R_{U_{2}}+R_{V_{2}})$
$\mbox{Subject to:}\penalty 10000\ \penalty 10000\ \penalty 10000\ P_{U_{1}}+P_{V_{1}}$ $=$ $P_{1}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ P_{U_{2}}+P_{V_{2}}$ $=$ $P_{2}.$ (8)
Solving optimization problem in 8 entails: (a) For each user, dividing the power between public and private messages, called power allocation over messages [^4] This is in contrast to power allocation over time which governs how each users’ power should be divided among time samples to optimize time-sharing.. (b) Finding the optimum density functions for each message code-book. (c) Finding encoding/decoding procures for each user. The term coding strategy is used to specify encoding/decoding procedures for each user at a respective point on the boundary. In Section 4.2, the encoding and decoding procedures are limited to single letter code-books (a single sample of $X_{1}$ and a single sample of $X_{2}$), where constraints governing power allocation over messages, see 8, are forced to be satisfied with equality. Then, in Section 5, it is shown that such single letter encoding, in conjunction with two time phases, is adequate for realizing the capacity region.
Capacity region (in the single letter case) is traversed by starting from the point with maximum $R_{1}$ and moving counterclockwise along the lower part of the boundary, i.e., for $\mu<1$. It is known that the point maximizing $R_{1}$ is achieved using Gaussian code-books, where message $X_{1}$ is entirely private, message $X_{2}$ is entirely public, $Y_{1}$ uses successive decoding and $Y_{2}$ treats the interference as noise. Starting from the point with maximum $R_{1}$, in a sequence of infinitesimal steps, $R_{2}$ is gradually increased in a trade-off resulting in decreasing $R_{1}$. Each step involves changing the power allocation (over messages) by infinitesimal amounts. Amounts of reallocated power, $\delta P_{1}$ and $\delta P_{2}$, are small enough such that the coding strategy does not change within the step (can potentially change at the start of the next step).
Let us consider an infinitesimal step from a starting point, specified by superscript $s$, to an end point specified by superscript $e$. The slope $\mathbf{\Upsilon}$ of such a step is defined as:
$\mathbf{\Upsilon}=\frac{\Delta\!R_{2}}{\Delta\!R_{1}}=\frac{R^{e}_{V_{2}}+R^{e}_{U_{2}}-R^{s}_{V_{2}}-R^{s}_{U_{2}}}{R^{s}_{V_{1}}+R^{s}_{U_{1}}-R^{e}_{V_{1}}-R^{e}_{U_{1}}}\triangleq\frac{\mathbf{N}}{\mathbf{D}}$ (9)
where $(R^{s}_{U_{1}},R^{s}_{V_{1}})$, $(R^{s}_{U_{2}},R^{s}_{V_{2}})$ are public and private rates of user 1 and user 2, respectively, at the starting point, likewise, $(R^{e}_{U_{1}},R^{e}_{V_{1}})$, $(R^{e}_{U_{2}},R^{e}_{V_{2}})$ are public and private rates at the end point. Note that $\Delta\!R_{1}$ and $\Delta\!R_{2}$ are defined to be positive, in particular $\Delta\!R_{1}$ is defined as the rate $R_{1}$ at the starting point, minus the rate $R_{1}$ at the end point. Optimality of boundary points is captured in $\mathbf{\Gamma}$ defined as
$\mathbf{\Gamma}=\sqrt{(\Delta\!R_{1})^{2}+(\Delta\!R_{2})^{2}}.$ (10)
The case of $\mu>1$ follows similarly by starting from a point with maximum $R_{2}$ and moving clockwise along the boundary.
Remark 2: The problem of finding the capacity region is complex, since: (a) Power allocation (over messages/time) affects the selection of code-books’ densities. (b) The value of weight $\mu$ changes as one moves along the boundary. (c) One needs to define the infinitesimal steps such that the boundary is covered continuously, and there are unique optimizing single letter code-books for each boundary point. This article does not claim that the coding strategy and its associated code-books’ densities for realizing an achievable rate pair $(R_{1},R_{2})$ are unique, nor that corresponding density functions are limited to be zero-mean Gaussian. The main result to be established in Section 4 (limited to the single letter case) is as follows: For power reallocation vectors which satisfy condition of Theorem 9, or of Theorem 10, zero-mean Gaussian code-books for public and private messages provide a unique solution maximizing $\gamma$ for a given $\upsilon$ in $(\mathbf{\Upsilon},\mathbf{\Gamma})=(\upsilon,\gamma)$. The solution also maximizes $\Delta\!R_{2}$ for a given $\Delta\!R_{1}$. This results in a unique point on the boundary (for single letter code-books). $\blacksquare$
Theorems 1 to 10 in Section 4.2 apply to a component GIC with given $P_{1}$, $P_{2}$.
4.2 Boundary for Single Letter Code-books
Theorem 1 establishes how $\mathbf{\Gamma}$ and $\mathbf{\Upsilon}<1$ change as one moves counterclockwise along the boundary. In essence, it shows that the single-letter rate region based on 8 is convex (in the absence of time-sharing).
Theorem 1.
For $\mu<1$, consider a set of consecutive steps, in counterclockwise direction, along the boudnary of the single letter capacity region for the component GIC based on 8. Corresponding values for $\mathbf{\Upsilon}$ in 9 will be monotonically decreasing, while $\mathbf{\Gamma}$ in 10 will be monotonically increasing.
Proof.
Consider two consecutive infinitesimal steps from point $\mathbb{U}$ to point $\mathbb{V}$ and from point $\mathbb{V}$ to point $\mathbb{W}$, where $\mathbb{U}$, $\mathbb{V}$, $\mathbb{W}$ are within the same boundary segment, i.e., rely on the same strategy. Let us assume $\Delta\!R_{1}$ for the first and second steps are equal to $\delta$, and corresponding $\Delta\!R_{2}$ values are equal to $\hat{\delta}$ and $\check{\delta}$, respectively. Since the boundary is continuous, it is possible to form such two consecutive steps. Noting that the lower part of the boudnary starts from a point with maximum $R_{1}$, and then moves counter-clock wise, we can conclude
$\delta>0,\penalty 10000\ \penalty 10000\ \hat{\delta}>0,\penalty 10000\ \penalty 10000\ \check{\delta}>0.$ (11)
For $\mu<1$, we should have
$\hat{\delta}>\check{\delta}\Longrightarrow\frac{\hat{\delta}}{\delta}>\frac{\check{\delta}}{\delta}.$ (12)
The reason is: (a) Points on the line connecting $\mathbb{U}$ to $\mathbb{W}$ correspond to a sequence of valid power reallocation vectors. (b) Since all steps along a given segment of the boundary rely on the same strategy, one could merge the power reallocation vectors for the first and second steps into a single power reallocation vector, and move directly from $\mathbb{U}$ to $\mathbb{W}$. (c) The weighted sum-rate can be optimized for each power allocation over messages corresponding to points on the line connecting $\mathbb{U}$ to $\mathbb{W}$. Noting (a), (b) and (c), if 12 is violated, it would entail $\mathbb{V}$ falls inside an achievable rate region, violating the initial assumption. From 9, 10, 11 and 12, it follows that
$\hat{\mathbf{\Upsilon}}$ $<$ $\check{\mathbf{\Upsilon}}$ (13)
$\hat{\mathbf{\Gamma}}$ $>$ $\check{\mathbf{\Gamma}}$ (14)
[…]
Theorem 1 above captures the point that, as one moves (counter-clockwise) along a segment of the boundary, $\mu$ will be continuous and monotonically increasing.
Theorem 2.
In at least one of the receivers, $Y_{1}$ and/or $Y_{2}$, public messages $U_{1}$ and $U_{2}$ are recovered using successive decoding.
Proof.
Regardless of code-books’ densities and the method used in recovering $U_{1}$ and $U_{2}$, i.e., joint or successive decoding, we have
$R_{V_{1}}$ $\leq$ $I(V_{1};Y_{1}|U_{1},U_{2})$ (15)
$R_{V_{2}}$ $\leq$ $I(V_{2};Y_{2}|U_{1},U_{2})$ (16)
$R_{U_{1}}$ $\leq$ $I(U_{1};Y_{1}|U_{2})$ (17)
$R_{U_{1}}$ $\leq$ $I(U_{1};Y_{2}|U_{2})$ (18)
$R_{U_{2}}$ $\leq$ $I(U_{2};Y_{1}|U_{1})$ (19)
$R_{U_{2}}$ $\leq$ $I(U_{2};Y_{2}|U_{1})$ (20)
$R_{U_{1}}+R_{U_{2}}$ $\leq$ $I(U_{1},U_{2};Y_{1})$ (21)
$R_{U_{1}}+R_{U_{2}}$ $\leq$ $I(U_{1},U_{2};Y_{2}).$ (22)
Consider a linear programming problem for maximizing $R_{U_{1}}+R_{V_{1}}+\mu R_{U_{2}}+\mu R_{U_{2}}$ subject to constrains in 15 to 22. Noting the objective function is composed of four variables, in the optimum solution (at least) four of the constraints should be active, i.e., satisfied with equality. Constraints 15 and 16 should be active, otherwise $R_{V_{1}}$ and/or $R_{V_{2}}$ could be increased without affecting constraints in 17 to 22. This means at least two of the constraints in 17 to 22 should be active. If any of the constraints 17 to 20 is active, it entails successive decoding at a respective receiver,and proof is complected. Otherwise, 21 and 22 should be both active, i.e.,
$R_{U_{1}}+R_{U_{2}}$ $=$ $I(U_{1},U_{2};Y_{1})\,=\,I(U_{2};Y_{1})+I(U_{1};Y_{1}|U_{2})$ (23)
$=$ $I(U_{1},U_{2};Y_{2})\,=\,I(U_{2};Y_{2})+I(U_{1};Y_{2}|U_{2})$ (24)
$=$ $I(U_{1},U_{2};Y_{1})\,=\,I(U_{1};Y_{1})+I(U_{2};Y_{1}|U_{1})$ (25)
$=$ $I(U_{1},U_{2};Y_{2})\,=\,I(U_{1};Y_{2})+I(U_{2};Y_{2}|U_{1}).$ (26)
We need to show that, in addition to 21 and 22, at least one of the constraints 17 to 20 becomes active. Let us assume 17 to 20 are satisfied with strict inequality. Combining 17, 18, 19, 20 with 23, 24, 25, 26, respectively, would result in
[…]
where $\varkappa_{1},\varkappa_{2},\varkappa_{3},\varkappa_{4}>0$. Replacing 27, 28, 29, 30 in $R_{U_{1}}+\mu R_{U_{2}}$, we have
$R_{U_{1}}+\mu R_{U_{2}}$ $=$ $I(U_{1};Y_{1}|U_{2})+\mu\,I(U_{2};Y_{1})-\varkappa_{1}\,(1-\mu)$ (31)
$R_{U_{1}}+\mu R_{U_{2}}$ $=$ $I(U_{1};Y_{2}|U_{2})+\mu\,I(U_{2};Y_{2})-\varkappa_{2}\,(1-\mu)$ (32)
$R_{U_{1}}+\mu R_{U_{2}}$ $=$ $I(U_{1};Y_{1})+\mu\,I(U_{2};Y_{1}|U_{1})-\varkappa_{3}\,(\mu-1)$ (33)
$R_{U_{1}}+\mu R_{U_{2}}$ $=$ $I(U_{1};Y_{2})+\mu\,I(U_{2};Y_{2}|U_{1})-\varkappa_{4}\,(\mu-1).$ (34)
Noting 31 and 32, it is concluded that for $\mu<1$, the weighted sum-rate is maximized when $\varkappa_{1}=0$ and/or $\varkappa_{2}=0$. Likewise, noting 33 and 34, for $\mu>1$, the weighted sum-rate is maximized when $\varkappa_{3}=0$ and/or $\varkappa_{4}=0$. These conditions entail successive decoding at respective receiver(s). ∎
Remark 3: Without loss of generality, in the rest of the paper it is assumed that 18 is the active constraint among 17 to 20. Relying on steps similar to Theorem 2, it follows that 21 and 22 will be both active as well. This results in
$R_{U_{1}}$ $=$ $I(U_{1};Y_{2}|U_{2})$ (35)
$R_{U_{2}}$ $=$ $I(U_{2};Y_{2})$ (36)
$R_{U_{1}}+R_{U_{2}}$ $=$ $I(U_{2};Y_{2})+I(U_{1};Y_{2}|U_{2})$ (37)
$R_{U_{1}}+R_{U_{2}}$ $=$ $I(U_{1},U_{2};Y_{1})$ (38)
where 35, 36, 37 capture successive decoding of $U_{2}$ followed by $U_{1}$ at $Y_{2}$, and 38 captures joint decoding of $(U_{1},U_{2})$ at $Y_{1}$. Given 35, 36, 37, we have
$R_{U_{1}}$ $=$ $\overbrace{I(U_{1};Y_{2}|U_{2})}^{(a)}\geq\overbrace{I(U_{1};Y_{2})}^{(b)}$ (39)
$R_{U_{2}}$ $=$ $\overbrace{I(U_{2};Y_{2})}^{(c)}\leq\overbrace{I(U_{2};Y_{2}|U_{1})}^{(d)}$ (40)
$R_{U_{1}}+R_{U_{2}}$ $=$ $\overbrace{I(U_{1};Y_{2}|U_{2})}^{(a)}+\overbrace{I(U_{2};Y_{2})}^{(c)}$ (41)
$=$ $\overbrace{I(U_{1};Y_{2})}^{(b)}+\overbrace{I(U_{2};Y_{2}|U_{1})}^{(d)}.$ (42)
Noting 41 and 42 are satisfied with equality, it can be concluded that that
$\overbrace{I(U_{1};Y_{2}|U_{2})-I(U_{1};Y_{2})}^{(e)}=\overbrace{I(U_{2};Y_{2}|U_{1})-I(U_{2};Y_{2})}^{(f)}.$ (43)
In 43, term (e) captures the (potential) increase in $R_{U_{1}}$ based on the decoding order $U_{2}$ followed by $U_{1}$ at $Y_{2}$ and term (f) captures the (potential) increase in $R_{U_{2}}$ based on the decoding order $U_{1}$ followed by $U_{2}$ at $Y_{2}$. Noting that (e) and (f) are equal, it follows that for $\mu<1$, the former case corresponding to term (e) results in a higher value for $R_{U_{1}}+\mu R_{U_{2}}$. This means 35 to 38 apply to the lower part of the boundary. $\blacksquare$
Theorem 3 establishes that core random variables $U_{1}$$V_{1}$$U_{2}$$V_{2}$ are a unique linear combination of compound random variables occurring in successive decoding at $Y_{1}$ or at $Y_{2}$. This property will be used to show that if such compound random variables are jointly Gaussian, then $U_{1}$, $V_{1}$, $U_{2}$, $V_{2}$ will be Gaussian as well.
Theorem 3.
There exits at least one invertible $4\times 4$ matrix allowing to express core random variables in terms of compound random variables.
Proof.
Without loss of generality, let us assume $a\neq 0$, and focus on the case that successive decoding of public message(s) is performed at $Y_{2}$. Consider compound random variables $C_{1}$ to $C_{4}$ involved in successive decoding at $Y_{2}$. We have
$C_{1}$ $=$ $\sqrt{a}U_{1}+\sqrt{a}V_{1}+U_{2}+V_{2}$ (44)
$C_{2}$ $=$ $\sqrt{a}U_{1}+\sqrt{a}V_{1}+V_{2}$ (45)
$C_{3}$ $=$ $\sqrt{a}V_{1}+V_{2}$ (46)
$C_{4}$ $=$ $\sqrt{a}V_{1}.$ (47)
Matrix of linear coefficients forming 44, 45, 46, 47 is equal to
$\begin{bmatrix}\sqrt{a}&\sqrt{a}&1&1\\ \sqrt{a}&\sqrt{a}&0&1\\ 0&\sqrt{a}&0&1\\ 0&\sqrt{a}&0&0\end{bmatrix}$. (48)
It easily follows that the matrix in 48 is invertible $\forall a\neq 0$. A similar result can be concluded for $a=0,\forall b\neq 0$. Note that Theorem 3 includes all core random variables $U_{1},V_{1},U_{2},V_{2}$. A similar result follows if $U_{1}$ or $U_{2}$ is zero. ∎
It is assumed that the decoding strategy in Remark 3 (captured in 35 to 38), which is compatible with Theorem 3 (captured in 44 to 47), with $a\neq 0$ applies throughout this article. This means $R_{U_{1}}$, $R_{V_{1}}$, $R_{U_{2}}$ are governed by a cascade of additive noise channels due to successive decoding at $Y_{2}$ and $R_{V_{1}}$ is governed by an additive noise channel at $Y_{1}$. As a result, rate values contributing to $\mathbf{\Upsilon}$, $\mathbf{\Gamma}$ in 9, 10, respectively, correspond to independent additive noise channels depicted in Fig. 3.
Figure 3: Channel models depicting decoding methods discussed in Theorem 3 where 3(a) corresponds to successive decoding of $U_{2}$, $U_{1}$ and $V_{2}$ at $Y_{2}$, and 3(b) corresponds to joint decoding of $(U_{1},U_{2})$ followed by decoding of $V_{1}$ at $Y_{1}$.
Since power constraints are forced to be satisfied with equality, a stationary solution may include cases that code-books’ densities have a non-zero statistical mean. Following example clarifies this point.
Example:
Figure 4: Example of a channel where the stationary solution for mutual information may result in a maximum or a minimum, according to the statistical mean of $\tilde{Z}$.
Consider the channel in Fig. 4, where $\tilde{X}$, $\tilde{Z}$ and $Z$ are independent, and $\int\vartheta^{2}f_{\tilde{Z}}(\vartheta)d\vartheta=P_{\!\tilde{Z}}$. Let us define
$\hat{f}$ $=$ $f_{\tilde{X}}\ast f_{\tilde{Z}}\ast{\cal N}(0,1)$ (49)
$=$ $f_{\tilde{Z}}+{\cal N}(0,2)$ (50)
$\check{f}$ $=$ $f_{\tilde{Z}}\ast{\cal N}(0,1)$ (51)
where ${\cal N}(\mathfrak{u},\mathfrak{s})$ is a Gaussian density with statistical average $\mathfrak{u}$ and variance $\mathfrak{s}$. We have
$I(\tilde{X};\tilde{Y})=\mathsf{H}^{\hat{f}}-\mathsf{H}^{\check{f}}.$ (52)
It follows that
[…]
$\max_{\hat{f},\check{f}}I(\tilde{X};\tilde{Y})\penalty 10000\ \penalty 10000\ \mbox{is achieved for}$ $\!\!f_{\tilde{Z}}={\cal N}(\sqrt{P_{\!\tilde{Z}}},0).$ (54)
A non-zero statistical mean entails the power $P_{\!\tilde{Z}}$ is intentionally wasted to avoid interference. $\blacksquare$
Theorem 4 establishes that code-books densities for $U_{1}$, $V_{1}$, $U_{2}$ and $V_{2}$ are zero-mean. This entails a case similar to the above example will not be encountered in code-books’ densities forming the capacity region in this work.
Theorem 4.
Code-books’ densities for (single letter) $U_{1}$, $V_{1}$, $U_{2}$, $V_{2}$ are zero mean for $\forall P_{1},\forall P_{2}$.
Proof.
Code-books’ densities for public messages $U_{1}$ and $U_{2}$ are zero mean. The reason is that, instead of wasting the allocated power values $P_{U_{1}}$ and/or $P_{U_{2}}$ relying on a non-zero mean value, the variance of corresponding code-book(s) can be increased which in turn increases $R_{U_{1}}$ and/or $R_{U_{2}}$ while satisfying the condition that public messages should be recoverable at both receivers. On the other hand, if the code-books’ densities for private messages $V_{1}$ and/or $V_{2}$ have a non-zero mean, the wasted power can be allocated to the corresponding public message, increasing $R_{U_{1}}$ and/or $R_{U_{2}}$, while guaranteeing public and private messages can be decoded. ∎
Relying on Theorem 4, all optimization problems involving application of calculus of variation are formulated in terms of zero-mean densities.
Theorem 5 shows that Gaussian code-books result in a stationary solution for $\mathbf{\Upsilon}$, $\mathbf{\Gamma}$, $\Delta\!R_{1}$ and $\Delta\!R_{2}$.
Theorem 5.
Gaussian densities for $U_{1}$, $V_{1}$, $U_{2}$, $V_{2}$ result in a stationary solution for $\mathbf{\Upsilon}$, $\mathbf{\Gamma}$, $\Delta\!R_{1}$ and $\Delta\!R_{2}$.
Proof.
Appendix B establishes that Gaussian densities for compound random variables result in a stationary solution for $\mathbf{\Upsilon}$, as well as for $\mathbf{\Gamma}$. In the following, it is established that densities for core random variables will be Gaussian as well. Let us focus on $\mathbf{N}$, i.e.,
$\mathbf{N}\equiv R^{e}_{V_{2}}+R^{e}_{U_{2}}-R^{s}_{V_{2}}-R^{s}_{U_{2}}$ (55)
where $R^{s}_{V_{2}},R^{s}_{U_{2}}$ are fixed and $R^{e}_{V_{2}},R^{e}_{U_{2}}$ should be optimized. Let us focus on the case shown in channel models in Fig. 3. This means $R_{U_{2}}^{e}$ and $R_{V_{2}}^{e}$, forming $\mathbf{N}$ in 55, are mutual information terms across two channels formed at $Y_{2}$, each with an additive noise independent of its input. Mutual information terms forming $R^{e}_{V_{2}}$ and $R^{e}_{U_{2}}$ are each composed of two entropy terms (likewise for $R^{e}_{V_{1}}$ and $R^{e}_{U_{1}}$ appearing in $\mathbf{D}$). For simplicity, formulations do not explicitly include the role of Gaussian noise terms added at $Y_{1}$ and $Y_{2}$. Let us use notations $\mathbf{p}_{i}$, $i=1,2,3,4$ to refer to densities of compound random variables appearing in entropy terms in $R^{e}_{U_{2}}$ and $R^{e}_{V_{2}}$. From Fig. 3, we have
$\mathbf{p}_{1}:$ density function of compound random variable (signal): $\!\!\sqrt{a}U_{1}+\sqrt{a}V_{1}+U_{2}+V_{2}$ (56)
$\mathbf{p}_{2}:$ density function of compound random variable (noise): $\!\!\sqrt{a}U_{1}+\sqrt{a}V_{1}+V_{2}$ (57)
$\mathbf{p}_{3}:$ density function of compound random variable (signal): $\!\!\sqrt{a}V_{1}+V_{2}$ (58)
$\mathbf{p}_{4}:$ density function of compound random variable (noise): $\!\!\sqrt{a}V_{1}.$ (59)
Likewise, in $\mathbf{D}$, term $R_{U_{1}}^{e}$ is governed by an additive noise channel formed at $Y_{2}$ with signal/noise captured in 57/58, respectively, and term $R_{V_{1}}^{e}$ is governed by an additive noise channel formed at $Y_{1}$ (after $U_{1},U_{2}$ are jointly decoded). Relevant entropy terms include two additional compound random variables with densities $\mathbf{p}_{5}$, $\mathbf{p}_{6}$ where
$\mathbf{p}_{5}:$ density function of compound random variable (signal): $\!\!V_{1}+\sqrt{b}V_{2}$ (60)
$\mathbf{p}_{6}:$ density function of compound random variable (noise): $\!\!\sqrt{b}V_{2}.$ (61)
Since $U_{1}$, $U_{2}$, $V_{1}$, $V_{2}$ are independent of each other, each $\mathbf{p}_{i}$, $i=1,2,3,4,5,6$ can be expressed in terms of a convolution. In applying calculus of variation, density functions of core random variables $U_{1}$, $U_{2}$, $V_{1}$, $V_{2}$ are assumed to be zero mean, and constraints on “power" and “area under each density function" are added to the objective function using Lagrange multipliers. Then, density functions of $U_{1}$, $U_{2}$, $V_{1}$, $V_{2}$ are perturbed using $\epsilon_{1}h_{1}$, $\epsilon_{2}h_{2}$, $\epsilon_{3}h_{3}$ and $\epsilon_{4}h_{4}$. Setting the derivatives of 9 with respect to $\epsilon_{i}$, $i=1,2,3,4$ equal to zero results in
$\frac{\partial\mathbf{\Upsilon}}{\partial\epsilon_{i}}\Big|_{\epsilon_{i}=0}=0\penalty 10000\ \penalty 10000\ \Longrightarrow\penalty 10000\ \penalty 10000\ \left(\frac{\partial\mathbf{N}}{\partial\epsilon_{i}}\mathbf{D}-\frac{\partial\mathbf{D}}{\partial\epsilon_{i}}\mathbf{N}\right)\Big|_{\epsilon_{i}=0}=0.$ (62)
Constraints on powers of core random variables are
$P_{U_{1}}+P_{V_{1}}$ $=$ $P_{1}$ (63)
$P_{U_{2}}+P_{V_{2}}$ $=$ $P_{2}.$ (64)
Power constraints in 63, 64 are expressed in terms a larger set, with each constraint limiting the power of a compound random variable. Power constraints in this larger set are linearly dependent, causing redundancy. However, since constraints in the enlarged set are consistent, imposing redundancy does not affect the validity of the final solution. A similar set of redundant constraints are used in imposing the restriction that the area under each density function should be equal to one. Under these conditions, relying on a formulation similar to [43] (see page 335), it follows that
$\frac{\partial\mathbf{N}}{\partial\epsilon_{i}}\Big|_{\epsilon_{i}=0}\penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{and}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \frac{\partial\mathbf{D}}{\partial\epsilon_{i}}\Big|_{\epsilon_{i}=0}$ (65)
in 62 will be zero if densities of $U_{1}$, $U_{2}$, $V_{1}$, $V_{2}$ are zero-mean Gaussian. Appendix B includes some details in applying calculus of variation to $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$. It follows that the same Gaussian densities for core random variables which result in a stationary solution for $\mathbf{\Upsilon}$, also result in a stationary solution for $\mathbf{\Gamma}$ (see Remark 4). Applying calculus of variation directly to $(\Delta\!R_{1},\Delta\!R_{2})$, it follows that Gaussian densities for core random variables result in a stationary solution for $\Delta\!R_{1}$ and $\Delta\!R_{2}$. As a result, starting from a point on the boundary, $R_{1}$ and $R_{2}$ at the end point will be stationary solutions as well. ∎
Remark 4: Noting Theorem 5, since the same Gaussian code-books result in a stationary solution for $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$, it will be a stationary solution to the optimization problem:
$\mbox{Maximize}\penalty 10000\ \penalty 10000\ \mathbf{\Gamma}+\mathfrak{L}_{\Upsilon}\mathbf{\Upsilon}$ (66)
where $\mathfrak{L}_{\Upsilon}$ is a Lagrange multiplier in maximizing $\mathbf{\Gamma}$ subject to $\mathbf{\Upsilon}\geq\hat{\Upsilon}$ for given $\hat{\Upsilon}$. Theorem 10 provides more details. It is concluded that $\mu$ in $R_{1}+\mu R_{2}$ is a Lagrange multiplier in maximizing $R_{1}$ subject to $R_{2}\geq\hat{R}_{2}$ for given $\hat{R}_{2}$. $\blacksquare$
Next, Theorem 6 shows that the second order variation for $(\Delta\!R_{1},\Delta\!R_{2})$ is non-zero, even though entropy terms are added with positive and negative values. This means stationary solution is a (potentially local ) maximum/minimum. Note that: (a) Derivations in Theorem 6 are applicable for any valid power reallocation vector. (b) Theorem 8 shows that one can take a step along a boundary segment relying on a “unique”, boundary achieving, power reallocation vector. Considering (a) and (b), in conjunctions with Theorems 1, 6, 7 (also see Theorem 10), it is concluded that, given a unique boundary achieving power reallocation vector, such a stationary solution is a unique (potentially local) maximizer.
Theorem 6.
Given a valid power reallocation vector, the second order variations for $\Delta\!R_{1}$ and $\Delta\!R_{2}$ are non-zero.
Proof.
Fixing power reallocation vector, let us limit the proof to cases where both $\Delta\!R_{1}$ and $\Delta\!R_{2}$ are stationary solutions (in terms of density functions for the given power reallocation vector). Otherwise (i.e., if only one of the two, $\Delta\!R_{1}$ or $\Delta\!R_{2}$, is a stationary solution), one could move along a vertical (or a horizontal) line by perturbing density functions to reach to a point where both $\Delta\!R_{1}$ and $\Delta\!R_{2}$ are stationary solutions. This is possible unless $a=b=0$ for which capacity region is rectangular [^5] Moving along a vertical (or a horizontal) line is possible only if the point is within the boundary. Since the movement along the lower part of the boundary is counterclockwise, it avoids the initial vertical line with $\Delta\!R_{1}=0$, $\Delta\!R_{2}\neq 0$. A similar argument applies to the upper part for which the movement is clockwise and the initial horizontal line is avoided.. To have an inflection point for $R_{1}+\mu R_{2}$, second order variation should be zero for both $\Delta\!R_{1}$ and $\Delta\!R_{2}$. In addition, noting the starting point is optimum, and $\Delta\!R_{1}$, $\Delta\!R_{2}$ are both stationary solutions, it follows that end point is associated with a stationary solution for both $R_{1}$ and $R_{2}$. Proof proceeds by expressing the stationary conditions in terms $R_{1}$ and $R_{2}$.
Consider a valid power reallocation vector, i.e., one resulting in an end point with $P_{U_{1}}\geq 0$, $P_{V_{1}}\geq 0$, $P_{U_{2}}\geq 0$, $P_{V_{2}}\geq 0$, and let us focus on
$\mathsf{H}^{\digamma_{\!\!1}}-\mathsf{H}^{\digamma_{\!\!2}}.$ (67)
In the following, rates $R_{U_{1}}$, $R_{U_{2}}$, $R_{V_{1}}$, $R_{V_{2}}$ are expressed in terms of 67 by identifying compound random variables acting as arguments of $\digamma_{\!\!1}$ and $\digamma_{\!\!2}$. Any given (valid) power reallocation vector is associated with a (valid) set of values for the power of associated compound random variables. For simplicity of notation, terms corresponding to Gaussian noise are ignored. From Theorem 2, we conclude
$R_{U_{1}}\Longrightarrow$ $\digamma_{\!\!1}\thicksim\sqrt{a}U_{1}+\sqrt{a}V_{1}+V_{2},$ $\digamma_{\!\!2}\thicksim\sqrt{a}V_{1}+V_{2}$ (68)
$R_{V_{1}}\Longrightarrow$ $\digamma_{\!\!1}\thicksim V_{1}+\sqrt{b}V_{2},$ $\digamma_{\!\!2}\thicksim\sqrt{b}V_{2}$ (69)
$R_{U_{2}}\Longrightarrow$ $\digamma_{\!\!1}\thicksim\sqrt{a}U_{1}+\sqrt{a}V_{1}+U_{2}+V_{2},$ $\digamma_{\!\!2}\thicksim\sqrt{a}U_{1}+\sqrt{a}V_{1}+V_{2}$ (70)
$R_{V_{2}}\Longrightarrow$ $\digamma_{\!\!1}\thicksim\sqrt{a}V_{1}+V_{2},$ $\digamma_{\!\!2}\thicksim\sqrt{a}V_{1}.$ (71)
Let us use $\epsilon_{U_{1}}h_{U_{1}}$, $\epsilon_{V_{1}}h_{V_{1}}$, $\epsilon_{U_{2}}h_{U_{2}}$, $\epsilon_{V_{2}}h_{V_{2}}$ to perturb the density functions of $U_{1},V_{1},U_{2},V_{2}$, respectively. From 265, second order variations for compound random variables in 68 to 71 are
[…]
$\overbrace{\frac{\left[h_{U_{1}}\left(\frac{u_{1}}{\sqrt{a}}\right)\ast f_{V_{1}}\left(\frac{v_{1}}{\sqrt{a}}\right)\ast f_{V_{2}}(v_{2})\right]^{2}}{a\,f_{U_{1}}\left(\frac{u_{1}}{\sqrt{a}}\right)\ast f_{V_{1}}\left(\frac{v_{1}}{\sqrt{a}}\right)\ast f_{V_{2}}(v_{2})}}^{T_{1}}$
[…]
Consider the variation of $R_{U_{1}}+R_{V_{1}}$ with respect to $\epsilon_{U_{1}}$ and $\epsilon_{V_{1}}$. We have
[…]
Consider second order variation in terms of $\epsilon_{U_{1}}$, i.e., when $P_{U_{1}}\neq 0$. From 81, the second variation for $R_{U_{1}}+R_{V_{1}}$ will be zero if $T_{1}=0$. Replacing $T_{1}=0$ in 83, this requires $T_{2}=0$. Condition $T_{1}=T_{2}=0$ can be satisfied only if $h_{U_{1}}(u_{1})=0$, resulting in a trivial case. Next, consider the second order variation due to $\epsilon_{V_{1}}$. From LABEL:E98, the condition requires $T_{3}=T_{4}-T_{5}$, and from 84, the condition requires $T_{6}=T_{3}-T_{4}$. Combining the two conditions, we obtain the condition $T_{6}=-T_{5}$, which cannot happen (unless in the trivial case of $h_{V_{1}}(v_{1})=0$) since $T_{5}\geq 0$ and $T_{6}\geq 0$. Relaying on similar derivations, it follows that second order variation for $f_{U_{2}}+\epsilon_{U_{2}}h_{U_{2}}(u_{2})$ and $f_{V_{2}}+\epsilon_{V_{2}}h_{V_{2}}(v_{2})$ will be non-zero as well.
As mentioned in Remark 1, capacity region is potentially enlarged by forming the convex hull of component GICs. Next, Theorem 7, focusing on a component GIC with given $P_{1}$ and $P_{2}$, shows that there is a single power allocation (over messages) for any given point on the corresponding (single letter) boundary. Theorem 7 is used in conjunction with: (a) Theorem 5, and noting a zero mean Gaussian density is entirely specified by its second moment, to establish that each boundary point is achieved by a unique single letter Gaussian code-book $\forall P_{1}$, $\forall P_{2}$, $\forall\mu<1$. (b) Theorems 8 and 9 to establish that the boundary of each component GIC can be traversed by (incrementally) moving thorough a unique set of points. (c) Theorem 10 to establish that the stationary solution corresponding to Theorems 5 and 6 is a (potentially local) maximizer for $R_{1}+\mu R_{2}$.
Theorem 7.
Given $\mu<1$, there is a single power allocation over messages achieving a point on the boundary of the component GIC corresponding to $P_{1}$ and $P_{2}$, $\forall P_{1}$, $\forall P_{2}$.
Proof.
Given $\mu<1$, let us use $\mathsf{W}(\mu)$ to refer to the corresponding optimum weighted sum-rate, i.e.,
$\mathsf{W}(\mu)\equiv\max(R_{1}+\mu R_{2}).$ (85)
Consider the following two power allocations (over messages) for users $X_{1}$, $X_{2}$, refereed to as $1^{st}$ and $2^{nd}$, and distinguished by superscripts 1,2:
$1^{st}\mbox{\penalty 10000\ power allocation for user\penalty 10000\ }X_{1}:(P^{1}_{U_{1}},P^{1}_{V_{1}})=(\mathsf{t}^{1}_{1},1-\mathsf{t}^{1}_{1})P_{1}$ $\equiv$ $\mathbf{p}_{1}^{1}$ (86)
$1^{st}\mbox{\penalty 10000\ power allocation for user\penalty 10000\ }X_{2}:(P^{1}_{U_{2}},P^{1}_{V_{2}})=(\mathsf{t}^{1}_{2},1-\mathsf{t}^{1}_{2})P_{2}$ $\equiv$ $\mathbf{p}_{2}^{1}$ (87)
$2^{nd}\mbox{\penalty 10000\ power allocation for user\penalty 10000\ }X_{1}:(P^{2}_{U_{1}},P^{2}_{V_{1}})=(\mathsf{t}^{2}_{1},1-\mathsf{t}^{2}_{1})P_{1}$ $\equiv$ $\mathbf{p}_{1}^{2}$ (88)
$2^{nd}\mbox{\penalty 10000\ power allocation for user\penalty 10000\ }X_{2}:(P^{2}_{U_{2}},P^{2}_{V_{2}})=(\mathsf{t}^{2}_{2},1-\mathsf{t}^{2}_{2})P_{2}$ $\equiv$ $\mathbf{p}_{2}^{2}$ (89)
where $\mathsf{t}^{1}_{1},\mathsf{t}^{1}_{2},\mathsf{t}^{2}_{1},\mathsf{t}^{2}_{2}\in[0,1]$. Consider applying calculus of variation in conjunction with power allocation (over messages) 4-tuples $(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})$, as well as in conjunction with power allocation (over messages) 4-tuples $(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})$. According to Theorem 4, the corresponding stationary solutions rely on zero-mean Gaussian code-books for core random variables $U_{1}$, $V_{1}$, $U_{2}$, $V_{2}$. Let us assume the two solutions result in the same point on the boundary, i.e.,
$\mathsf{W^{1}}(\mu)=\mathsf{W^{2}}(\mu)=\max(R_{1}+\mu R_{2})$ (90)
where superscripts 1,2 correspond to power allocations (over messages) corresponding to $(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})$ and $(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})$, respectively. Consider power allocation (over messages) 4-tuples obtained by time sharing between $(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})$ and $(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})$, i.e.,
$(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})\mathsf{T}+(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})(1-\mathsf{T}),\penalty 10000\ \mathsf{T}\in[0,1].$ (91)
Such a time-sharing results in the same value of $\mathsf{W^{1}}(\mu)=\mathsf{W^{2}}(\mu)$ for the weighted sum-rate. On the other hand, if
$(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})\neq(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})$ (92)
it follows that
$(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})\neq(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})\neq(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})\mathsf{T}+(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})(1-\mathsf{T})\penalty 10000\ \mbox{for}\penalty 10000\ \mathsf{T}\neq 0,1.$ (93)
For given $\mu$, let us apply calculus of variation in conjunction with power allocation (over messages) $(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})\mathsf{T}+(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})(1-\mathsf{T})\penalty 10000\ \mbox{for}\penalty 10000\ \mathsf{T}\neq 0,1$. This results in a solution, using zero-mean Gaussian densities for core random variables, optimized subject to power allocation (over messages) given in 91. If the first and the second power allocation vectors result in the same point on the boundary, the weighted sum-rate obtained through time sharing between them, denoted as $\mathsf{\breve{W}}(\mu)$, should be equal to $\mathsf{\breve{W}}(\mu)=\mathsf{W^{1}}(\mu)=\mathsf{W^{2}}(\mu)$, $\forall\mathsf{T}:0<\mathsf{T}<1$, i.e., a continuum of power allocations (over messages) result in the same value for $R_{1}+\mu R_{2}$. This can happen only in the trivial case of $a=b=0$. This entails, for a non-trivial component GIC, $(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})$ and $(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})$ with $(\mathbf{p}_{1}^{1},\mathbf{p}_{2}^{1})\neq(\mathbf{p}_{1}^{2},\mathbf{p}_{2}^{2})$, cannot result in the same point on the boundary. ∎
Next, the condition for a power reallocation vector to be boundary achieving is discussed. Let us consider a step along the boundary which is small enough such that the coding strategy remains the same within the step. Let us assume $(\hat{\Delta}\!P_{1},\hat{\Delta}\!P_{2})$ is the power reallocation vector corresponding to an end point beyond which a change in strategy is needed, and consider
$(\Delta\!P_{1},\Delta\!P_{2}):\Delta\!P_{1}\leq\hat{\Delta}\!P_{1}\penalty 10000\ \penalty 10000\ \mbox{and}\penalty 10000\ \penalty 10000\ \Delta\!P_{2}\leq\hat{\Delta}\!P_{2}.$ (94)
Let us define $\upsilon\leq\mu^{s}$, where $\mu^{s}$ is the value of $\mu$ at the starting point, and the set $\bar{\mathsf{S}}_{\upsilon}$ as
$\bar{\mathsf{S}}_{\upsilon}=\left\{f_{U_{1}},f_{V_{1}},f_{U_{2}},f_{V_{2}}\!:\!\penalty 10000\ \mbox{outgoing slope at the starting point is}\penalty 10000\ \upsilon\triangleq\min_{(\delta P_{1},\delta P_{2})\in[0,\Delta\!P_{1}]\times[0,\Delta\!P_{2}]}\mathbf{\Upsilon}\right\}.$ (95)
Set $\bar{\mathsf{S}}_{\upsilon}$ is defined over all possible code-books’ densities, including Gaussian. Each member of 95 corresponds to a power reallocation vector $(\delta P_{1},\delta P_{2})\in[0,\Delta\!P_{1}]\times[0,\Delta\!P_{2}]$. This correspondence is potentially many-to-one since multiple choices for densities $(f_{U_{1}},f_{V_{1}},f_{U_{2}},f_{V_{2}})$, with the same $(\delta P_{1},\delta P_{2})$, may achieve the same $\mathbf{\Upsilon}=\upsilon$. Given $\mathbf{\Upsilon}=\upsilon$, the size of the set $\bar{\mathsf{S}}_{\upsilon}$ is reduced by limiting it to choice(s) which maximize $\mathbf{\Gamma}$. Maximum value of $\mathbf{\Gamma}$ over the set $\bar{\mathsf{S}}_{\upsilon}$ is denoted as $\varkappa_{\upsilon}$. Let us consider a second set $\bar{\bar{\mathsf{S}}}_{\upsilon}$ where
$\bar{\bar{\mathsf{S}}}_{\upsilon}\subseteq\bar{\mathsf{S}}_{\upsilon}:\penalty 10000\ \mathbf{\Gamma}=\varkappa_{\upsilon}.$ (96)
The set $\bar{\bar{\mathsf{S}}}_{\upsilon}$ includes a point on the boundary with
$\mathbf{\Upsilon}=\upsilon\penalty 10000\ \penalty 10000\ \mbox{and}\penalty 10000\ \penalty 10000\ \mathbf{\Gamma}=\varkappa_{\upsilon}\triangleq\max_{\mathbf{\Upsilon}=\upsilon}\mathbf{\Gamma}.$ (97)
We are interested in establishing that the size of $\bar{\bar{\mathsf{S}}}_{\upsilon}$ can be reduced, by increasing $\upsilon$, such that the shrunken set includes a single element, say $\zeta$. Since $\bar{\bar{\mathsf{S}}}_{\upsilon}$ always includes a point on the boundary, it follows that $\zeta$ falls on the boundary. In addition, we need to show that: (i) $\zeta$ is realized using Gaussian code-books, and (ii) the rest of the boundary can be covered starting from $\zeta$. Theorem 8 addresses these requirements.
Theorem 8.
Cardinality of the set $\bar{\bar{\mathsf{S}}}_{\upsilon}$ can be reduced, by increasing $\upsilon<\mu^{s}$, in a recursive manner, such that the final set is associated with a single $(\delta P_{1},\delta P_{2})$.
Proof.
Let us assume the original set $\bar{\bar{\mathsf{S}}}_{\upsilon}$ is associated with $M$ distinct vectors $(\delta^{m}P_{1},\delta^{m}P_{2})$, $m=1,\ldots,M$. Each of these $M$ vectors is associated with a respective set of code-books’ densities. Consider
$(\breve{\delta}P_{1},\breve{\delta}P_{2})=(\min_{m}\delta^{m}P_{1},\min_{m}\delta^{m}P_{2}).$ (98)
The pair $(\breve{\delta}P_{1},\breve{\delta}P_{2})$ is called the Pareto minimal point corresponding to the set $(\delta^{m}P_{1},\delta^{m}P_{2})$, $m=1,\ldots,M$. Let us use $(\breve{\delta}P_{1},\breve{\delta}P_{2})$ to compute new values for $(\mathbf{\Upsilon},\mathbf{\Gamma})$ and select the subset with smallest value of $\upsilon$ denoted as $\breve{\upsilon}$. Accordingly, let us form the sets $\bar{\mathsf{S}}_{\breve{\upsilon}}$ and $\bar{\bar{\mathsf{S}}}_{\breve{\upsilon}}$. Starting from the power reallocation vector $(\breve{\delta}P_{1},\breve{\delta}P_{2})$, each of the pairs $(\delta^{m}P_{1},\delta^{m}P_{2})$, $m=1,\ldots,M$, can be reached relying on a step with power reallocation $(\delta^{m}P_{1}-\breve{\delta}P_{1},\delta^{m}P_{2}-\breve{\delta}P_{2})$. This is possible since $\delta^{m}P_{1}-\breve{\delta}P_{1}\geq 0$ and $\delta^{m}P_{2}-\breve{\delta}P_{2}\geq 0$. This means relying on $\bar{\bar{\mathsf{S}}}_{\breve{\upsilon}}$ to achieve the next point on the boundary does not contradict the possibility of further moving counterclockwise to achieve the boundary point corresponding to $\bar{\bar{\mathsf{S}}}_{\upsilon},\penalty 10000\ \upsilon<\breve{\upsilon}$. Now let us shrink the range for power reallocation vector by setting
${\Delta}\!P_{1}=\breve{\delta}P_{1}\penalty 10000\ \penalty 10000\ \mbox{and}\penalty 10000\ \penalty 10000\ {\Delta}\!P_{2}=\breve{\delta}P_{2}.$ (99)
Accordingly, let us construct new sets following 95 and 96. Having multiple elements in $\bar{\bar{\mathsf{S}}}_{\breve{\upsilon}}$ allows recursively moving in clockwise direction, where $\mathbf{\Upsilon}$ increases and $\mathbf{\Gamma}$ decreases in each step. This procure can continue until one of the following cases occurs. Case (i): The value of $\mathbf{\Gamma}$ at the final point is zero. Case (ii): The final set includes a single Pareto minimal power reallocation vector achieving a single point on the boundary. Case (i) entails no further counterclockwise step along the boundary is feasible, requiring a change in the strategy. In Case (ii), from Theorems 1, 5 and 8, it follows that there is a Pareto minimal power reallocation which, in conjunction with zero-mean Gaussian code-books for compound random variables, results in a unique point on the boundary.
In summary, referring to Theorem 8, using $(\breve{\delta}P_{1},\breve{\delta}P_{2})$ instead of $(\delta^{m}P_{1},\delta^{m}P_{2})$, $m=1,\ldots,M$ is accompanied by a movement in clockwise direction, i.e., reaching from $(\upsilon,\gamma)$ to $(\breve{\upsilon},\breve{\gamma})$, where
$(\upsilon,\varkappa_{\upsilon})\leadsto(\breve{\upsilon},\varkappa_{\breve{\upsilon}}):\penalty 10000\ \penalty 10000\ \breve{\upsilon}>\upsilon\penalty 10000\ \penalty 10000\ \mbox{and}\penalty 10000\ \penalty 10000\ \varkappa_{\breve{\upsilon}}<\varkappa_{\upsilon}.$ (100)
Such a movement can continue in a recursive manner until the step size is small enough to include a single power reallocation vector, i.e.,
$\exists i\in[1,\ldots,M]:\penalty 10000\ (\breve{\delta}P_{1},\breve{\delta}P_{2})=(\delta^{i}P_{1},\delta^{i}P_{2})$ (101)
with the resulting $(\breve{\delta}P_{1},\breve{\delta}P_{2})$ achieving to a unique point on the boundary. Theorem 8 entails, relying on Pareto minimal power reallocation, the past history in moving counterclockwise along the boundary is captured solely by the starting point in each step. This means, considering two nested Pareto minimal power reallocation vectors $(\dot{\delta}P_{1},\dot{\delta}P_{2})$ and $(\ddot{\delta}P_{1},\ddot{\delta}P_{2})$, where
$\dot{\delta}P_{1}\leq\ddot{\delta}P_{1}\penalty 10000\ \penalty 10000\ \mbox{and}\penalty 10000\ \penalty 10000\ \dot{\delta}P_{2}\leq\ddot{\delta}P_{2}.$ (102)
These power reallocation vectors, in conjunction with Gaussian code-books, achieve two successive points on the boundary
$(\mathbf{\Upsilon}_{\!1},\mathbf{\Gamma}_{\!1})=(\dot{\upsilon},\varkappa_{\dot{\upsilon}})\penalty 10000\ \penalty 10000\ \mbox{and}\penalty 10000\ \penalty 10000\ (\mathbf{\Upsilon}_{\!2},\mathbf{\Gamma}_{\!2})=(\ddot{\upsilon},\varkappa_{\ddot{\upsilon}})$ (103)
satisfying
$\ddot{\upsilon}\leq\dot{\upsilon}\penalty 10000\ \penalty 10000\ \mbox{and}\penalty 10000\ \penalty 10000\ \varkappa_{\ddot{\upsilon}}\geq\varkappa_{\dot{\upsilon}}.$ (104)
Theorem 9.
Assume power reallocation vector is Pareto minimal. Then, the stationary solution obtained using Gaussian densities for core random variables results in an end-point which falls on the boundary (boundary corresponding to single letter code-books).
Proof.
Consider a power reallocation vector achieving a unique end point on the boundary. For such a power reallocation vector, consider applying calculus of variation to $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$ by perturbing densities of core random variables. Setting the derivatives of underlying functional to zero results in a system of equality constraints, which are satisfied if compound random variables are jointly Gaussian. Each compound random variable is a linear combination of core random variables, and linear expressions obtained using different sets of compound random variables are consistent with each other. Theorem 3 shows that the matrix of corresponding linear coefficients is invertible. This in turn means core random variables can be expressed as a unique linear combination of compound random variables. This means core random variables should be Gaussian, and the correspondence is unique. From Theorems 5 and 6, the stationary solution based on Gaussian densities for compound random variables either maximizes or minimizes $\mathbf{\Upsilon}$. A similar conclusion applies to $\mathbf{\Gamma}$. Combining these arguments with Theorem 1 and Theorem 7, it is concluded that for the Gaussian code-books in conjunction with a Pareto minimal power reallocation vector, $\mathbf{\Upsilon}$ is minimized while $\mathbf{\Gamma}$ is maximized. By applying calculus of variation directly to $\Delta\!R_{1}$ and $\Delta\!R_{2}$, it follows that Gaussian code-books, in conjunction with a Pareto minimal power reallocation vector, result in a stationary solution for $\Delta\!R_{1}$, $\Delta\!R_{2}$ with non-zero second order variation. ∎
Figure 5: $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$ as a function of $\omega$ (related to Theorem 10).
Remark 5: The optimum Pareto minimal power reallocation vector is not unique. However, the corresponding set has a nested structure, and relying on any member of the set will be associated with a unique set of Gaussian code-books (see Theorem 9). Different members of the set of Pareto minimal power reallocation pairs correspond to different step sizes. This property allows covering the boundary in a continuous manner. $\blacksquare$
Next, Theorem 10, in conjunction with Fig. 5, provides more details concerning the the stationary solutions based on Theorems 5 and 6. Consider a segment on the boundary from a starting point $\mathbb{S}$ to an end point $\mathbb{E}$ as depicted in Fig. 5. With some misuse of notation, superscripts are used to refer to points inside or on the boundary. Assume power reallocation vector for point $\mathbb{E}$ is equal to $({\delta}P^{\mathbb{E}}_{U_{1}},{\delta}P^{\mathbb{E}}_{V_{1}},{\delta}P^{\mathbb{E}}_{U_{2}},{\delta}P^{\mathbb{E}}_{V_{2}})$. Consider time sharing between points $\mathbb{S}$ and $\mathbb{E}$ with a time sharing factor $\omega\in[0,1]$ where $\omega=0$ and $\omega=1$ correspond to points $\mathbb{S}$ and $\mathbb{E}$, respectively. Time sharing achieves point $\mathbb{T}$ inside the capacity region, corresponding to an effective (time-shared) power reallocation vector $\omega({\delta}P^{\mathbb{E}}_{U_{1}},{\delta}^{\mathbb{E}}P_{V_{1}},{\delta}P^{\mathbb{E}}_{U_{2}},{\delta}P^{\mathbb{E}}_{V_{2}})$. Let us assume the power reallocation vector $\omega({\delta}P^{\mathbb{E}}_{U_{1}},{\delta}^{\mathbb{E}}P_{V_{1}},{\delta}P^{\mathbb{E}}_{U_{2}},{\delta}P^{\mathbb{E}}_{V_{2}})$, with optimum code-books’ densities, results in the point $\mathbb{B}$ on the boudnary corresponding to $\mathbf{\Upsilon}^{\mathbb{B}}(\omega)$, $\mathbf{\Gamma}^{\mathbb{B}}(\omega)$. Relying on code-books’ densities obtained through time sharing for point $\mathbb{T}$ and optimum code-books’ densities for points $\mathbb{E}$ and $\mathbb{B}$, we have
$\mathbf{\Upsilon}^{\mathbb{B}}$ $>$ $\mathbf{\Upsilon}^{\mathbb{T}}\penalty 10000\ =\penalty 10000\ \mathbf{\Upsilon}^{\mathbb{E}}$ (105)
$\mathbf{\Gamma}^{\mathbb{B}}$ $>$ $\mathbf{\Gamma}^{\mathbb{T}}\penalty 10000\ =\penalty 10000\ \omega\mathbf{\Gamma}^{\mathbb{E}}.$ (106)
Theorem 10.
Consider a segment [^6] This means the strategy remains the same from $\mathbb{S}$ to $\mathbb{E}$. of the boundary (limited to single-letter code-books) starting from a point $\mathbb{S}$ to a point $\mathbb{E}$. Power reallocation corresponding to $\omega({\delta}P^{\mathbb{E}}_{U_{1}},{\delta}^{\mathbb{E}}P_{V_{1}},{\delta}P^{\mathbb{E}}_{U_{2}},{\delta}P^{\mathbb{E}}_{V_{2}})$, in conjunction with independent and identical Gaussian code-books, solves the following constrained optimization problems:
Maximize $\penalty 10000\ \mathbf{\Gamma}(\omega)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{subject to}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathbf{\Upsilon}(\omega)\,\geq\hat{\mathbf{\Upsilon}}$ (107)
Maximize $\Delta\!R_{2}(\omega)\penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{subject to}\penalty 10000\ \penalty 10000\ \penalty 10000\ \Delta\!R_{1}(\omega)\leq\hat{\Delta}\!R_{1}$ (108)
Minimize $\Delta\!R_{1}(\omega)\penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{subject to}\penalty 10000\ \penalty 10000\ \penalty 10000\ \Delta\!R_{2}(\omega)\geq\hat{\Delta}\!R_{2}$ (109)
Maximize $\penalty 10000\ R_{2}(\omega)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{subject to}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ R_{1}(\omega)\geq\hat{R}_{1}$ (110)
Maximize $\penalty 10000\ R_{1}(\omega)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{subject to}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ R_{2}(\omega)\geq\hat{R}_{2}$ (111)
for given $\hat{\mathbf{\Upsilon}}$, $\hat{\Delta}\!R_{1}$, $\hat{\Delta}\!R_{2}$, $\hat{R}_{1}$ and $\hat{R}_{2}$.
Proof.
Noting Theorem 1, for independent and identical Gaussian code-books, we have
$\mathbf{\Gamma}\,(\omega)$ is a monotonically increasing function of $\omega$ (112)
$\penalty 10000\ \mathbf{\Upsilon}\,(\omega)$ is a monotonically decreasing function of $\omega.$ (113)
Noting 112 and 113, for power reallocation $\omega({\delta}P^{\mathbb{E}}_{U_{1}},{\delta}^{\mathbb{E}}P_{V_{1}},{\delta}P^{\mathbb{E}}_{U_{2}},{\delta}P^{\mathbb{E}}_{V_{2}})$, independent and identical Gaussian code-books result in
[…]
All discussions so far were limited to only one component GIC, where encoding and decoding procedures relied on single letter code-books. Since the single letter analysis did not impose any restrictions on $P_{1}$ and $P_{2}$, it follows that a simple time-sharing involving several single letter rate regions, equipped with power allocation (over messages/time), can be realized. Next, in Section 5, Theorem 11 shows that the maximum weighted sum-rate for each component GIC (with given $P_{1}$ and $P_{2}$) can be realized without the need to use vector encoding. Theorems 12 and 13 entail each point on the upper concave envelope can be achieved by time-sharing between (at most) two unique single letter Gaussian code-books, one corresponding to a two users GIC and the one, if it exists, to a single user Gaussian channel. As a result, the overall problem is simplified by bypassing the need to explicitly compute the optimum time sharing and the associated optimum power allocation over messages/time, i.e., computing the upper concave envelope.
5 Optimality of Single Letter Code-books
In time-sharing, time axis is divided into multiple non-overlapping segments, called phases hereafter. Each phase uses a fraction of time, a fraction of $P_{1}$ and a fraction of $P_{2}$, to maximize its relative contribution to the cumulative weighted sum-rate. Let us assume there are $\aleph$ phases indexed by $\mathtt{n}=1,\ldots,\aleph$ with time duration $\mathtt{t}_{1}\leq\mathtt{t}_{2}\leq\mathtt{t}_{3}\ldots\leq\mathtt{t}_{\aleph}$. To simplify arguments, phases are changed to pairs of equal duration; the first pair includes phase $\mathtt{n}=1$ and a part of the phase $\mathtt{n}=2$. Remaining phases, including what is left from phase $\mathtt{n}=2$, are ordered and pairing continues recursively. Let us focus on one such pair. Superscripts $(\,\bar{\cdot}\,)$ and $(\,\bar{\bar{\cdot}}\,)$ refer to the first phase and the second phase in the pair. Power of user 1 allocated to the two phases forming the pair are denoted as $\bar{\wp}_{{}_{1}}$ and $\bar{\bar{\wp}}_{{}_{1}}$. Likewise, power of user 2 allocated to the two phases are denoted as $\bar{\wp}_{{}_{2}}$ and $\bar{\bar{\wp}}_{{}_{2}}$. Notations $\mathtt{u}_{1},\mathtt{v}_{1},\mathtt{u}_{2},\mathtt{v}_{2}$, $\mathtt{x}_{1},\mathtt{x}_{2}$ refer to (vector) code-books and $\mathtt{y}_{1},\mathtt{y}_{2}$ to corresponding outputs. Components of a vector are indexed using a superscript, e.g., components of $\mathtt{y}_{1}$ are denoted as $y^{i}_{1}$. Corresponding rate values are denoted as $\bar{\mathfrak{R}}_{1}=\bar{\mathfrak{r}}_{\mathtt{u}_{1}}+\bar{\mathfrak{r}}_{\mathtt{v}_{1}}$, $\bar{\mathfrak{R}}_{2}=\bar{\mathfrak{r}}_{\mathtt{u}_{2}}+\bar{\mathfrak{r}}_{\mathtt{v}_{2}}$ and $\bar{\bar{\mathfrak{R}}}_{1}=\bar{\bar{\mathfrak{r}}}_{\mathtt{u}_{1}}+\bar{\bar{\mathfrak{r}}}_{\mathtt{v}_{1}}$, $\bar{\bar{\mathfrak{R}}}_{2}=\bar{\bar{\mathfrak{r}}}_{\mathtt{u}_{2}}+\bar{\bar{\mathfrak{r}}}_{\mathtt{v}_{2}}$.
For a block length $\mathtt{t}$, consider vector code-books $\mathtt{u_{1}}=[u_{1}^{i},i=1,\ldots\mathtt{t}]$, $\mathtt{v_{1}}=[v_{1}^{i},i=1,\ldots\mathtt{t}]$, $\mathtt{u_{2}}=[u_{2}^{i},i=1,\ldots\mathtt{t}]$, $\mathtt{v_{2}}=[v_{2}^{i},i=1,\ldots\mathtt{t}]$ generated using densities $\mathfrak{p}(\mathtt{u}_{1})$, $\mathfrak{p}(\mathtt{v}_{1})$, $\mathfrak{p}(\mathtt{u}_{2})$ and $\mathfrak{p}(\mathtt{v}_{2})$, with marginals $p_{i}(u_{1}^{i}),p_{i}(v_{1}^{i}),p_{i}(u_{2}^{i}),p_{i}(v_{2}^{i}),i=1,\ldots\mathtt{t}$, respectively.
Consider perturbations in 115 to 120:
$\mathfrak{p}(\mathtt{u}_{1})$ $\rightarrow$ $\mathfrak{p}(\mathtt{u}_{1})+\epsilon_{\mathtt{u}_{1}}\mathfrak{h}(\mathtt{u}_{1})$ (115)
$\mathfrak{p}(\mathtt{v}_{1})$ $\rightarrow$ $\mathfrak{p}(\mathtt{v}_{1})+\epsilon_{\mathtt{v}_{1}}\mathfrak{h}(\mathtt{v}_{1})$ (116)
$\mathfrak{p}(\mathtt{u}_{2})$ $\rightarrow$ $\mathfrak{p}(\mathtt{u}_{2})+\epsilon_{\mathtt{u}_{2}}\mathfrak{h}(\mathtt{u}_{2})$ (117)
$\mathfrak{p}(\mathtt{v}_{2})$ $\rightarrow$ $\mathfrak{p}(\mathtt{v}_{2})+\epsilon_{\mathtt{v}_{2}}\mathfrak{h}(\mathtt{v}_{2})$ (118)
subject to: (a) power constraints
$E(\|\mathtt{u}_{1}\|^{2})+E(\|\mathtt{v}_{1}\|^{2})\,=\,\mathtt{t}P_{1}$ (119)
$E(\|\mathtt{u}_{2}\|^{2})+E(\|\mathtt{v}_{2}\|^{2})\,=\,\mathtt{t}P_{2}$ (120)
and (b) integration of densities $\mathfrak{p}(\mathtt{u}_{1}),\mathfrak{p}(\mathtt{v}_{1}),\mathfrak{p}(\mathtt{u}_{2}),\mathfrak{p}(\mathtt{v}_{2})$ over their respective domains. Also, consider perturbations in 121 to 126:
$\mathfrak{p}(\mathtt{u}_{1})=\prod_{i=1}^{\mathtt{t}}p_{i}(u_{1}^{i})$ $\rightarrow$ $\prod_{i=1}^{\mathtt{t}}p_{i}(u_{1}^{i})+\epsilon_{\mathtt{u}_{1}}\mathfrak{h}(\mathtt{u}_{1})$ (121)
$\mathfrak{p}(\mathtt{v}_{1})=\prod_{i=1}^{\mathtt{t}}p_{i}(v_{1}^{i})$ $\rightarrow$ $\prod_{i=1}^{\mathtt{t}}p_{i}(v_{1}^{i})+\epsilon_{\mathtt{v}_{1}}\mathfrak{h}(\mathtt{v}_{1})$ (122)
$\mathfrak{p}(\mathtt{u}_{2})=\prod_{i=1}^{\mathtt{t}}p_{i}(u_{2}^{i})$ $\rightarrow$ $\prod_{i=1}^{\mathtt{t}}p_{i}(u_{2}^{i})+\epsilon_{\mathtt{u}_{2}}\mathfrak{h}(\mathtt{u}_{2})$ (123)
$\mathfrak{p}(\mathtt{v}_{2})=\prod_{i=1}^{\mathtt{t}}p_{i}(v_{2}^{i})$ $\rightarrow$ $\prod_{i=1}^{\mathtt{t}}p_{i}(v_{2}^{i})+\epsilon_{\mathtt{v}_{2}}\mathfrak{h}(\mathtt{v}_{2}).$ (124)
subject to: (a) power constraints
$\sum_{i=1}^{\mathtt{t}}E(|u_{1}^{i}|^{2})\,+\,\sum_{i=1}^{\mathtt{t}}E(|v_{1}^{i}|^{2})\,=\,\mathtt{t}P_{1}$ (125)
$\sum_{i=1}^{\mathtt{t}}E(|u_{2}^{i}|^{2})\,+\,\sum_{i=1}^{\mathtt{t}}E(|v_{2}^{i}|^{2})\,=\,\mathtt{t}P_{2}$ (126)
and (b) area under densities $p_{i}(u_{1}^{i}),p_{i}(v_{1}^{i}),p_{i}(u_{2}^{i}),p_{i}(v_{2}^{i}),i=1,\ldots\mathtt{t}$. Relying on calculus of variation, it follows that independent and identical Gaussian densities provide a stationary solution for 115 to 120, as well as for 121 to 126, both with a non-zero second order variation.
Theorem 11.
Consider a phase where both users are active. Independent and identically distributed Gaussian code-books (single letter) optimize the weighted sum-rate.
Proof.
Let us assume there is a segment on the boundary where vector encoding over a phase (where both users are active) is strictly superior to independent and identical Gaussian code-books, as shown in Fig. 6. Consider a point $\mathbb{A}_{1}$ on the vector-coded curve (v) corresponding to a weight $\mu_{1}$. Point $\mathbb{G}_{1}$ on the Gaussian-coded curve (g) corresponds to the same value of $\mu_{1}$. Let us move from point $\mathbb{G}_{1}$ to a point $\mathbb{A}_{2}$ by increasing $\mathfrak{R}_{2}$, corresponding to the value of $\mu_{2}<\mu_{1}$, which in turn maps to a point $\mathbb{G}_{2}$ on the Gaussian curve. This procedure of moving clock-wise along the two curves can continue until reaching to points $\mathbb{A}_{n}$, $\mathbb{G}_{n}$ and $\mathbb{G}_{n+1}$, such that: (a) $\mathbb{G}_{n+1}$ falls on the end point of the previous segment, (b) $\mathbb{G}_{n+1}$ and $\mathbb{G}_{n}$ are in a neighborhood formed by an infinitesimal boundary achieving power reallocation, and (c) $\mathbb{A}_{n}$ and $\mathbb{G}_{n+1}$ form a neighborhood due to perturbations in 115 to 120 at $\mathbb{A}_{n}$, and/or perturbations in 121 to 126 at $\mathbb{G}_{n+1}$ . Note that: (d) $\mathbb{G}_{n}$ and $\mathbb{G}_{n+1}$ remain separated due to the non-zero power reallocation, (e) $\mathbb{A}_{n}$ and $\mathbb{G}_{n}$ remain separated, since $\mathbb{A}_{n}$ is obtained from $\mathbb{G}_{n}$ by a small, but non-zero, increase in $\mathfrak{R}_{2}$, and (f) $\mathbb{G}_{n+1}$ and $\mathbb{A}_{n}$ correspond to the same value of weight $\mu_{n}$. Considering formation of these continuous neighborhoods, it follows that the value of weighted sum-rate at the point $\mathbb{A}_{n}$ on the boundary approaches that of the (potentially) local optimum, single letter, solution $\mathbb{G}_{n+1}$.
Figure 6: Segment of the boundary capturing the weighted sum-rate over the two-user phase where vector encoding captured by the curve denoted as (v) is strictly superior to independent and identical Gaussian code-books captured by the curve denoted as (g). The curve denoted by (v) would be used in optimizing the upper concave envelope. Consequently, (potential) improvements due to vector encoding is studied without the need to include time sharing.
Theorem 12.
Consider two phases over which both users are active. An optimum solution exists for which the two phases can be merged into one.
Proof.
Without loss of generality, (a) consider a pair of phases of equal duration where both users are active, and (b) limit the analysis to using single letter code-books over each phase. Consider a solution where power levels $\bar{\wp}_{{}_{1}},\bar{\bar{\wp}}_{{}_{1}},\bar{\wp}_{{}_{2}},\bar{\bar{\wp}}_{{}_{2}}$ are strictly positive, $\bar{\wp}_{{}_{1}}\neq\bar{\bar{\wp}}_{{}_{1}}$ and/or $\bar{\wp}_{{}_{2}}\neq\bar{\bar{\wp}}_{{}_{2}}$. Consider applying perturbations in 115 to 120 to find a stationary solution for the cumulative weighted sum-rate, i.e., $\bar{\mathfrak{R}}_{1}+\mu\bar{\mathfrak{R}}_{2}+\bar{\bar{\mathfrak{R}}}_{1}+\mu\bar{\bar{\mathfrak{R}}}_{2}$ subject to $\bar{\wp}_{{}_{1}}+\bar{\bar{\wp}}_{{}_{1}}$ for the total power of user 1 and $\bar{\wp}_{{}_{2}}+\bar{\bar{\wp}}_{{}_{2}}$ for the total power of user 2, respectively. It follows that independent and identical Gaussian densities, with uniform power allocation over messages, as well as over time, results in a stationary solution which, noting Theorem 11, is the global optimum solution to formulations in Theorem 10. ∎
Theorem 13.
Assume the optimum solution includes a phase where both users are active. There is at most one additional phase over which a single user is active.
Proof.
Consider a phase 1 where both users are active, composed of $\bar{\mathtt{t}}_{[1,2]}>0$ time samples. The statement of theorem fails if, in addition to phase 1, there are two single-user phases, 2 and 3, of duration $\bar{\bar{\mathtt{t}}}_{[1]}\geq 0$ and $\bar{\bar{\bar{\mathtt{t}}}}_{[2]}\geq 0$, occupied by users 1 and 2, respectively. Constraint on total time is expressed as $\mathtt{T}=\bar{\mathtt{t}}_{[1,2]}+\bar{\bar{\mathtt{t}}}_{[1]}+\bar{\bar{\bar{\mathtt{t}}}}_{[2]}$ where $\mathtt{T}$ is fixed. Let us rely on notations $p_{1}\geq 0$ and $p_{2}\geq 0$ to refer to the power values allocated, from power budgets of user 1 and user 2, to phases 2 and 3, respectively. This means power values allocated to phase 1 are equal to $P_{1}-p_{1}$ and $P_{2}-p_{2}$. We have: (a) Weighted sum-rate corresponding to phase 1 is monotonically decreasing function of $p_{1}$, $p_{2}$, $\bar{\bar{\mathtt{t}}}_{[1]}$ and $\bar{\bar{\bar{\mathtt{t}}}}_{[2]}$. (b) Rate corresponding to phase 2 is a monotonically increasing function of $p_{1}$ and $\bar{\bar{\mathtt{t}}}_{[1]}$. Likewise, rate corresponding to phase 3 is a monotonically increasing function of $p_{2}$ and $\bar{\bar{\bar{\mathtt{t}}}}_{[2]}$. The boundary of the feasible region is formed by the following planes: Power Planes: $p_{1}=0$ and/or $p_{2}=0$. Time Planes: $\bar{\bar{\mathtt{t}}}_{[1]}=0$ and/or $\bar{\bar{\bar{\mathtt{t}}}}_{[2]}=0$. Since it is assumed $\bar{\mathtt{t}}_{[1,2]}>0$, the cases of $p_{1}=P_{1}$ and/or $p_{2}=P_{2}$ are not included in forming the boundary. The rate of user 1 over phases 1 and 2 add to $R_{1}$, and the rate of user 2 over phases 1 and 3 add to $R_{2}$. Noting (a) and (b) above, $R_{1}$, $R_{2}$, and consequently, $R_{1}+\mu R_{2}$ for a given $\mu$, change monotonically over the feasible region. Noting $\bar{\mathtt{t}}_{[1,2]}>0$, Power Planes and Time Planes coincide at a single point formed by $(\bar{\bar{\mathtt{t}}}_{[1]},p_{1})=(0,0)$ and/or $(\bar{\bar{\bar{\mathtt{t}}}}_{[2]},p_{2})=(0,0)$. Noting $R_{1}+\mu R_{2}$ is monotonically changing over the feasible region, there is an optimum solution which falls on the boundary corresponding to either $(\bar{\bar{\mathtt{t}}}_{[1]},p_{1})=(0,0)$ and/or $(\bar{\bar{\bar{\mathtt{t}}}}_{[2]}),p_{2})=(0,0)$. ∎
Remark 6: Theorem 12 can be applied to merge any number of phases over which both users are active into a single two-user phase. The phase occupied by a single user, if it exists, corresponds to a simple point-to-point Gaussian noise channel, for which single-letter Gaussian code-book maximizes the rate. Noting a zero mean Gaussian density is entirely specified by its second moment, Theorems 11, 12 and 13 entail any point on the upper concave envelope can be achieved by time-sharing between (at most) two single letter Gaussian code-books, corresponding to one two-user phase and (at most) one single-user phase. $\blacksquare$
6 Converse Results
Expressing 35, 36, 37 in terms of vectors, we have
$\mathfrak{r}_{\mathtt{u}_{1}}$ $=$ $I(\mathtt{u}_{1};\mathtt{y}_{2}|\mathtt{u}_{2})$ (127)
$\mathfrak{r}_{\mathtt{u}_{2}}$ $=$ $I(\mathtt{u}_{2};\mathtt{y}_{2})$ (128)
$\mathfrak{r}_{\mathtt{u}_{1}}+\mathfrak{r}_{\mathtt{u}_{2}}$ $=$ $I(\mathtt{u}_{1},\mathtt{u}_{2};\mathtt{y}_{1})=I(\mathtt{u}_{2};\mathtt{y}_{2})+I(\mathtt{u}_{1};\mathtt{y}_{2}|\mathtt{u}_{2})$ (129)
$\mathfrak{r}_{\mathtt{v}_{1}}$ $=$ $I(\mathtt{v}_{1};\mathtt{y}_{1}|\mathtt{u}_{1},\mathtt{u}_{2})$ (130)
$\mathfrak{r}_{\mathtt{v}_{2}}$ $=$ $I(\mathtt{v}_{2};\mathtt{y}_{2}|\mathtt{u}_{1},\mathtt{u}_{2})$ (131)
$\mathfrak{R}_{1}$ $=$ $I(\mathtt{u}_{1};\mathtt{y}_{2}|\mathtt{u}_{2})+I(\mathtt{v}_{1};\mathtt{y}_{1}|\mathtt{u}_{1},\mathtt{u}_{2})$ (132)
$\mathfrak{R}_{2}$ $=$ $I(\mathtt{u}_{2};\mathtt{y}_{2})+I(\mathtt{v}_{2};\mathtt{y}_{2}|\mathtt{u}_{1},\mathtt{u}_{2}).$ (133)
Theorem 14.
If probability of error in recovering $\mathtt{u}_{1},\mathtt{u}_{2},\mathtt{v}_{1}$ at $\mathtt{y}_{1}$ and $\mathtt{u}_{1},\mathtt{u}_{2},\mathtt{v}_{2}$ at $\mathtt{y}_{2}$ tend to zero as $\mathtt{t}\rightarrow\infty$, then the rate vector $(\mathfrak{R}_{1}=\mathfrak{r}_{\mathtt{u}_{1}}+\mathfrak{r}_{\mathtt{v}_{1}},\mathfrak{R}_{2}=\mathfrak{r}_{\mathtt{u}_{2}}+\mathfrak{r}_{\mathtt{v}_{2}})$ should fall within the optimum region with independent and identically distributed single letter Gaussian code-books.
Proof.
Consider a rate pair $\mathfrak{R}_{1},\mathfrak{R}_{2}$ satisfying 127 to 133, and rate pairs $(\mathfrak{R}_{1}+\mathfrak{\Delta},\mathfrak{R}_{2})$ and $(\mathfrak{R}_{1},\mathfrak{R}_{2}+\mathfrak{\Delta})$ with $\mathfrak{\Delta}>0$. Let us consider the channel models in Figs. 3 in conjunction with vectors $\mathtt{u}_{1},\mathtt{v}_{1},\mathtt{u}_{2},\mathtt{v}_{2}$. In Fig. 3, the input $\mathtt{u}_{2}$ is subject to an additive noise term, and consequently, if $\mathfrak{r}_{\mathtt{u}_{2}}>I(\mathtt{u}_{2};\mathtt{y}_{2})$, the probability of error for $\mathtt{u}_{2}$, denoted as $P^{\,e}_{\mathtt{u}_{2}}$, will be bounded away from zero, i.e., $P^{\,e}_{\mathtt{u}_{2}}\nrightarrow 0$. If $\mathfrak{r}_{\mathtt{u}_{2}}>I(\mathtt{u}_{2};\mathtt{y}_{2})$, noting 127 to 133, $\mathtt{u}_{2}$ acts as noise in: (a) decoding of $\mathtt{u}_{1}$ at $\mathtt{y}_{2}$; see 127, (b) joint decoding of $(\mathtt{u}_{1},\mathtt{u}_{1})$ at $\mathtt{y}_{1}$; see 129, (c) decoding of $\mathtt{v}_{1}$ at $\mathtt{y}_{1}$; see 130, and (d) decoding of $\mathtt{v}_{2}$ at $\mathtt{y}_{2}$; see 131. In such a case, even if rates $\mathfrak{r}_{\mathtt{u}_{1}}$, $\mathfrak{r}_{\mathtt{v}_{1}}$, $\mathfrak{r}_{\mathtt{v}_{2}}$ remain limited to mutual information terms over their respective channels in Figs. 3, we have $P^{\,e}_{\mathtt{u}_{1}}\nrightarrow 0$, $P^{\,e}_{\mathtt{v}_{1}}\nrightarrow 0$ and $P^{\,e}_{\mathtt{v}_{2}}\nrightarrow 0$. On the other hand, if $\mathfrak{r}_{\mathtt{u}_{2}}=I(\mathtt{u}_{2};\mathtt{y}_{2})$, then $\mathtt{u}_{2}$ can be decoded at $\mathtt{y}_{2}$. In this case, if $\mathfrak{r}_{\mathtt{u}_{1}}>I(\sqrt{a}\mathtt{u}_{1};\mathtt{y}_{2}|\mathtt{u}_{2})$, noting the channel in Fig. 3, we conclude $P^{\,e}_{\mathtt{u}_{1}}\nrightarrow 0$, $P^{\,e}_{\mathtt{v}_{2}}\nrightarrow 0$. Likewise, for the channel in Fig. 3, $P^{\,e}_{(\mathtt{u}_{1},\mathtt{u}_{2})}\nrightarrow 0$, $P^{\,e}_{\mathtt{v}_{1}}\nrightarrow 0$. Next consider the case that the rates of public messages are limited to their respective mutual information terms, and hence public messages can be decoded error-free at $\mathtt{y}_{1}$ and $\mathtt{y}_{2}$. Removing $\mathtt{u}_{1}$, $\mathtt{u}_{2}$ from channel models in Figs. 3, it follows that $P^{\,e}_{\mathtt{v}_{1}}\nrightarrow 0$ unless $R_{\mathtt{v}_{1}}\leq I(\mathtt{v}_{1};\mathtt{y}_{1}|\mathtt{u}_{1},\mathtt{u}_{2})$ and $P^{\,e}_{\mathtt{v}_{2}}\nrightarrow 0$ unless $R_{\mathtt{v}_{2}}\leq I(\mathtt{v}_{2};\mathtt{y}_{2}|\mathtt{u}_{1},\mathtt{u}_{2})$. The proof follows noting the region formed based on 127 to 133 is maximally enlarged using independent and identically distributed single letter Gaussian code-books (see Theorem 11). ∎
Next, it will be shown that the Han-Kobayashi (HK) achievable rate region, upon shrinking its feasible region by imposing some restrictive but consistent constraints, achieves the boundary of the capacity region.
7 Optimality of the HK Region with Gaussian Code-books
Noting Theorem 12, let us limit our attention to single letter case. Expanded Han-Kobayashi constraints [^7] See expressions 3.2 to 3.15 on page 51 of [5], with the changes ($\mbox{current\penalty 10000\ article}\leftrightarrow[2]$): $U_{1}\leftrightarrow W_{1}$, $U_{2}\leftrightarrow W_{2}$, $V_{1}\leftrightarrow U_{1}$, $V_{2}\leftrightarrow U_{2}$, $R_{U_{1}}\leftrightarrow T_{1}$, $R_{U_{2}}\leftrightarrow T_{2}$, $R_{V_{1}}\leftrightarrow S_{1}$, $R_{V_{2}}\leftrightarrow S_{2}$. can be expressed as [5],
Maximize: $R_{1}+\mu R_{2}\penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{where}$ (134)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{1}|U_{2},V_{1})$ (135)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{2}|U_{2},V_{2})$ (136)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{1}|U_{1},V_{1})$ (137)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{2}|U_{1},V_{2})$ (138)
$R_{V_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(V_{1};Y_{1}|U_{1},U_{2})$ (139)
$R_{V_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(V_{2};Y_{2}|U_{1},U_{2})$ (140)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{1}|V_{1})$ (141)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{2}|V_{2})$ (142)
$R_{U_{1}}+R_{V_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1},V_{1};Y_{1}|U_{2})=I(U_{1};Y_{1}|U_{2})+I(V_{1};Y_{1}|U_{1},U_{2})$ (143)
$R_{U_{2}}+R_{V_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{2},V_{2};Y_{2}|U_{1})=I(U_{2};Y_{2}|U_{1})+I(V_{2};Y_{2}|U_{1},U_{2})$ (144)
$R_{U_{2}}+R_{V_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{2},V_{1};Y_{1}|U_{1})=I(U_{2};Y_{1}|U_{1})+I(V_{1};Y_{1}|U_{1},U_{2})$ (145)
$R_{U_{1}}+R_{V_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1},V_{2};Y_{2}|U_{2})=I(U_{1};Y_{2}|U_{2})+I(V_{2};Y_{2}|U_{1},U_{2})$ (146)
$R_{U_{1}}+R_{U_{2}}+R_{V_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2},V_{1};Y_{1})=I(U_{1},U_{2};Y_{1})+I(V_{1};Y_{1}|U_{1},U_{2})$ (147)
$R_{U_{1}}+R_{U_{2}}+R_{V_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2},V_{2};Y_{2})=I(U_{1},U_{2};Y_{2})+I(V_{2};Y_{2}|U_{1},U_{2})$ (148)
$E(X_{1}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{1}$ (149)
$E(X_{2}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{2}.$ (150)
Since the above formulation results in an achievable weighted sum-rate, any set of restrictive assumptions, if consistent with 134 to 150, results in an achievable (potentially inferior) solution. Let us restrict $U_{1}$, $U_{2}$, $V_{1}$, $V_{2}$ to be independent, $X_{1}=U_{1}+V_{1}$, $X_{2}=U_{2}+V_{2}$. We have $E(X_{1}^{2})=E(U_{1}^{2})+E(V_{1}^{2})$ and $E(X_{2}^{2})=E(U_{2}^{2})+E(V_{2}^{2})$. For given power allocation and encoding/decoding strategies (determining the values of mutual information terms on right hand sides of 135 to 148), optimization problem in 134 to 148 will be a parametric linear programming problem with four variables, i.e., $R_{U_{1}}$, $R_{U_{2}}$, $R_{V_{1}}$, $R_{V_{2}}$. This means, in the optimum solution, at least 4 constraints among 135 to 148 will be satisfied with equality, resulting in zero value for the corresponding slack variables. It turns out, with optimized power allocation and encoding/decoding strategies, a higher number of slack variables will be zero. In view of the dual linear program, these additional zero-valued slack variables will be advantageous in increasing the value of the objective function.
Let us shrink the HK region by restrictive assumptions
$R_{V_{1}}=I(V_{1};Y_{1}|U_{1},U_{2}),\penalty 10000\ \penalty 10000\ \penalty 10000\ R_{V_{2}}=I(V_{2};Y_{2}|U_{1},U_{2}),\penalty 10000\ \penalty 10000\ \penalty 10000\ R_{U_{1}}+R_{U_{2}}=I(U_{1},U_{2};Y_{1})=I(U_{1},U_{2};Y_{2}).$ (151)
We have
Maximize: $R_{1}+\mu R_{2}\penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{where}$ (152)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{1}|U_{2})\penalty 10000\ \stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\penalty 10000\ I(U_{1};Y_{1}|U_{2},V_{1})$ (153)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{2}|U_{2})\penalty 10000\ \stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\penalty 10000\ I(U_{1};Y_{2}|U_{2},V_{2})$ (154)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{1}|U_{1})\penalty 10000\ \stackrel{{\scriptstyle\text{(c)}}}{{\leq}}\penalty 10000\ I(U_{2};Y_{1}|U_{1},V_{1})$ (155)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{2}|U_{1})\penalty 10000\ \stackrel{{\scriptstyle\text{(d)}}}{{\leq}}\penalty 10000\ I(U_{2};Y_{2}|U_{1},V_{2})$ (156)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ \stackrel{{\scriptstyle\text{(e)}}}{{=}}\penalty 10000\ I(U_{1},U_{2};Y_{1})$ (157)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ \stackrel{{\scriptstyle\text{(f)}}}{{=}}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{2})$ (158)
$R_{V_{1}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(V_{1};Y_{1}|U_{1},U_{2})$ (159)
$R_{V_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(V_{2};Y_{2}|U_{1},U_{2})$ (160)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ \stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{1}|U_{2},V_{1})$ (161)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ \stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{2}|U_{2},V_{2})$ (162)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ \stackrel{{\scriptstyle\text{(c)}}}{{\leq}}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{1}|U_{1},V_{1})$ (163)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ \stackrel{{\scriptstyle\text{(d)}}}{{\leq}}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{2}|U_{1},V_{2})$ (164)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{1}|V_{1})\penalty 10000\ \stackrel{{\scriptstyle\text{(e)}}}{{\leq}}\penalty 10000\ I(U_{1},U_{2};Y_{1})$ (165)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{2}|V_{2})\penalty 10000\ \stackrel{{\scriptstyle\text{(f)}}}{{\leq}}\penalty 10000\ I(U_{1},U_{2};Y_{2})$ (166)
$E(X_{1}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{1}$ (167)
$E(X_{2}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{2}.$ (168)
Noting relationships specified by (a),(b),(c),(d),(e) and (f) in 152 to 166, it follows that 161 to 166 are redundant. Upon removing redundant constraints from 152 to 168, we obtain
Maximize: $R_{1}+\mu R_{2}\penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{where}$ (169)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{1}|U_{2})$ (170)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{2}|U_{2})$ (171)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{1}|U_{1})$ (172)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{2}|U_{1})$ (173)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{1})$ (174)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{2})$ (175)
$R_{V_{1}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(V_{1};Y_{1}|U_{1},U_{2})$ (176)
$R_{V_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(V_{2};Y_{2}|U_{1},U_{2})$ (177)
$E(X_{1}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{1}$ (178)
$E(X_{2}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{2}.$ (179)
Let us consider the following two problems with solutions which are potentially inferior to that of the original problem in 134 to 150.
Maximize: $R_{1}+\mu R_{2}\penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{where}$ (180)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{1}|U_{2})$ (181)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{2}|U_{2})$ (182)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{1}|U_{1})$ (183)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{2})\penalty 10000\ \leq\penalty 10000\ I(U_{2};Y_{2}|U_{1})$ (184)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{1})$ (185)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{2})$ (186)
$R_{V_{1}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(V_{1};Y_{1}|U_{1},U_{2})$ (187)
$R_{V_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(V_{2};Y_{2}|U_{1},U_{2})$ (188)
$E(X_{1}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{1}$ (189)
$E(X_{2}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{2}$ (190)
and
Maximize: $R_{1}+\mu R_{2}\penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{where}$ (191)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{1})\penalty 10000\ \leq\penalty 10000\ I(U_{1};Y_{1}|U_{2})$ (192)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{2}|U_{2})$ (193)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{1}|U_{1})$ (194)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{2}|U_{1})$ (195)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{1})$ (196)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{2})$ (197)
$R_{V_{1}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(V_{1};Y_{1}|U_{1},U_{2})$ (198)
$R_{V_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(V_{2};Y_{2}|U_{1},U_{2})$ (199)
$E(X_{1}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{1}$ (200)
$E(X_{2}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{2}$ (201)
The problem in 180 to 190 becomes the same as the one in 191 to 201 by swapping $U_{1}\longleftrightarrow U_{2}$, $V_{1}\longleftrightarrow V_{2}$. This means one of the two results in a higher value for $R_{1}+\mu R_{2}$ with $\mu<1$ and the other in a higher value for $R_{1}+\mu R_{2}$ for $\mu>1$. Let us focus on 180 to 190 and set
$I(U_{1};Y_{2}|U_{2})$ $\penalty 10000\ \penalty 10000\ \stackrel{{\scriptstyle\text{(e)}}}{{\leq}}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{1}|U_{2})$ (202)
$I(U_{2};Y_{1}|U_{1})$ $\penalty 10000\ \penalty 10000\ \stackrel{{\scriptstyle\text{(g)}}}{{\leq}}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{2}|U_{1}).$ (203)
This results in
Maximize: $R_{1}+\mu R_{2}\penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{where}$ (204)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{2}|U_{2})\penalty 10000\ \stackrel{{\scriptstyle\text{(e)}}}{{\leq}}\penalty 10000\ I(U_{1};Y_{1}|U_{2})$ (205)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{1}|U_{2})$ (206)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {\leq}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{1}|U_{1})\penalty 10000\ \stackrel{{\scriptstyle\text{(g)}}}{{\leq}}\penalty 10000\ I(U_{2};Y_{2}|U_{1})$ (207)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{2})\penalty 10000\ \leq\penalty 10000\ I(U_{2};Y_{2}|U_{1})$ (208)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{1})$ (209)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{2})$ (210)
$R_{V_{1}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(V_{1};Y_{1}|U_{1},U_{2})$ (211)
$R_{V_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(V_{2};Y_{2}|U_{1},U_{2})$ (212)
$E(X_{1}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{1}$ (213)
$E(X_{2}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{2}$ (214)
where 205, 207 are from 202 and 203, respectively. Removing redundant constraints from 204 to 214, we obtain
Maximize: $R_{1}+\mu R_{2}\penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{where}$ (215)
$R_{U_{1}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1};Y_{2}|U_{2})$ (216)
$R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{2};Y_{2})$ (217)
$R_{U_{1}}+R_{U_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(U_{1},U_{2};Y_{2})\penalty 10000\ =\penalty 10000\ I(U_{1},U_{2};Y_{1})$ (218)
$R_{V_{1}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(V_{1};Y_{1}|U_{1},U_{2})$ (219)
$R_{V_{2}}$ $\penalty 10000\ \penalty 10000\ {=}\penalty 10000\ \penalty 10000\ I(V_{2};Y_{2}|U_{1},U_{2})$ (220)
$E(X_{1}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{1}$ (221)
$E(X_{2}^{2})$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ P_{2}$ (222)
Solution to 215 to 222 results: (1) an achievable solution for which constraints in 134 to 150 are not violated, and (2) the corresponding solution coincides with optimum boundary established in Section 4.2 for $\mu<1$. This entails Han-Kobayashi region with Gaussian code-books is optimum.
Note that the formulation in 215 to 222 corresponds to the case that both users have public and private messages. For $\mu<1$, boundary includes segments where user 1 sends only a private message and user 2 sends both public and private messages. Likewise, for $\mu>1$, boundary includes segments where user 2 sends only a private message and user 1 sends both public and private messages. Formulations and proofs of optimality for these cases follow similarly.
Appendix
In the following, to simplify expressions, entropy values are computed in base “e".
Appendix A Constrained Maximization of Entropy Functions
A.1 Entropy Term Involving a Single Density Function
Consider the following constrained optimization problem:
$\mbox{Find function}\penalty 10000\ \penalty 10000\ f>0\penalty 10000\ \penalty 10000\ \mbox{to maximize}$ $\displaystyle-\int f\log(f)$ (223)
subject to: $\displaystyle\int x^{2}f=P$ (224)
and $\displaystyle\int f=1.$ (225)
Using Lagrange multipliers to add 224 and 225 to 223, we obtain
$\int f\log(f)+\lambda\int x^{2}f+\gamma\int f.$ (226)
Using a perturbation term $\epsilon h$ in 226 results in
$\int(f+\epsilon h)\log(f+\epsilon h)+\lambda\int x^{2}(f+\epsilon h)+\gamma\int f+\epsilon h.$ (227)
Derivative of 227 with respect to $\epsilon$ is equal to
$\int h\left[\log(f+\epsilon h)+1-\lambda x^{2}-\gamma\right].$ (228)
Setting 228 to zero for $\epsilon=0$, it follows that a Gaussian density for $f$ results in a stationary solution for constrained optimization problem in 223, 224 and 225. Next, it is shown that such a stationary solution is the maximum by using second order perturbation. Derivative of 228 with respect to $\epsilon$ at $\epsilon=0$ is equal to
[…]
Referring to reference [1], the condition in 229 implies that Gaussian density for $f$, computed relying on calculus of variation, is the global maximum solution for optimization problem in 223, 224 and 225.
A.2 Entropy Term Involving a Convolution of Density Functions
[…]
$\digamma=f_{1}\ast f_{2}.$ (230)
[…]
$\mathsf{H}^{\digamma}=-\int{\digamma}\ln({\digamma}).$ (231)
[…]
$p\digamma=(f_{1}+\epsilon_{1}h_{1})\ast(f_{2}+\epsilon_{2}h_{2})$ (232)
with an entropy of
$\mathsf{H}^{p\digamma}=-\int(f_{1}+\epsilon_{1}h_{1})\ast(f_{2}+\epsilon_{2}h_{2})\ln[(f_{1}+\epsilon_{1}h_{1})\ast(f_{2}+\epsilon_{2}h_{2})].$ (233)
[…]
$\frac{\partial\mathsf{H}^{p\digamma}}{\partial\epsilon_{1}}\Big|_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $0$ (234)
$\frac{\partial\mathsf{H}^{p\digamma}}{\partial\epsilon_{2}}\Big|_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $0$ (235)
$\frac{\partial\mathsf{H}^{p\digamma}}{\partial\epsilon_{1}}\Big|_{\epsilon_{1}=0,\epsilon_{2}=0}=-\int(h_{1}\ast f_{2})\ln(f_{1}\ast f_{2})-\int(h_{1}\ast f_{2})$ $=$ (236)
$-\int(h_{1}\ast f_{2})[\ln(f_{1}\ast f_{2})+1].$ (237)
Constraints on power and probability density function are expressed as:
$\mathsf{E}^{f_{1}\ast f_{2}}$ $=$ $\int x^{2}[f_{1}(x)\ast f_{2}(x)]dx\penalty 10000\ \penalty 10000\ \mbox{is a constant}$ (238)
$\mathsf{A}^{f_{1}\ast f_{2}}$ $=$ $\int f_{1}(x)\ast f_{2}(x)dx=1.$ (239)
We have
$\frac{\partial\mathsf{E}^{f_{1}\ast f_{2}}}{\partial\epsilon_{1}}\Big|_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $\int x^{2}(h_{1}\ast f_{2})$ (240)
$\frac{\partial\mathsf{A}^{f_{1}\ast f_{2}}}{\partial\epsilon_{1}}\Big|_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $\int(h_{1}\ast f_{2}).$ (241)
Adding 240 and 241 with Lagrange multipliers $\lambda_{1}$ and $\lambda_{2}$ to 237, we obtain
$\int(h_{1}\ast f_{2})[\ln(f_{1}\ast f_{2})+1-\lambda_{1}x^{2}-\lambda_{2}].$ (242)
Similarly, for derivative with respect to $\epsilon_{2}$, we obtain
$\int(f_{1}\ast h_{2})[\ln(f_{1}\ast f_{2})+1-\lambda_{3}x^{2}-\lambda_{4}].$ (243)
Setting 242 and 243 to zero, it follows that a Gaussian density for $f_{1}(x_{1})\ast f_{2}(x_{2})$ is a stationary point for the entropy of $\digamma=f_{1}(x_{1})\ast f_{2}(x_{2})$. Derivation is very similar to [43] (see page 335). The final conclusion is that $x_{1}+x_{2}$ is Gaussian. However, having a Gaussian density for $x_{1}+x_{2}$ does not mean $x_{1}$ and $x_{2}$ should be Gaussian as well. This problem does not occur in the case of interest here, since, having a Gaussian density for compound random variables can occur only if core random variables are Gaussian. This point is established in Theorem 3.
A.3 Effect of Scaling of Random Variables
Let us consider
$\mathsf{H}^{\digamma_{\!\!1}}-\mathsf{H}^{\digamma_{\!\!2}}$ (244)
with
$\digamma_{\!\!1}$ $=$ $f_{1}(x)\ast\frac{1}{\gamma}f_{2}\left(\frac{x}{\gamma}\right)\ast n$ (245)
$\digamma_{\!\!2}$ $=$ $f_{2}(x)\ast n$ (246)
where $f_{1}$ and $f_{2}$ are densities of $x_{1}$ and $x_{2}$, respectively, and $n$ is Gaussian. Let us consider perturbing $f_{2}$ with $\epsilon_{2}h_{2}(x)$. We have
$f_{2}(x)\ast n$ $\Longrightarrow$ $\left[f_{2}(x)+\epsilon_{2}h_{2}(x)\right]\ast n$ (247)
$f_{1}(x)\ast\frac{1}{\gamma}f_{2}\left(\frac{x}{\gamma}\right)\ast n$ $\Longrightarrow$ $f_{1}(x)\ast\frac{1}{\gamma}\left[f_{2}\left(\frac{x}{\gamma}\right)+\epsilon_{2}h_{2}\left(\frac{x}{\gamma}\right)\right]\ast n.$ (248)
It turns out that the effect of $n$ does not impact conclusions (see Appendix A.4). For simplicity of notation, $n$ is ignored in the following derivations. As a result, 247 and 248 are simplified to
$f_{2}(x)$ $\Longrightarrow$ $f_{2}(x)+\epsilon_{2}h_{2}(x)$ (249)
$\frac{1}{\gamma}f_{1}(x)\ast f_{2}\left(\frac{x}{\gamma}\right)$ $\Longrightarrow$ $\frac{1}{\gamma}f_{1}(x)\ast f_{2}\left(\frac{x}{\gamma}\right)+\frac{\epsilon_{2}}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right).$ (250)
Corresponding entropy terms are:
$-\int[f_{2}(x)+\epsilon_{2}h_{2}(x)]\ln[f_{2}(x)+\epsilon_{2}h_{2}(x)]\penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{and}$ (251)
$-\int\left[\frac{1}{\gamma}f_{1}(x)\ast f_{2}\left(\frac{x}{\gamma}\right)+\frac{\epsilon_{2}}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right]\ln\left[\frac{1}{\gamma}f_{1}(x)\ast f_{2}\left(\frac{x}{\gamma}\right)+\frac{\epsilon_{2}}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right].$ (252)
A.3.1 First Order Variations
Derivatives of 251, 252 with respect to $\epsilon_{2}$ are, respectively, equal to
$-\int h_{2}(x)\ln\left[f_{2}(x)+\epsilon_{2}h_{2}(x)\right]+h_{2}(x)\penalty 10000\ \penalty 10000\ \mbox{and}$ (253)
$-\int\left[\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right]\ln\left(\frac{1}{\gamma}f_{1}(x)\ast\left[f_{2}\left(\frac{x}{\gamma}\right)+\epsilon_{2}h_{2}\left(\frac{x}{\gamma}\right)\right]\right)+\left[\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right].$ (254)
Setting $\epsilon_{2}=0$ in 253, 254, we obtain
$-\int h_{2}(x)\ln f_{2}(x)+h_{2}(x)\penalty 10000\ \penalty 10000\ \mbox{and}$ (255)
$-\int\left(\left[\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right]\ln\left[\frac{1}{\gamma}f_{1}(x)\ast f_{2}\left(\frac{x}{\gamma}\right)\right]+\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right).$ (256)
Corresponding constraints on power are expressed as
$\int x^{2}f_{2}(x)$ $\Longrightarrow$ $\int x^{2}[f_{2}(x)+\epsilon_{2}h_{2}(x)]\penalty 10000\ \penalty 10000\ \mbox{and}$ (257)
$\int x^{2}\left[f_{1}(x)\ast\frac{1}{\gamma}f_{2}\left(\frac{x}{\gamma}\right)\right]$ $\Longrightarrow$ $\int x^{2}\left[f_{1}(x)\ast\frac{1}{\gamma}\left(f_{2}\left(\frac{x}{\gamma}\right)+\epsilon_{2}h_{2}\left(\frac{x}{\gamma}\right)\right)\right].$ (258)
Likewise, constraints on areas under density functions are expressed as
$\int f_{2}(x)$ $\Longrightarrow$ $\int f_{2}(x)+\epsilon_{2}h_{2}(x)\penalty 10000\ \penalty 10000\ \mbox{and}$ (259)
$\int\left[f_{1}(x)\ast\frac{1}{\gamma}f_{2}\left(\frac{x}{\gamma}\right)\right]$ $\Longrightarrow$ $\int\left[f_{1}(x)\ast\frac{1}{\gamma}\left(f_{2}\left(\frac{x}{\gamma}\right)+\epsilon_{2}h_{2}\left(\frac{x}{\gamma}\right)\right)\right].$ (260)
Computing derivatives of 257 and 258 with respect to $\epsilon_{2}$ and setting $\epsilon_{2}=0$ in the results, we obtain
$\int x^{2}h_{2}(x)\penalty 10000\ \penalty 10000\ \mbox{and}$ (261)
$\int x^{2}\left[\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right].$ (262)
Similar to 261 and 262, constraints on areas under density functions result in
$\int h_{2}(x)\penalty 10000\ \penalty 10000\ \mbox{and}$ (263)
$\int\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right).$ (264)
Then, using Lagrange multipliers, 261, 263 are added to 255 and 262, 264 to 256. Note that the term $h_{2}(x)$ is common in 255, 261 and 263 and can be factored out. Likewise, the term $\frac{1}{\gamma}f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)$ is common in 256, 262 and 264 and can be factored out. It follows that relying on Gaussian densities with proper variances for $f_{1}$ and $f_{2}$ results in a stationary point for the entropy terms in 251 and 252.
A.3.2 Second Order Variations
Noting 253 and 254, it follows that the second order derivative of 244 with respect to $\epsilon_{2}$, at $\epsilon_{2}=0$, is equal to
$\frac{\left[f_{1}(x)\ast h_{2}\left(\frac{x}{\gamma}\right)\right]^{2}}{\gamma f_{1}(x)\ast f_{2}\left(\frac{x}{\gamma}\right)}+\frac{\left[h_{2}(x)\right]^{2}}{f_{2}(x)}.$ (265)
As will be discussed in Appendix B, for the objective function $\mathbf{\Upsilon}$ defined in 9, perturbations are formed using functions $\epsilon_{i}h_{i}$. Each second order derivative of the form
$\frac{\partial^{2}\mathbf{\Upsilon}}{\partial\epsilon_{i}^{2}}$ (266)
is composed of multiple terms, each of the form given in 265. The term corresponding to perturbation $\epsilon_{i}h_{i}$ will be zero only if $h_{i}=0$. This means collection of Gaussian density functions for compound random variables, each obtained from
$\frac{\partial\mathbf{\Upsilon}}{\partial\epsilon_{i}}=0\penalty 10000\ \penalty 10000\ \mbox{at}\penalty 10000\ \penalty 10000\ \epsilon_{i}=0$ (267)
result in a non-zero value for 266. This means the corresponding stationary solution is either a minimum or a maximum.
A.4 Functional of Compound Random Variables
Let us assume $f_{1}(x_{1})$ and $f_{2}(x_{2})$ are density functions for two core random variables, forming compound random variables $x_{1}+x_{2}$ and $x_{2}$. Let us define
$\digamma_{\!\!1}$ $=$ $f_{1}\ast f_{2}\ast n$ (268)
$\digamma_{\!\!2}$ $=$ $f_{2}\ast n$ (269)
where $n$ is the probability density function of the additive Gaussian noise. Then,
$\mathsf{H}^{\digamma_{\!\!1}}-\mathsf{H}^{\digamma_{\!\!2}}$ (270)
is the mutual information over an additive noise channel where $f_{1}\ast f_{2}\ast n$ is the channel output, and $f_{2}\ast n$ is the additive noise. We are interested to find a stationary solution for 270. In the following, variations of $\digamma_{\!\!1}$, $\digamma_{\!\!2}$ are denoted as $p\digamma_{\!\!1}$, $p\digamma_{\!\!2}$, respectively. Constrains on power are expressed as:
$\mathsf{E}^{f_{1}\ast f_{2}\ast n}$ $=$ $\int x^{2}(f_{1}\ast f_{2}\ast n)\penalty 10000\ \penalty 10000\ \mbox{is a constant}$ (271)
$\mathsf{E}^{f_{2}\ast n}$ $=$ $\int x^{2}(f_{2}\ast n)\penalty 10000\ \penalty 10000\ \mbox{is a constant}$ (272)
$\mathsf{A}^{f_{1}\ast f_{2}\ast n}$ $=$ $\int(f_{1}\ast f_{2}\ast n)=1.$ (273)
$\mathsf{A}^{f_{2}\ast n}$ $=$ $\int(f_{2}\ast n)=1.$ (274)
We have
$\frac{\partial\mathsf{H}^{p\digamma_{\!\!1}}}{\partial\epsilon_{1}}\Big|_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $-\int(h_{1}\ast f_{2}\ast n)[\ln(f_{1}\ast f_{2}\ast n)+1]$ (275)
$\frac{\partial\mathsf{E}^{p\digamma_{\!\!1}}}{\partial\epsilon_{1}}\Big|_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $\int x^{2}(h_{1}\ast f_{2}\ast n)$ (276)
$\frac{\partial\mathsf{A}^{p\digamma_{\!\!1}}}{\partial\epsilon_{1}}\Big|_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $\int h_{1}\ast f_{2}\ast n$ (277)
$\frac{\partial\mathsf{H}^{p\digamma_{\!\!1}}}{\partial\epsilon_{2}}\Big|_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $-\int(f_{1}\ast h_{2}\ast n)[\ln(f_{1}\ast f_{2}\ast n)+1]$ (278)
$\frac{\partial\mathsf{E}^{p\digamma_{\!\!1}}}{\partial\epsilon_{2}}\Big|_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $\int x^{2}(f_{1}\ast h_{2}\ast n)$ (279)
$\frac{\partial\mathsf{A}^{p\digamma_{\!\!1}}}{\partial\epsilon_{2}}\Big|_{\epsilon_{1}=0,\epsilon_{2}=0}$ $=$ $\int f_{1}\ast h_{2}\ast n$ (280)
and
$\frac{\partial\mathsf{H}^{p\digamma_{\!\!2}}}{\partial\epsilon_{2}}\Big|_{\epsilon_{2}=0}$ $=$ $-\int(h_{2}\ast n)\ln(f_{2}\ast n)$ (281)
$\frac{\partial\mathsf{E}^{p\digamma_{\!\!2}}}{\partial\epsilon_{2}}\Big|_{\epsilon_{2}=0}$ $=$ $\int x^{2}(h_{2}\ast n)$ (282)
$\frac{\partial\mathsf{A}^{p\digamma_{\!\!2}}}{\partial\epsilon_{2}}\Big|_{\epsilon_{2}=0}$ $=$ $\int h_{2}\ast n.$ (283)
Adding 276, 277 with Lagrange multipliers to 275; 279, 280 with Lagrange multipliers to 278; and 282, 283 with Lagrange multipliers to 281, then setting the results to zero, it follows that Gaussian distributions for $f_{1}\ast f_{2}$ and $f_{2}$ result in a stationary solution for 270. Denoting arguments of $f_{1}$, $f_{2}$, $n$ as $x_{1}$, $x_{2}$, $z$, respectively, this entails random variables $y_{1}=x_{1}+x_{2}+z$ and $y_{2}=x_{2}+z$ are jointly Gaussian, and consequently, $w_{1}=x_{1}+x_{2}$ and $w_{2}=x_{2}$ are jointly Gaussian as well. In general, if a linear combination of some random variables is Gaussian, it does not necessarily mean each random variable should be Gaussian as well. However, in this example, we can uniquely express $x_{1}$, $x_{2}$ in terms of $w_{1}$, $w_{2}$, i.e., $x_{1}=w_{1}-w_{2}$ and $x_{2}=w_{2}$. This entails $x_{1}$ and $x_{2}$ should be Gaussian as well.
Remark 7: Let use rely on indices $\{1,\ldots,\mathbf{c}_{1}\}$, $\{1,\ldots,\mathbf{c}_{2}\}$ to specify elements of core and compound random variables, respectively. In this work, $\mathbf{c}_{2}\geq\mathbf{c}_{1}$. To obtain a stationary solution, it is shown that compound random variables should be jointly Gaussian. In addition, there are subset(s) of $\{1,\ldots,\mathbf{c}_{2}\}$ of size $\mathbf{c}_{1}$ such that corresponding matrix of linear coefficients is full rank. This property allows expressing core random variables as a linear combination of a subset of compound random variables. Also, expressions corresponding to different subsets of size $\mathbf{c}_{1}$ from $\{1,\ldots,\mathbf{c}_{2}\}$ are consistent. The conclusion is that each core random variable should be Gaussian. $\blacksquare$
Appendix B Stationary Solutions for $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$
This appendix shows that density functions resulting in stationary solutions for $\mathbf{\Upsilon}$ and $\mathbf{\Gamma}$ are Gaussian. Let us focus on $\mathbf{\Upsilon}$, since the derivation for $\mathbf{\Gamma}$ is very similar. Terms forming the numerator and the denominator of $\mathbf{\Upsilon}$ are rate across channels with additive noise (see Fig. 3). Each entropy term, corresponding to a compound random variable, is based on the convolution of densities of the underlying core random variables. As will be shown in Appendix A.3, scale factors for core random variables do not affect the derivations to follow. For this reason, such scale factors are not included in this Appendix. Notation $\circledast_{i=1}^{\mathfrak{q}}\mathfrak{p}_{i}$ denotes the convolution $\mathfrak{p}_{1}\ast\mathfrak{p}_{2}\ast...\ast\mathfrak{p}_{q}$, called multi-convolution hereafter. Recall that calculus of variation is applied by perturbing density function of each core random variable. For this reason, multi-convolution terms which involve same core random variables appear in the derivations. Since derivatives related to perturbation of different core random variables are handled separately, one can limit derivations to multi-convolution terms which have (at least) one common core random variable, denoted by the generic notation $\mathfrak{g}$ hereafter. If multi-convolution terms include two or more common core random variables, say $\mathfrak{g}_{1}$ and $\mathfrak{g}_{2}$, since such common terms are perturbed separately, derivations for each term will be similar to what is presented here. It is also enough to consider only four entropy terms (to define a reduced/generic expression for $\mathbf{\Upsilon}$) as given in 284. Derivations for more general cases will be similar (due to linearity of multi-convolution with respect to its terms).
$\frac{\mathfrak{N}}{\mathfrak{D}}=\frac{\displaystyle-\int\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}\right)+\int\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}\right)}{\displaystyle-\int\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}\right)+\int\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}\right)}.$ (284)
Let us perturb $\mathfrak{g}\Rightarrow\mathfrak{g}+\ell\mathfrak{h}$ resulting in
$\!\int\!\!\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}\right)$ $\!\!\Rightarrow\!\!$ $\!\int\!\!\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{h}\right)\log\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{h}\right)$
$\!\int\!\!\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}\right)$ $\!\!\Rightarrow\!\!$ $\!\int\!\!\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{h}\right)\log\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{h}\right)$
$\!\int\!\!\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}\right)$ $\!\!\Rightarrow\!\!$ $\!\int\!\!\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{h}\right)\log\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{h}\right)$
$\!\!\!\int\!\!\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}\right)$ $\!\!\Rightarrow\!\!$ $\!\!\int\!\!\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{h}\right)\log\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{h}\right).$ (285)
Derivatives of right hand terms in 285, with respect to $\ell$, are equal to
$\textbf{\ T}_{1}(\ell)\penalty 10000\ =\penalty 10000\ \frac{\partial}{\partial\ell}\int\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{h}\right)\log\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{h}\right)$ $=$
$\int\circledast_{i=1}^{\mathfrak{m}}(\mathfrak{a}_{i}\ast\mathfrak{h})\log\left(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{h}\right)+\circledast_{i=1}^{\mathfrak{m}}(\mathfrak{a}_{i}\ast\mathfrak{h})$ (286)
$\textbf{\ T}_{2}(\ell)\penalty 10000\ =\penalty 10000\ \frac{\partial}{\partial\ell}\int\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{h}\right)\log\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{h}\right)$ $=$
$\int\circledast_{i=1}^{\mathfrak{n}}(\mathfrak{b}_{i}\ast\mathfrak{h})\log\left(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{h}\right)+\circledast_{i=1}^{\mathfrak{n}}(\mathfrak{b}_{i}\ast\mathfrak{h})$ (287)
$\textbf{\ T}_{3}(\ell)\penalty 10000\ =\penalty 10000\ \frac{\partial}{\partial\ell}\int\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{h}\right)\log\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{h}\right)$ $=$
$\int\circledast_{i=1}^{\mathfrak{p}}(\mathfrak{e}_{i}\ast\mathfrak{h})\log\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{h}\right)+\circledast_{i=1}^{\mathfrak{p}}(\mathfrak{e}_{i}\ast\mathfrak{h})$ (288)
$\textbf{\ T}_{4}(\ell)\penalty 10000\ =\penalty 10000\ \frac{\partial}{\partial\ell}\int\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{h}\right)\log\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{h}\right)$ $=$
$\int\circledast_{i=1}^{\mathfrak{q}}(\mathfrak{f}_{i}\ast\mathfrak{h})\log\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{h}\right)+\circledast_{i=1}^{\mathfrak{q}}(\mathfrak{f}_{i}\ast\mathfrak{h}).$ (289)
It follows that
$\textbf{\ T}_{1}(0)$ $=$ $\int(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{h})[1+\log(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g})]$ (290)
$\textbf{\ T}_{2}(0)$ $=$ $\int(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{h})[1+\log(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g})]$ (291)
$\textbf{\ T}_{3}(0)$ $=$ $\int(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{h})[1+\log(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g})]$ (292)
$\textbf{\ T}_{4}(0)$ $=$ $\int(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{h})[1+\log(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g})].$ (293)
Noting the expression for $\mathfrak{D}$ in 284, let us define
$\textbf{\ D}_{1}$ $=$ $-\int\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}\right)$ (294)
$\textbf{\ D}_{2}$ $=$ $\int\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}\right)\log\left(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}\right)$ (295)
$\mathfrak{D}$ $=$ $\textbf{\ D}_{1}+\textbf{\ D}_{2}.$ (296)
It follows that
[…]
where
[…]
Constraints on power corresponding to terms  $\textbf{\ T}_{1}(\ell)$,  $\textbf{\ T}_{2}(\ell)$,  $\textbf{\ T}_{3}(\ell)$ and  $\textbf{\ T}_{4}(\ell)$ are
$\int x^{2}(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g})$ $\Rightarrow$ $\int x^{2}(\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}\ast\mathfrak{h})$ (302)
$\int x^{2}(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g})$ $\Rightarrow$ $\int x^{2}(\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}\ast\mathfrak{h})$ (303)
$\int x^{2}(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g})$ $\Rightarrow$ $\int x^{2}(\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}\ast\mathfrak{h})$ (304)
$\int x^{2}(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g})$ $\Rightarrow$ $\int x^{2}(\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{g}+\ell\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}\ast\mathfrak{h}).$ (305)
Computing the derivatives of above terms with respect to $\ell$ for $\ell=0$, and including Lagrange multipliers $\varsigma_{1}$, $\varsigma_{2}$, $\varsigma_{3}$ and $\varsigma_{4}$, we obtain
$\varsigma_{1}\int x^{2}(\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i})$ (306)
$\varsigma_{2}\int x^{2}(\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i})$ (307)
$\varsigma_{3}\int x^{2}(\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i})$ (308)
$\varsigma_{4}\int x^{2}(\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{q}}\mathfrak{f}_{i}).$ (309)
Likewise, constraints on areas under density functions can be expressed as
$\iota_{1}\int\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{m}}\mathfrak{a}_{i}$ (310)
$\iota_{2}\int\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{n}}\mathfrak{b}_{i}$ (311)
$\iota_{3}\int\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{p}}\mathfrak{e}_{i}$ (312)
$\iota_{4}\int\mathfrak{h}\ast\circledast_{i=1}^{\mathfrak{q}}\mathfrak{h}_{i}$ (313)
[…]
Appendix C Detailed Derivations - First Step
The point on the capacity region with maximum $R_{1}$ is achieved at a corner point using Gaussian densities where user 1 allocates its power $P_{1}$ to a private message and user 2 allocates its power $P_{2}$ to a public message. Stating from this corner point, density functions at the end point of the first incremental step are studied. With some misuse of notations, in specifying the entropy of a compound random variable, the subscript in $H$ shows the corresponding linear combination, the superscripts $s$ and $e$ show if it is a starting point or an end point, and the argument shows the total power, e.g., $H^{s}_{V_{1}+\sqrt{b}U_{2}+Z}(P_{1}+bP_{2}+1)$ denotes the entropy of $V_{1}+\sqrt{b}U_{2}+Z$ at the starting point, where the power values of $V_{1}$, $U_{2}$ are equal to $P_{1}$, $P_{2}$, respectively. Notations $R^{s}_{U_{1}}(.)$, $R^{s}_{U_{2}}(.)$, $R^{s}_{V_{1}}(.)$, $R^{s}_{V_{2}}(.)$ and $R^{e}_{U_{1}}(.)$, $R^{e}_{U_{2}}(.)$, $R^{e}_{V_{1}}(.)$, $R^{e}_{V_{2}}(.)$ refer to the rate associated with $U_{1}$}, $U_{2}$}, $V_{1}$}, $V_{2}$} at the starting point and at the end point on a step, respectively (function of relevant power values). Movement is achieved by reallocating a small power value of $\delta P_{2}$ form $U_{2}$ to $V_{2}$. Figure 7 shows such a power reallocation. For the first step, we have:
$R^{s}_{U_{1}}$ $=$ $0$ (314)
$R^{s}_{V_{1}}$ $=$ ${\complement}(P_{1},1)$ (315)
$R^{s}_{U_{2}}$ $=$ $\complement(bP_{2},P_{1}+1)$ (316)
$R^{s}_{V_{2}}$ $=$ $0$ (317)
where
$\complement(\alpha,\beta)=0.5\log_{2}\left(1+\frac{\alpha}{\beta}\right).$ (318)
At the end point, $U_{2}$ at $Y_{1}$ is subject to the noise
$\frac{1}{\sqrt{b}}V_{1}+V_{2}+\frac{1}{\sqrt{b}}Z$ (319)
while $U_{2}$ at $Y_{2}$ is subject to the noise
$\sqrt{a}V_{1}+V_{2}+Z.$ (320)
Comparing 319 with 320, since $a<1$ and $b<1$, it is concluded that the rate of $U_{2}$ is governed by the mutual information between $U_{2}$ and $Y_{1}$. As a result
$R^{e}_{U_{1}}$ $=$ $0$ (321)
$R^{e}_{V_{1}}$ $=$ $H^{e}_{V_{1}+\sqrt{b}V_{2}+Z}(P_{1}+b\delta P_{2}+1)-H^{e}_{\sqrt{b}V_{2}+Z}(b\delta P_{2}+1)$ (322)
$R^{e}_{U_{2}}$ $=$ $H^{e}_{V_{1}+\sqrt{b}V_{2}+\sqrt{b}U_{2}+Z}(P_{1}+bP_{2}+1)-H^{e}_{V_{1}+\sqrt{b}V_{2}+Z}(P_{1}+b\delta P_{2}+1)$ (323)
$R^{e}_{V_{2}}$ $=$ $H^{e}_{\sqrt{a}V_{1}+V_{2}+Z}(aP_{1}+\delta P_{2}+1)-H^{e}_{\sqrt{a}V_{1}+Z}(aP_{1}+1)$ (324)
and
$\mathbf{\Upsilon}=\frac{R^{e}_{U_{2}}+R^{e}_{V_{2}}-R^{s}_{U_{2}}-R^{s}_{V_{2}}}{R^{s}_{U_{1}}+R^{s}_{V_{1}}-R^{e}_{U_{1}}-R^{e}_{V_{1}}}.$ (325)
We have
$R^{e}_{U_{2}}+R^{e}_{V_{2}}-R^{s}_{U_{2}}-R^{s}_{V_{2}}$ $=$ (326)
$H^{e}_{V_{1}+\sqrt{b}U_{2}+\sqrt{b}V_{2}+Z}(P_{1}+bP_{2}+1)-H^{e}_{V_{1}+\sqrt{b}V_{2}+Z}(P_{1}+b\delta P_{2}+1)$ $+$ (327)
$H^{e}_{\sqrt{a}V_{1}+V_{2}+Z}(aP_{1}+\delta P_{2}+1)-H^{e}_{\sqrt{a}V_{1}+Z}(aP_{1}+1)$ - (328)
$\complement(bP_{2},P_{1}+1)$ (329)
and
$R^{s}_{U_{1}}+R^{s}_{V_{1}}-R^{e}_{U_{1}}-R^{e}_{V_{1}}=\complement(P_{1},1)-H^{e}_{V_{1}+\sqrt{b}V_{2}+Z}(P_{1}+b\delta P_{2}+1)+H^{e}_{\sqrt{b}V_{2}+Z}(b\delta P_{2}+1).$ (330)
Compound random variables appearing in 321330 are
$\hat{C}_{1}$ $=$ $\sqrt{a}V_{1}+V_{2}$ (331)
$\hat{C}_{2}$ $=$ $\sqrt{a}V_{1}$ (332)
$\hat{C}_{3}$ $=$ $V_{1}+\sqrt{b}U_{2}+\sqrt{b}V_{2}$ (333)
$\hat{C}_{4}$ $=$ $V_{1}+\sqrt{b}V_{2}$ (334)
$\hat{C}_{5}$ $=$ $\sqrt{b}V_{2}.$ (335)
To find density functions for the end point, we rely on calculus of variations. Each compound random variable is accompanied by a constraint on its second moment, and a constraint on the area under its density. Relying on calculus of variation, it is concluded that densities of compound random variables are Gaussian. Using equations 331, 332, 333; or 332, 333, 334; or 332, 333, 335; or 333, 334, 335 from 331, 332, 333, 334, 335, one can express $V_{1},U_{2},V_{2}$ in terms of three of the compound random variables. The existence of such an invertible mapping means $V_{1},U_{2},V_{2}$ should be Gaussian as well. Figure 7 depicts the structure of Gaussian code-books for the start and end points. Note that, since $\delta P_{1}=0$, the condition of Pareto minimality is satisfied.
Figure 7: First step moving counterclockwise from the corner point with maximum $R_{1}$. (a),(b) correspond to the starting point, and (c) corresponds to the end point on the first step.